### v1

# 22affiliationtext: Booth School of Business

Abstract
In designing generative models, it is commonly believed that in order to learn useful latent structure, we face a fundamental tension between expressivity and structure. In this paper we challenge this view by proposing a new approach to training arbitrarily expressive generative models that simultaneously learn disentangled latent structure. This is accomplished by adding a simple decoder-only module to the head of an existing decoder block that can be arbitrarily complex. The module learns to process concept information by implicitly inverting linear representations from an encoder. Inspired by the notion of intervention in causal graphical models, our module selectively modifies its architecture during training, allowing it to learn a compact joint model over different contexts. We show how adding this module leads to disentangled representations that can be composed for out-of-distribution generation. To further validate our proposed approach, we prove a new identifiability result that extends existing work on identifying structured representations in nonlinear models.
1 Introduction
Generative models have transformed information processing and demonstrated remarkable capacities for creativity in a variety of tasks ranging from vision to language to audio. The success of these models has been largely driven by modular, differentiable architectures based on deep neural networks that learn useful representations for downstream tasks. Recent years have seen increased interest in understanding and exploring the representations produced by these models through evolving lines of work on structured representation learning, identifiability and interpretability, disentanglement, and causal generative models. This work is motivated in part by the desire to produce performant generative models that also capture meaningful, semantic latent spaces that enable out-of-distribution generation under perturbations.
A key conceptual driver of this line of work is the tradeoff between flexibility and structure, or expressivity and interpretability: It is widely believed that to learn structured, interpretable representations, model capacity must be constrained, sacrificing flexibility and expressivity. As described in Section 3, this intuition is supported by a growing body of theoretical work on nonlinear ICA (Hyvärinen and Pajunen, 1999), disentanglement (Bengio, 2013), and causal representation learning (Schölkopf et al., 2021). On the practical side, methods that have been developed to learn structured latent spaces tend to be bespoke to specific data types and models, and typically impose significant limitations on the flexibility and expressivity of the underlying models. Moreover, the most successful methods for learning structured representations typically impose fixed, known structure a priori, as opposed to learning this structure from data.
At the same time, there is a growing body of work that suggests generative models already learn surprisingly structured latent spaces (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015, see also Section 3). This in turn suggests that existing models are already “close” to capturing the desired structure, and perhaps only small modifications are needed. Since we already have performant models that achieve state-of-the-art results in generation and discrimination, do we need to re-invent the wheel to achieve these goals? Our hypothesis is that we should be able to leverage the expressiveness of these models to build new models that learn latent structure from scratch. We emphasize that our goal is not to explain or interpolate the latent space of a pre-trained model, but rather to leverage known architectures to train a new model end-to-end from scratch.
Figure 1: Overview of our approach. Given a black-box encoder-decoder architecture (blue), we propose to append a context module (red) to the head of the decoder. (left) Instead of passing the output of the encoder directly to the decoder (blue box + blue arrow), the embeddings $\mathbf{e}$ are passed through the context module, consisting of three distinct layers. The output of this module is then passed into the decoder. (right) The model learns to compose different concepts OOD, in this case object and background colour, to values that never appear in the training data.
In this paper, we adopt this perspective: We start with a black-box encoder-decoder architecture, upon which we make no assumptions, and then augment this model to learn structured representations. The idea is that the black-box is an architecture that is already known to perform well on downstream tasks such as sampling and prediction, and so is flexible enough to capture complex patterns in data. We then introduce a modular, end-to-end differentiable architecture for learning structured representations, augmenting the original model with a simple decoder-only module that retains its existing capabilities while enabling concept identification and intervention as well as composing together multiple concept interventions for out-of-distribution (OOD) generation. The resulting latent structure is learned fully end-to-end, with no fixed structure imposed a priori. The motivation is to provide a framework for taking existing performant architectures and to train a new model that performs just as well, but with the added benefit of learning structured representations without imposing specific prior knowledge or structure a priori. These representations may be useful for interpretability, downstream tasks, or other purposes. Since the model, unlike previous approaches, imposes no latent structure a priori, we theoretically validate the approach with a novel identifiability result for structured representations recovered by the model.
To explore the implications of this modification, we run experiments including OOD generation. An attractive feature of our approach is that it makes running carefully controlled experiments and ablations particularly simple: We can carefully control hyperparameters across different models, ensuring that any differences in performance are attributable to specifically controlled architectural modifications. As a result, we contribute to a growing body of empirical work on the OOD capabilities of generative models by introducing a new set of ablations that can be used to guide the development of new methods.
Contributions
Our main contributions can be summarized as follows:
  1. 1.
    We introduce a simple module that can be attached to the head of an existing decoder block to learn concept representations by splitting each concept into a tensor slice, where each slice represents an interventional context in a reduced form structural equation model (SEM). This module can be attached to the head of any decoder block and trained end-to-end.
  2. 2.
    We conduct experiments on disentanglement and OOD generation on several benchmark datasets (3DIdent, CelebA, MNIST) as well as controlled simulations. To this end, we introduce a simple simulated visual environment for testing disentanglement and OOD generation. This allows us to carefully test different ablations under controlled settings.
  3. 3.
    To illustrate the adaptivity of our module to different architectures, we include experiments with several different VAE architectures including NVAE (Vahdat and Kautz, 2020).
As a matter of independent interest, we prove an identifiability result under concept interventions, and discuss how the proposed architecture can be interpreted as an approximation to this model. Full technical proofs are deferred to the appendix.
2 Preliminaries
Let $\mathbf{x}$ denote the observations, e.g., pixels in an image, and $\mathbf{z}$ denote hidden variables, e.g. latent variables that are to be inferred from the pixels. We are interested in training generative models with an encoder-decoder architecture, such as a variational autoencoder (VAE). A typical generative model consists of a decoder $p_{\theta}(\mathbf{x}\,|\,\mathbf{z})$ and an encoder $q_{\phi}(\mathbf{z}\,|\,\mathbf{x})$. After specifying a prior $p_{\theta}(\mathbf{z})$, this defines a likelihood by
[…]
Both the decoder and encoder are specified by deep neural networks that will be trained end-to-end via standard techniques (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015). These networks are often chosen to capture implicit assumptions about the structure of the latent space, such as hierarchical or graphical structure, equivariance, etc. In this work, we focus on VAEs since they are often use to represent structured and semantic latent spaces, which is our focus. Unlike traditional structured VAEs, we do not impose a specific structure a priori.
Our starting point is the following incongruence: On the one hand, it is well-established that the latent spaces of VAEs are typically entangled and semantically misaligned, fail to generalize OOD, and suffer from posterior collapse (which has been tied to latent variable nonidentifiability, Wang et al., 2021). On the other hand, VAEs still learn highly structured latent spaces that can be traversed and interpolated, are “nearly” identifiable (Willetts and Paige, 2021; Reizinger et al., 2022), and represent abstract concepts linearly. Thus, our hypothesis is that sufficiently flexible models do capture meaningful structure internally, just not in a way that is interpretable or meaningful in practice. So, to learn latent structure, we build off the already performant embeddings of a generative model.
Our set-up is the following: We are given a black-box VAE, consisting of an encoder and a decoder, on which we make no assumptions other than the encoder outputs a latent code corresponding to $\mathbf{x}$. In the notation of (2), this corresponds to $\mathbf{z}$, however, to distinguish the black-box embeddings from our model, we will denote the black-box embeddings hereafter by $\mathbf{e}$. Our plan is to work solely with the embeddings $\mathbf{e}$ and learn how to extract linear concept representations from $\mathbf{e}$ such that distinct concepts can be intervened upon and composed together to create novel, OOD samples. Crucially, we do not modify the black-box architecture in any way.
3 Related work
Structured generative models
To provide structure such as hierarchical, graphical, causal, and disentangled structures as well as other inductive biases in the latent space, there has been a trend towards building structured generative models that directly impose this structure a priori. Early work looked at incorporating fixed, known structure into generative models, such as autoregressive, graphical, and hierarchical structure (Germain et al., 2015; Johnson et al., 2016; Sønderby et al., 2016; Webb et al., 2018; Weilbach et al., 2020; Ding et al., 2021; Mouton and Kroon, 2023). This was later translated into known causal structure (Kocaoglu et al., 2017; Markham et al., 2023). When the latent structure is unknown, several techniques have been developed to learn useful (not necessarily causal) structure from data (Li et al., 2018; He et al., 2019; Wehenkel and Louppe, 2021; Kivva et al., 2022; Moran et al., 2022). More recently, based on growing interest in disentangled (Bengio, 2013) and/or causal (Schölkopf et al., 2021) representation learning, methods that learn causal structure have been developed (Markham and Grosse-Wentrup, 2020; Moraffah et al., 2020; Yang et al., 2021; Ashman et al., 2022; Shen et al., 2022; Kaltenpoth and Vreeken, 2023).
Causal representation learning
Causal representation learning (Schölkopf et al., 2021), which involves learning causal structure in the latent space, is a rapidly developing area that has produced fundamental results on theoretical aspects of identifiability (Brehmer et al., 2022; Shen et al., 2022; Lachapelle et al., 2022; Moran et al., 2022; Kivva et al., 2022; Buchholz et al., 2023; Gresele et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023; Morioka and Hyvärinen, 2023; Liu et al., 2023; Mameche et al., 2023; Yao et al., 2023; Varici et al., 2023; Sturma et al., 2023; Xu et al., 2024; Zhang et al., 2024; Li et al., 2024; Varici et al., 2024; Bing et al., 2024; Ahuja et al., 2024; Talon et al., 2024; Yao et al., 2024). Recent work has also pushed in the direction of identifying concepts (Leemann et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025). Our work draws inspiration from this line of work, which articulates precise conditions under which latent structure can be recovered in principle. By contrast, our focus is on methodological aspects of exploiting learned causal invariances in practice.
Linear representations
Generative models are known to represent concepts linearly in embedding space (e.g. Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015); see also Levy and Goldberg (2014); Arora et al. (2016); Gittens et al. (2017); Allen and Hospedales (2019); Ethayarajh et al. (2018); Seonwoo et al. (2019). This phenomenon has been well-documented over the past decade in both language models (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Conneau et al., 2018; Tenney et al., 2019; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Li et al., 2023; Park et al., 2023; Gurnee et al., 2023; Jiang et al., 2024) and computer vision (Radford et al., 2015; Raghu et al., 2017; Bau et al., 2017; Engel et al., 2017; Kim et al., 2018; Chen et al., 2020; Wang et al., 2023; Trager et al., 2023). Our approach actively exploits this tendency by searching for concept representations as linear projections of the embeddings learned by a black-box model.
OOD generalization
A growing line of work studies the OOD generalization capabilities of generative models, with the general observation being that existing methods struggle to generalize OOD (Xu et al., 2022; Mo et al., 2024; Montero et al., 2022, 2021; Schott et al., 2022). It is worth noting that most if not all of this work evaluates OOD generalization using reconstruction on held-out OOD samples, as opposed to generation. For example, a traditional VAE may be able to reconstruct held-out samples, but it is not possible to actively sample OOD. See Section 5.1 for more discussion.
4 Architecture
We start with a black-box encoder-decoder architecture and assume there are $d_{c}$ concepts of interest, denoted by $\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$. Our objectives are two-fold:
  1. 1.
    To learn structure between concepts from black-box embeddings $\mathbf{e}$ as linear projections $C\mathbf{e}$;
  2. 2.
    To compose concepts together in a single, transparent model that captures how different concepts are related.
We seek to learn these concepts in an end-to-end, differentiable manner.
A key intuition is that composition can be interpreted as a type of intervention in the latent space. This is sensible since interventions in a causal model are a type of nontrivial distribution shift. Thus, the problem takes on a causal flavour which we exploit to build our architecture. The difficulty with this from the causal modeling perspective is that encoding structural assignments and/or causal mechanisms directly into a feed-forward neural network is tricky, because edges between nodes within the same layer aren’t allowed in a feed-forward network but are required for the usual DAG representation of a causal model. To bypass this, we use the reduced form of a causal model which has a clear representation as a bipartite directed graph, with arrows only from exogenous to endogenous variables, making it conducive to being embedded into a feedforward NN. The tradeoff is that we do not learn a causal structure (i.e., a causal DAG), however, this enables the model to perform interventions directly in the latent space and to exploit invariances between interventional contexts.
4.1 Overview
Before outlining the architectural details, we provide a high-level overview of the main idea. A traditional decoder transforms the embeddings $\mathbf{e}$ into the observed variables $\mathbf{x}$, and the encoder operates in reverse by encoding $\mathbf{x}$ into $\mathbf{e}$. Thus,
[…]
Consistent with existing empirical work (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015), we represent concepts as linear projections of the embeddings $\mathbf{e}$: Each concept $\mathbf{c}_{j}$ can be approximated as $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$. This is modeled via a linear layer $\mathbf{c}\to\mathbf{e}$ that implicitly inverts this relationship in the decoder.
As a result, it makes sense to model the relationships between concepts with a linear SEM:
$\mathbf{c}_{j}=\sum_{k=1}^{d_{c}}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{k},\quad\alpha_{kj}\in\mathbb{R}.$
By reducing this SEM, we deduce that
$\mathbf{c}=A_{0}\boldsymbol{\varepsilon},\text{ where }\left\{\begin{aligned} \mathbf{c}&=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})\\ \boldsymbol{\varepsilon}&=(\boldsymbol{\varepsilon}_{1},\ldots,\boldsymbol{\varepsilon}_{d_{c}})\end{aligned}\right..$ (1)
We model this with a linear layer $\boldsymbol{\varepsilon}\to\mathbf{c}$, where the weights in this layer correspond to the matrix $A_{0}$. This layer will be used to encode the SEM between the concepts, which will be used to implement causal interventions.
Remark 4.1.
Due to the reduced form SEM above, our approach does not and cannot model the structural causal model encoded by the $\alpha_{kj}$. What is important is that the reduced form $\mathbf{c}=A_{0}\boldsymbol{\varepsilon}$ still encodes causal invariances and interventions, which is enough in our setting, without directly estimating a causal graph.
In principle, since the exogenous variables $\boldsymbol{\varepsilon}_{j}$ are independent, we could treat $\boldsymbol{\varepsilon}$ as the input latent space to the generative model. Doing this, however, incurs two costs: 1) To conform to standard practice, $\boldsymbol{\varepsilon}$ would have to follow an isotropic Gaussian prior, and 2) It enforces artificial constraints on the latent dimension $\dim(\boldsymbol{\varepsilon})$. For this reason, we use a second expressive layer $\mathbf{z}\to\boldsymbol{\varepsilon}$ that gradually transforms $\mathbf{z}\sim\mathcal{N}(0,I)$ into $\boldsymbol{\varepsilon}$. This allows for $\dim(\mathbf{z})$ to be larger and more expressive than $d_{c}$ (in practice, we set $\dim(\mathbf{z})$ to be a multiple of $\dim(\boldsymbol{\varepsilon})$), and for $\boldsymbol{\varepsilon}$ to be potentially non-Gaussian.
The final decoder architecture can be decomposed, at a high-level, as follows:
$\mathbf{z}\longrightarrow\boldsymbol{\varepsilon}\longrightarrow\mathbf{c}\longrightarrow\mathbf{e}\longrightarrow\mathbf{x}.$ (2)
Implementation details for each of these layers can be found in the next section.
4.2 Details
Our goal is to augment the decoder of a generative model (without modifying the internals of the existing decoder) in such a way to identify $d_{c}$ latent concepts from a black-box decoder. Moreover, we seek to accomplish this without sacrificing expressivity: The input layer to the decoder as well as the output layer of the encoder should both be arbitrary. This not only allows for arbitrary flexibility in both the latent dimension as well as the structure of the black-box encoder-decoder and its latent space, but also makes implementation significantly simpler: The learning of latent concepts is completely abstracted away from the encoding step.
To accomplish this, we assume given a black-box encoder-decoder pair, denoted by $\operatorname{enc}(\mathbf{x})$ and $\operatorname{dec}(\mathbf{z})$, respectively. We modify the decoder by appending a context module to the head of the decoder. This context module is divided into three layers: A representation layer, an intervention layer, and an expressive layer.
  1. 1.
    The first representation layer, as the name suggests, learns to represent concepts by implicitly inverting their linear representations $\mathbf{c}_{j}=C_{j}\mathbf{e}$ from the embeddings of the encoder. This is a linear layer between $\mathbf{c}\to\mathbf{e}$.
  2. 2.
    The intervention layer embeds these concept representations into an overall reduced form SEM as suggested by (1). Each concept corresponds to its own context where it has been intervened upon. A key part of this layer is how it can be used to learn and enforce interventional semantics through this shared SEM. This layer corresponds to the second layer $\boldsymbol{\varepsilon}\to\mathbf{c}$ in (2).
  3. 3.
    The expressive layer is used to reduce the (potentially very large) input latent dimension of independent Gaussian inputs down to a smaller space of non-Gaussian exogenous noise variables for the intervention layer, corresponding to the first layer $\mathbf{z}\to\boldsymbol{\varepsilon}$ in (2). This is implemented as independent, deep MLPs that gradually reduce the dimension in each layer: If it is desirable to preserve Gaussianity for the exogenous variables, linear activations can be used in place of nonlinear activations (e.g., ReLU).
Because the intervention layer is modeled after an SEM, it is straightforward to perform latent concept interventions using the calculus of interventions in an SEM.
Remark 4.2.
Crucially, no part of this SEM is fixed or known—everything is trained end-to-end. In particular, we do not assume a known causal DAG or even a known causal order. This stands in contrast to previous work on causal generative models such as CausalGAN (Kocaoglu et al., 2017) and CausalVAE (Yang et al., 2021).
Concept interventions
For simplicity, assume here that the concepts are each one-dimensional with $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$; generalization to multi-dimensional concepts is more or less straightforward and explained below. The structural coefficients $\alpha_{kj}\in\mathbb{R}$ capture direct causal effects between concepts, with $\alpha_{kj}\neq 0$ indicating the presence of an edge $\mathbf{c}_{k}\to\mathbf{c}_{j}$. An intervention on the $j$th concept requires deleting all these incoming edges; i.e., setting $\alpha_{\cdot j}=0$, updating the outgoing edges from $\boldsymbol{\varepsilon}_{j}$, as well as replacing $\boldsymbol{\varepsilon}_{j}$ with a new $\boldsymbol{\varepsilon}_{j}^{\prime}$, which results from updating the incoming edges to $\boldsymbol{\varepsilon}_{j}$ in the preceding layer. Thus, during training, when we intervene on $\mathbf{c}_{j}$, we zero out the row $\alpha_{\cdot j}$ while replacing the column $\alpha_{j\cdot}$ with a new column $\beta_{j\cdot}$ that is trained specifically for this interventional setting. The result is $d_{c}+1$ intervention-specific layers: $A_{0}$ for the observational setting, and $A_{1},\ldots,A_{d_{c}}$ for each interventional setting. This is captured as a three-way tensor where each slice of this tensor captures a different interventional context that corresponds to intervening on different concepts. At inference time, to generate interventional samples, we simply swap out the $A_{0}$ for $A_{j}$. Moreover, multi-target interventions can be sampled by zero-ing out multiple rows and substituting multiple $\beta_{j\cdot}$’s and $\boldsymbol{\varepsilon}_{j}$’s.
Multi-dimensional concepts
In the causal model described above, we assumed $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$ for simplicity. Everything carries through if we allow additional flexibility with $\dim(\mathbf{c}_{j})>1$ and $\dim(\boldsymbol{\varepsilon}_{j})>1$, potentially even with $\dim(\mathbf{c}_{j})\neq\dim(\boldsymbol{\varepsilon}_{j})$. In practice, we implement this via two width parameters $0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ and $0pt_{c}=\dim(\mathbf{c}_{j})$. The design choice of enforcing uniform dimensionality is not necessary, but is made here since in our experiments there did not seem to be substantial advantages to choosing nonuniform widths.
With these modifications, $\alpha_{kj}$ becomes a $0pt_{\varepsilon}\times 0pt_{c}$ matrix; instead of substituting rows and columns above, we now substitute slices in the obvious way. The three-way tensor $A$ is now $(d_{c}+1)\times 0pt_{\varepsilon}d_{c}\times 0pt_{c}d_{c}$-dimensional.
Identifiability
An appealing aspect of this architecture is that under certain assumptions, it can be viewed as an approximation to an identifiable model over disentangled concepts. Formally, we assume given observations $\mathbf{x}$, concepts $\mathbf{c}$, and a collection of embeddings $\mathbf{e}$ such that $\mathbf{x}=f(\mathbf{e})$. Additionally assume a noisy version of the linear representation hypothesis, i.e., $\mathbf{c}_{j}=C_{j}\mathbf{e}+\varepsilon_{j}$ with $\varepsilon_{j}\sim\mathcal{N}(0,\Omega_{j})$. Then we have the following identifiability result:
Theorem 4.1 (Identifiability of linear representations).
Assume that the rows of each $C_{j}$ are chosen from a linearly independent set and $f$ is injective and differentiable. Then, given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$.
We do not claim to identify a causal graph, only the latent distribution $p(\mathbf{c})$ over the concepts and its functional structure $C_{j}$ within the data-generating process. As opposed to solving a causal representation learning problem, the causal semantics used in this paper are simply being ported into the generative modeling framework to give solutions to the problem of learning latent distribution structure in a completely unsupervised fashion. Theorem 4.1 shows that the set problems where there exists a structured latent distribution to be identified is nonempty. Moreover, the structure of these identifiable distributions exactly fits our model architecture, meaning that set of problems in which the proposed method will approximate a fully disentangled latent space is nonempty. In other words, there exist data-modeling problems that are solvable by the novel approach introduced in this paper. The technical difficulty in the proof of Theorem 4.1 is analyzing the behaviour of interventions in this model; since our main focus is on methodological aspects and the technical details are fairly involved, these details are deferred to the appendix.
4.3 Summary of architecture
The context module thus proposed offers the following appealing desiderata in practice:
  1. 1.
    There is no coupling between the embedding dimension and the number of concepts, or the complexity of the SEM. In particular, this architecture allows for arbitrary black-box encoder-decoder architectures to be used to learn embeddings, from which a causal model is then trained on top of. This can be trained end-to-end, fine-tuned after pre-training, or even frozen, allowing for substantial computational savings. In our experiments, we illustrate this with both standard convolutional networks as well as NVAE (Vahdat and Kautz, 2020), which is a deep hierarchical VAE with millions of parameters.
  2. 2.
    The intervention layer is a genuine causal model that provides rigourous causal semantics that allow for sampling from arbitrary concept interventions, including interventions that have not been seen during training. But note that we do not claim to learn the structural coefficients, only the reduced model which is sufficient for concept interventions and sampling.
  3. 3.
    The architecture is based on an approximation to an identifiable concept model (Theorem 4.1), which provides formal justification for the intervention layer as well as reproducibility assurances.
  4. 4.
    The context module is itself arbitrarily flexible, meaning that there is no risk of information loss in representing concepts with the embeddings $\mathbf{e}$. Of course, information loss is possible if we compress this layer too much (e.g., by choosing $d_{c}$, $k$, $0pt_{\varepsilon}$, or $0pt_{c}$ too small), but this is a design choice and not a constraint of the architecture itself.
As a consequence, the only tradeoff between representational capacity and causal semantics is design-based: The architecture itself imposes no such constraints. The causal model can be arbitrarily flexible and chosen independent of the black-box encoder-decoder pair, which is also allowed to be arbitrarily flexible.
5 Empirical results
We divide the presentation here into two main parts: small models on (semi-)synthetic datasets (Section 5.1) versus larger models on more challenging datasets (Section 5.2). We emphasize key findings here but include complete results and experimental details in the appendix.
5.1 Small models on (semi-)synthetic data
To systematically investigate the interplay between architectural design, learned representations, and OOD performance, we begin with small-scale experiments with controlled environments. This setting enables rapid iteration, comprehensive ablation studies, and access to ground-truth metrics unavailable in complex real-world data. We evaluate models on two benchmarks:
  • MNIST: We use Morpho-MNIST (Castro et al., 2019) to allow for different interventional contexts, such as making digits thicker or thinner.
  • quad: A novel semi-synthetic visual environment with disentangled latents (colour, shape, size, orientation) and explicit OOD compositionality challenges (e.g., generating unseen combinations of background quadrant colours and central object attributes).
MNIST
On MNIST, we see (Table 1) results of our module across different ablations. The first column (Context module) is our module incorporated into a lightweight convolutional VAE. Ablation 1 is an analogously lightweight standard implementation of $\beta$-VAE trained only on the observational context. Ablation 2 is $\beta$-VAE trained on the full pooled dataset. Ablation 3 is similar to the context module, except that it ignores the contexts and pools everything into a single context (like Ablation 2). These results indicate that the cost of the structured representation learned by the context module is only a slight degradation in the numerical metrics.
Table 1: Ablation on MNIST (200 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
Reconstructed Distribution Evaluation
MMD (obs) ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
Generated Distribution Evaluation
MMD (obs) ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
Figure 2: Example images obtained from the quad dataset. (top) Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention). (bottom) Each subfigure corresponds to a different double-concept intervention context, indicated by the title. Double-concept interventions were not included in the training data, allowing for genuinely OOD evaluation.
Controlled simulations on quad
While evaluations on unsimulated data are the gold standard for evaluating models, simulated environments are valuable for conducting controlled experiments under given conditions. To provide a controllable, synthetic test bed for evaluating composition in a visual environment, we developed a simple semi-synthetic benchmark, visualized in Figure 2.
quad is a visual environment defined by 8 concepts:
  1. 1.
    quad1: The colour of the first quadrant;
  2. 2.
    quad2: The colour of the second quadrant;
  3. 3.
    quad3: The colour of the third quadrant;
  4. 4.
    quad4: The colour of the fourth quadrant;
  5. 5.
    size: The size of the center object;
  6. 6.
    orientation: The orientation (angle) of the center object;
  7. 7.
    object: The colour of the center object;
  8. 8.
    shape: The shape of the center object (circle, square, pill, triangle).
With the exception of shape, which is discrete, the remaining concepts take values in $[0,1]$.
Table 2: Ablation on Quad
Metric / Evaluation Context module Ablation 2
Generated Sample Evaluation
MMD (obs) ↓ 0.0099 (0.0058) 0.0187 (0.0063)
MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
Figure 3 shows OOD samples from a lightweight VAE trained on this dataset. Table 2 shows numerical results comparing our module (incorporated into a lightweight convolutional VAE) to Ablation 2 from above (analogously lightweight $\beta$-VAE trained on the full pooled dataset) for sample generation across the observational and interventional data (same distribution seen during training) along with OOD data. The quad data allows evaluating the sampled OOD distribution (which our module composes from seen contexts) with a ground truth OOD distribution (unseen during training).
Figure 3: Example generated OOD images from a simple, lightweight VAE on the quad benchmark. This shows a proof-of-concept that the concept module successfully composes concepts using even simple architectures. For performant architectures, see the NVAE results in Figure 5.2.
OOD generation vs. reconstruction
There is a crucial difference between OOD reconstruction and OOD generation. By OOD generation, we mean the ability to conditionally generate and combine novel interventions. For example, suppose during training the model only sees small, red objects along with big, blue objects. At test time, we wish to generate OOD samples of big, red objects or small, blue objects.
The crucial difference is that a model may be capable of OOD reconstruction, but not OOD generation. For example, we can always attempt to reconstruct an OOD sample and evaluate its reconstruction error. This is a common metric that has been used in previous work to evaluate the OOD capabilities of generative models (e.g. Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022). But unless the model learns specific latent concepts corresponding to size or colour, then we have no control over the size or colour of random samples from the model. We can always condition on the representations learned by the model, but these will not be disentangled or interpretable.
Thus, OOD generation not only evaluates the ability of a model to compose learned concepts in new ways, but implicitly also the ability of a model to identify and capture underlying concepts of interest. Accordingly, we argue that OOD generation is a more appropriate evaluation metric.
While the numerical results on synthetic data above indicate better performance on OOD generation using our module, a more intuitive visual evaluation on benchmark datasets is provided in Section 5.2, Figure 5.2.
5.2 Large-scale models and benchmarks
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 4: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right). In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together. The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
We now evaluate performance when our module is incorporated into a complex black-box model (NVAE of Vahdat and Kautz (2020)) on more challenging datasets, again including MNIST but now also 3DIdent (Zimmermann et al., 2021) and CelebA (Liu et al., 2015). Since CelebA does not have interventional data, we use this as an important ablation to test how well our module performs when used with conditional data (i.e. by conditioning on attributes in CelebA).
Table 3: Comparison of models (BPD = bits per dimension)
Dataset Type Black-box BPD
MNIST Observational NVAE 0.144
MNIST Interventional NVAE 0.149
3DIdent Observational NVAE 0.673
3DIdent Interventional NVAE 0.754
CelebA Observational NVAE 2.08
CelebA Conditional NVAE 2.13
Table 3 summarizes the results of these training runs, where “observational” means that the standard NVAE (i.e. without a context module) is trained on the entire datasets, whereas “interventional” means that the context module was used. We observe competitive reconstruction, with only slight degradation when the context module is attached to the NVAE decoder. More importantly, Figure 5.2 demonstrates generation capabilities enabled by our module. Each row corresponds to generated samples from the trained model in different contexts:
  1. 1.
    The first row shows generated observational samples;
  2. 2.
    The second and third row contain generated single-node interventions;
  3. 3.
    The final row shows generated samples where two distinct concepts are composed together.
For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon. CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data. Although there is a slight drop in OOD perceptual accuracy on CelebA, this is to be expected since the model was trained on conditional data as opposed to interventions. Finally, observe the last row of Figure 5.2, which demonstrates the OOD generation capability enabled by our module: these contexts did not appear in the training data but our module is able to sensibly generate them by composing and generalizing from the concepts it learned according to different data contexts.
Acknowledgements
The work by AM was supported by Novo Nordisk Foundation Grant NNF20OC0062897.
This research was supported in part through the computational resources and staff contributions provided for the Mercury high performance computing cluster at The University of Chicago Booth School of Business which is supported by the Office of the Dean.
Appendix A Identifiability of the proposed model
In this appendix we set up and prove Theorem 4.1.
A.1 Set-up
Our setup is the following: Let $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n}),\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$ and $\mathbf{e}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{d_{e}})$ be random vectors taking values in $\mathcal{X},\mathcal{C}$ and $\mathcal{E}$, respectively. For simplicity, we assume that $\mathcal{X}\subseteq\mathbb{R}^{n}$, $\mathbf{e}\in\mathcal{E}\subseteq\mathbb{R}^{d_{e}}$, and $\mathbf{x}=f(\mathbf{e})$ for some embeddings $\mathbf{e}\in\mathcal{E}$. We assume that $f$ is injective and differentiable. This map is allowed to be arbitrarily non-linear: We make no additional assumptions on $f$.
The random variables $\mathbf{c}_{j}$ will be referred to as “concepts”: Intuitively they represent abstract concepts such as shape, size, colour, etc. that may not be perfectly represented by any particular embedding in $\mathbf{e}$. The linear representation hypothesis (Section 3) is an empirical observation that abstract concepts can be approximately represented as linear projections of a sufficiently flexible latent space. This is operationalized by assuming that $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$ (see (A1) below). The matrix $C_{j}$ is referred to as the representation of the concept $\mathbf{c}_{j}$. If $\mathbf{c}_{j}\in\mathbb{R}$ (i.e. $C_{j}\in\mathbb{R}^{1\times d_{e}}$), then $\mathbf{c}_{j}$ is referred to as an atom and $C_{j}$ its atomic representation. We will often denote atoms by $\mathbf{a}_{j}$ for notational clarity.
We consider $(\mathbf{x},\mathbf{c},\mathbf{e})$ satisfying the following:
  1. (A1)
    $\mathbf{c}_{j}=C_{j}\mathbf{e}+\eta_{j}$ for all $j\in[d_{c}]$ for some real matrices $C_{1}\ldots,C_{d_{c}}$, where $\eta_{j}\sim\mathcal{N}(0,\Lambda_{j})$ for some diagonal matrix $\Lambda_{j}\succ 0$.
  2. (A2)
    $\mathbf{x}=f(\mathbf{e})$ for some injective and differentiable function $f$.
  3. (A3)
    There is a DAG $\mathsf{G}=([{d_{c}}],E)$ such that
    $\mathbf{c}_{j}=\sum_{k\in\operatorname{pa}_{\mathsf{G}}(j)}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{j},\quad\text{ for all $j\in[{d_{c}}]$}$ (25)
    and where $\boldsymbol{\varepsilon}_{1},\ldots,\boldsymbol{\varepsilon}_{d_{c}}$ are mutually independent and $\mathbf{e},\eta$ are independent of $\boldsymbol{\varepsilon}=(\boldsymbol{\varepsilon}_{1},\ldots,\boldsymbol{\varepsilon}_{d_{c}})$.
Note that these assumptions are not necessary in practice—as evidenced by our experiments in Section 5 which clearly violate these assumptions—and are merely used here to provide a proof-of-concept identifiability result based on our architectural choices.
Our goal is to identify the concept marginal $p(\mathbf{c})$ and concept representations $(C_{1},\ldots,C_{d_{c}})$ from concept interventions. Concept interventions are well-defined through the structural causal model defined by the DAG $\mathsf{G}$ and its structural equations. We consider single-node concept interventions where $\mathbf{c}_{j}$ is stochastically set to be centered at $\boldsymbol{\mu}_{j}$ with variance $\Omega_{j}\succ 0$ (making this precise requires some set-up; see Appendix A.3.1 for details).
By identifiability, we mean the usual notion of identifiability up to permutation, shift, and scaling from the literature. Namely, for any two sets of parameters $(f,C_{1},\ldots,C_{d_{c}})$ and $(\widetilde{f},\widetilde{C}_{1},\ldots,\widetilde{C}_{d_{c}})$ that generate the same observed marginal $p(\mathbf{x})$, there exists permutations $P_{j}$, diagonal scaling matrices $D_{j}$, and a shift $b\in\mathbb{R}^{d_{e}}$, such that
$C_{j}f^{-1}(\mathbf{x})=D_{j}P_{j}\widetilde{C}_{j}(\widetilde{f}^{-1}(\mathbf{x})+b),\quad\text{for all $j$ and $\mathbf{x}$.}$ (26)
Moreover, there exists an invertible linear transformation $L$ such that the concepts satisfy
$C_{j}=P_{j}\widetilde{C}_{j}L^{-1}.$ (27)
This definition of identifiability, which aligns with similar notions of identifiability that have appeared previously (e.g. Squires et al., 2023; von Kügelgen et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025), implies in particular that the concepts $\mathbf{c}_{j}$ are identifiable up to permutation, shift, and scale, and that their representations are identifiable up to a linear transformation. As a result, the concept marginal $p(\mathbf{c})$ is also identifiable (again up to permutation, shift, and scale as well). The shift ambiguity can of course be resolved by assuming the concepts are centered (zero-mean), and the scale ambiguity can be resolved by assuming concepts are normalized (e.g. unit norm).
We make the following assumptions, which are adapted from Rajendran et al. (2024) to the present setting. Note that Rajendran et al. (2024) do not consider a structural causal model over $\mathbf{c}$, and thus the notion of a concept intervention there is not well-defined.
Assumption 1 (Atomic representations).
There exists a set of atomic representations $\mathcal{A}=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{n}\}$ of linearly independent vectors $\mathbf{a}_{j}\in\mathbb{R}^{d_{e}}$ such that the rows of each representation $C_{j}$ are in $\mathcal{A}$. We denote the indices of $\mathcal{A}$ that appear as rows of $C_{j}$ by $S^{j}$, i.e. $S^{j}=\{i:\mathbf{a}_{i}\text{ is a row in $C_{j}$}\}$. We assume that $\cup_{j}S^{j}=\mathcal{A}$, i.e. every atom in $\mathcal{A}$ appears in some concept representation $C_{j}$.
Define a matrix $M\in\mathbb{R}^{n\times d_{c}}$ to track which concepts use which atoms, i.e.
$M_{ij}=\begin{cases}\frac{1}{\omega_{j}^{2}}&\text{if $i\in S^{j}$}\\ 0&\text{otherwise,}\end{cases}$ (28)
where $\omega_{j}^{2}$ are the diagonal entries of $\Omega_{j}$. Similarly, we define a matrix $B\in\mathbb{R}^{n\times d_{c}}$ given by
$B_{ij}=\begin{cases}\frac{(\boldsymbol{\mu}_{j})_{k}}{\omega_{j}^{2}}&\text{if $i\in S^{j}$ and the $k$th row of $C_{j}$ is $\mathbf{a}_{i}$,}\\ 0&\text{otherwise.}\end{cases}$ (29)
The second assumption ensures that the concepts and associated interventional environments are sufficiently diverse.
Assumption 2 (Concept diversity).
For every pair of atoms $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$ with $i\neq j$ there is a concept $\mathbf{c}_{k}$ such that $i\in S^{k}$ and $j\notin S^{k}$. Furthermore, $\operatorname{rank}(M)=n$ and there exists $v\in\mathbb{R}^{d_{c}}$ such that $v^{\top}M=0$ and $(v^{\top}B)_{i}\neq 0$ for each coordinate $i$.
A.2 Formal statement of Theorem 4.1
With this setup, we can now state and prove a more formal version of Theorem 4.1.
Theorem A.1.
Under Assumptions 1-2 and given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$, in the sense of (26-27).
In the context of learning disentangled representations, identifying the concept marginal $p(\mathbf{c})$ is particularly relevant, since this implies we also learn certain (in)dependence relationships between the concepts. In other words, if the “true” concept representations are disentangled (i.e. independent), then we identify disentangled concepts.
In relation to existing work, we make the following remarks:
  • While our proof ultimately relies on techniques from Rajendran et al. (2024), Theorem A.1 does not trivially follow from their results, which do not cover interventions over the concept marginal $p(\mathbf{c})$. Extending their results to the case of concept interventions, including making sure these interventions are well-defined causal objects, is the main technical difficulty in our proof. This is surprisingly tricky since we do not assume a full structural causal model over $(\mathbf{x},\mathbf{c},\mathbf{e})$: Interventions are only well-defined over $\mathbf{c}$, which makes deducing the downstream effects on the observed $\mathbf{x}$ somewhat subtle.
  • Various results in CRL (e.g. Buchholz et al., 2023; von Kügelgen et al., 2023; Varici et al., 2024, see Section 3 for more references) prove identifiability results under interventions, however, these results require interventions directly on the embeddings $\mathbf{e}$, as opposed to the concepts $\mathbf{c}$. Theorem A.1, by contrast, requires only $d_{c}\ll d_{e}$ total interventions and in doing so directly address the technical challenges associated with intervening directly on concepts as opposed to embeddings.
  • Leemann et al. (2023); Fokkema et al. (2025) study concept identifiability when the nonlinear function $f$ is known. We avoid this simplifying assumption (sometimes called post-hoc identifiability), and dealing with the potential nonidentifiability of $f$ is another technical complication our setting must deal with. Indeed, in our setting we can only identify the behaviour of $f$ on the concept subspaces defined by $C_{j}$, which is much weaker than identifying all of $f$ on $\mathcal{E}$. On the other hand, we have left finite-sample aspects to future work.
A.3 Proof of Theorem 4.1
To prove Theorem 4.1, we begin with some notation. Let $\mathcal{M}$ denote the collection of all distributions $P$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ satisfying (A1)-(A3) for some choice of $C_{1},\ldots,C_{d_{c}}$, $\eta$, $\mathsf{G}$, $f_{1},\ldots,f_{d_{c}}$, and $\boldsymbol{\varepsilon}$. For a fixed $\mathsf{G}$, let
[…]
The model $\mathcal{M}(\mathsf{G})$ is nonempty for any $\mathsf{G}$, as can be seen by letting $\mathbf{e},\eta$ and $\boldsymbol{\varepsilon}$ be normally distributed and taking all functions to be linear. Moreover, $\mathcal{M}=\cup_{\mathsf{G}}\mathcal{M}(\mathsf{G})$. Finally, let $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ denote the collection of all marginal distributions over $\mathbf{c}$ induced by the distributions in $\mathcal{M}(\mathsf{G})$.
A.3.1 Concept interventions
In practice, the edge structure of the graph $\mathsf{G}$ is unknown, and we do not assume this is known in advance. We work in the setting where we can only observe outcomes of the variables $\mathbf{x}$. However, we allow for the possibility to intervene on the variables $\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}}$. We consider consider only interventional distributions with single-node targets $I=\{j\}$ for $j\in[m]$. Note that, by definition of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, if $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ then there is a distribution $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ such that $\widetilde{p}(\mathbf{c})=p(\mathbf{c})$ for all $\mathbf{c}\in\mathcal{C}$. Although immediate from these definitions, we record this as a lemma for future use:
Lemma A.2.
If $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ then there is a distribution $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ such that $\widetilde{p}(\mathbf{c})=p(\mathbf{c})$ for all $\mathbf{c}\in\mathcal{C}$.
For the target $I=\{j\}$ and distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ we say that $P^{j}$ is an interventional distribution for $j$ and $P$ if in $P^{j}$ we have the following identities:
$\begin{split}\mathbf{c}_{i}&=\sum_{k\in\operatorname{pa}_{\mathsf{G}}(i)}\alpha_{ki}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{i},\\ \mathbf{c}_{j}&=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime},\quad\text{ where }\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2}).\end{split}$ (31)
That is, the structural equations defining $\mathbf{c}_{i}$ for $i\neq j$ are equal to those defining $\mathbf{c}_{i}$ in the observational distribution $P$, while the distribution of $\mathbf{c}_{j}$ is manipulated to be $\mathcal{N}(\boldsymbol{\mu}_{j},\omega_{j}^{2})$, independent of the other concepts.
The following is obvious from our assumptions:
Lemma A.3.
For any distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, the interventional distribution $P^{j}$ defined by (31) is also an element of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, i.e. $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
Thus, we have the usual, well-defined notion of intervention over the concepts $\mathbf{c}$. The next step is to translate the effect of these interventions onto $\mathbf{e}$ and $\mathbf{x}$.
The following proposition shows that an interventional distribution $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ always extends to a well-defined post-interventional joint distribution over $(\mathbf{x},\mathbf{c},\mathbf{e})$ that can be interpreted as an interventional distribution for the appropriate choice of joint observational distribution over $(\mathbf{x},\mathbf{c},\mathbf{e})$.
Proposition A.4.
Let $P^{j}$ be an interventional distribution for $j$ and $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Then there exists a distribution $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
  1. 1.
    $\widetilde{p}^{j}(\mathbf{c})=p^{j}(\mathbf{c})$ for all $\mathbf{c}\in\mathcal{C}$,
  2. 2.
    $\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})$ for all $\mathbf{e}\in\mathcal{E}$,
  3. 3.
    If $\widetilde{P}$ is induced by the functional relation $\mathbf{x}=f(\mathbf{e})$, and similarly $\widetilde{P}^{j}$ with $\mathbf{x}=f^{j}(\mathbf{e})$ then $f^{j}=f$.
Proof.
Lemma A.3 implies that $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Thus Lemma A.2 implies the existence of joint distributions $\widetilde{P},\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
$\widetilde{p}(\mathbf{c})=p(\mathbf{c})\quad\text{ and }\quad\widetilde{p}^{j}(\mathbf{c})=p^{j}(\mathbf{c})\text{ for all }\mathbf{c}\in\mathcal{C}.$ (32)
(1) follows from the definition of $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ for $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ from Lemma A.2 above; i.e. (32) above.
(2) follows from the definition of an interventional distribution and the factorization
$\widetilde{p}(\mathbf{x},\mathbf{c},\mathbf{e})=\widetilde{p}(\mathbf{x}\mid\mathbf{e})\widetilde{p}(\mathbf{c}\mid\mathbf{e})\widetilde{p}(\mathbf{e}),$
satisfied by any distribution $P\in\mathcal{M}(\mathsf{G})$. Namely, the interventional distribution $P^{j}$ satisfies
$\begin{split}\widetilde{p}^{j}(\mathbf{x},\mathbf{c},\mathbf{e})&=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ &=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e}){p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ \end{split}$
according to the definition of $\widetilde{P}$. Since the intervention only perturbs the conditional factors $\widetilde{p}(\mathbf{c}_{j}\mid\mathbf{c}_{\textrm{pa}_{\mathsf{G}}(j)},\mathbf{e})$ in $\widetilde{p}(\mathbf{c}\mid\mathbf{e})$ for $j$, we can always choose $\widetilde{P}$ and $\widetilde{P}^{j}$ such that $\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})$.
(3) follows similarly to (2). Namely, since the intervention only perturbs the conditional factors specified above, we can always choose $f^{j}=f$. ∎
We additionally have the following:
Proposition A.5.
Let $I_{j}:=\{j\}$ be the collection of single-node intervention targets, and consider the interventional distributions $P^{1},\ldots,P^{d_{c}}$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Then there exists $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}\in\mathcal{M}(\mathsf{G})$ such that
$\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})\qquad\textrm{ for all $j\in[d_{c}]$}$
and $f^{j}=f$ for all $j$.
Proof.
Fix a choice of $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ for $P$, which specifies a function $f$. For the chosen $\widetilde{P}$, we construct $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}$ according to Proposition A.4. The result follows by applying Proposition A.4 for every $j$. ∎
Proposition A.5 will allow us (below) to define more formally the environments that are used to identify concepts. Intuitively, each environment corresponds to single-node interventions on a single concept $\mathbf{c}_{j}$, however, we can only observe the effect this intervention has on the observed $\mathbf{x}$. The purpose of Propositions A.4-A.5 are to trace the dependence between $\mathbf{x}$ and $\mathbf{c}$—via the embeddings $\mathbf{e}$—according to the model (A1)-(A3).
Formally, let $I_{j}=\{j\}$ with $\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2})$ and $\mathbf{c}_{j}=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime}$ according to (31), i.e. each environment corresponds to a single-node intervention on the $j$th concept $\mathbf{c}_{j}$. By Proposition A.5, there exist post-intervention joint distributions $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ such that
$\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})\qquad\textrm{ for all $j\in[d_{c}]$}$
and $f^{j}=f$ for all $j$. Then the $j$th environment corresponds to sampling $\mathbf{x}=f(\mathbf{e})$ where $\mathbf{e}\sim p_{j}(\mathbf{e}):=\widetilde{p}^{j}(\mathbf{e}\,|\,\mathbf{c}_{j}=\boldsymbol{\mu}_{j})$.
A.3.2 Reduction to concept identification
We begin by computing the log-likelihood ratio between the observational and interventional environments. The idea is that the concept representations are revealed through fluctuations in this likelihood ratio. Using Proposition A.5, we have
$\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\log p_{0}(\mathbf{e})-\log\widetilde{p}^{j}(\mathbf{c}_{j}=\boldsymbol{\mu}_{j}\,|\,\mathbf{e})-\log\widetilde{p}^{j}(\mathbf{e})$ (33)
$\propto-\tfrac{1}{2}(\boldsymbol{\mu}_{j}-C_{j}\mathbf{e})^{T}\Omega_{j}^{-1}(\boldsymbol{\mu}_{j}-C_{j}\mathbf{e})$ (34)
$\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}(\mathbf{a}_{i}^{T}\mathbf{e})^{2}-B_{ij}\mathbf{a}_{i}^{T}\mathbf{e}+O(1).$ (35)
From here, we proceed as in Rajendran et al. (2024) (note that in their notation, $\mathbf{e}=\mathbf{z}$). For completeness, we sketch the main steps here. The basic idea is to analyze the system of $d_{c}$ equations induced by the log-likelihood ratios for each concept intervention. First, we reduce the system to a standard form by transforming the atoms into the standard basis via a change of coordinates. After this change of coordinates, the system can be rewritten as
$h_{j}(\mathbf{e}):=\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}\mathbf{e}_{i}^{2}-B_{ij}\mathbf{e}_{i}+O(1).$ (36)
These functions are convex, which allows us to identify the $|S_{T}|$ where $S_{T}=\cup_{j\in T}S_{j}$ by minimizing the convex function $\sum_{j\in T}h_{j}(\mathbf{e})$. Then a similar induction argument identifies the matrices $M$ and $B$. Now, given two distinct representations of $p(\mathbf{x})$ via nonlinear mixing functions $f$ and $\widetilde{f}$, define $\varphi=\widetilde{f}^{-1}\circ f$. Let $\mathbf{h}=(h_{1},\ldots,h_{n})$ and write $\mathbf{e}^{c}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{n})$ and $\mathbf{e}=(\mathbf{e}^{c},\mathbf{e}^{\perp})\in\mathbb{R}^{n\times(d_{c}-n)}$. Let $\iota^{\perp}(\mathbf{e}^{c})=(\mathbf{e}^{c},0)$, $\pi^{c}(\mathbf{e})=\mathbf{e}^{c}$, and $\varphi^{\perp}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\varphi^{\perp}(\mathbf{e}^{c})_{i}=\varphi(\mathbf{e},0)_{i}$. Note that via the change of variables formula, we can write $\mathbf{h}(\mathbf{e})=\mathbf{H}(\mathbf{x})$ for some function $\mathbf{H}$. Then
$\mathbf{h}(\mathbf{e}^{c},0)=\mathbf{H}(f(\mathbf{e}^{c},0))=\mathbf{H}(\widetilde{g}(\varphi^{\perp}(\mathbf{e}^{c})))=\mathbf{h}(\varphi^{\perp}(\mathbf{e}^{c})),$ (37)
which is the crucial relation used in the proof of Theorem 1 of Rajendran et al. (2024). From here, identifiability follows from a straightforward but lengthy linear algebraic argument based on the first-order conditions for minimizing $\mathbf{h}$ and the identifiability of $M,B$ proved above. The proof of Theorem A.1 follows.
Appendix B quad: A semi-synthetic benchmark for OOD composition
To provide a controllable, synthetic test bed for evaluating composition in a visual environment, we developed a simple semi-synthetic benchmark, visualized in Figures 56. quad is a visual environment defined by 8 concepts:
  1. 1.
    quad1: The colour of the first quadrant;
  2. 2.
    quad2: The colour of the second quadrant;
  3. 3.
    quad3: The colour of the third quadrant;
  4. 4.
    quad4: The colour of the fourth quadrant;
  5. 5.
    size: The size of the center object;
  6. 6.
    orientation: The orientation (angle) of the center object;
  7. 7.
    object: The colour of the center object;
  8. 8.
    shape: The shape of the center object (circle, square, pill, triangle).
With the exception of shape, which is discrete, the remaining concepts take values in $[0,1]$. This environment has the following appealing properties:
  • Composing multiple concepts is straightforward and perceptually distinct;
  • Concepts are continuous, which allows a variety of soft and hard interventions;
  • Sampling is fast and easy, and requires only a few lines of Python code and no external software;
  • Creating OOD hold-out validation datasets of any size is straightforward;
  • This can be simulated on any $n\times n$ grid (as long as $n$ is large enough to distinguish different shapes; $n\geq 16$ is large enough in practice). In our experiments, we used $n=64$.
These properties make quad especially useful for both quantitative and qualitative evaluation; indeed, it was created precisely for this reason. Example images are shown in Figure 5.
To describe our experimental setting, let $\mathbf{c}_{j}$ indicate the $j$th concept as listed above. We used seven different contexts, including an observational context and one context each for single-concept interventions on $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$, i.e. quad1, quad2, quad3, quad4, size, and orientation. We use a sample size of 50 000 images per context. In each context, $\mathbf{c}_{7}$ (object) and $\mathbf{c}_{8}$ (shape) are sampled uniformly at random. The remaining six concepts were sampled as follows:
  • The observational context was generated by randomly sampling $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$ independently from the range $[0,0.5]$,
  • The interventional contexts were generated by isolating a particular concept and sampling it uniformly from $[0.5,1]$, while sampling the rest from $[0,0.5]$.
For example, colours in $[0,0.5]$ include greens, yellows, and reds, and colours in $[0.5,1]$ include blues, purples, and magenta. Thus, the object colour takes on any value in every context, whereas the four quadrant colours are restricted in the observational context. When we intervene on a particular quadrant, we see that the colour palette noticeably shifts from greens to blues (Figure 5).
Similarly, we can compose multiple interventions together, as in Figure 6. This facilitates evaluation of the concept composition capabilities of models trained on single interventions, providing a ground truth distribution to compare against generated OOD samples like in Figure 7.
Figure 5: Example images obtained from the quad dataset. Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention).
We also made use of the easy-to-generate and perceptually distinct concept compositions in this dateset by evaluating models on held-out contexts (each containing 10 000 images) with double-concept interventions, including quad2_quad3, quad2_quad4, quad2_size, quad1_quad2, quad1_quad3, quad1_quad4, and quad1_size. Examples can be seen in Figure 6.
For comparison, see Figure D.1.2 for examples of generated multi-concept compositions (OOD) from a simple 3-layer convolutional model. The training data did not contain any multi-concept compositions.
Figure 6: Example images obtained from the quad dataset. Each subfigure corresponds to a different double-concept intervention context, indicated by the title.
Appendix C Experimental details
In this appendix we provide additional details on our experimental protocol and set-up.
C.1 Evaluation metrics
We evaluated each model on the following metrics:
  • Validation ELBO loss: This is the standard ELBO loss used for both training and validating VAEs. We report this in bits per dimension (bpd), which is the loss scaled by a factor of $\frac{1}{\mathrm{num\_pixels}\cdot\ln 2}$.
  • Reconstruction loss: This approximates the negative log-likelihood. We use binary cross-entropy and report bpd.
  • Covariance error: Unlike the losses above, which evaluate individual reconstructed images with respect to their corresponding ground truth images, this metric (and the ones below) evaluates a reconstructed distribution with respect to a ground truth distribution. Specifically, denoting the empirical covariance matrix across the ground truth datasets as $S^{T}$ and that over the reconstructed dataset as $S^{R}$, we report $\frac{1}{d^{2}}\sum_{i,j=1}^{d}|S^{T}_{ij}-S^{R}_{ij}|$.
  • Mean error: This is analogous to the previous metric but using sample means instead of covariances:$\frac{1}{d}\sum_{i=1}^{d}|\mu^{T}_{i}-\mu^{R}_{i}|$.
  • Maximum mean discrepancy (mmd): Another metric comparing distributions, MMD (Gretton et al., 2012) is a kernel-based metric for comparing two general distributions that may not be distinguishable by merely mean and covariance.
When reporting values for these metrics in tables, we show the mean, averaged over different runs/seeds, followed in parentheses by the standard error. We generally use a 70/30 training/validation split. When appropriate, evaluations are split up so that metrics are computed with respect to the observational context (obs), or averaged over single-concept interventional contexts (ivn), or averaged over double-concept interventional contexts (ood)—with the exception of Table 3, which is detailed in Section 5.2 and Appendix C.4.
C.2 Datasets
In addition to the quad benchmark introduced in Appendix B, we used the following benchmark datasets.
MNIST
We use a variation of the classic MNIST dataset (LeCun et al., 1998) known as Morpho-MNIST (Castro et al., 2019) along with additional affine transformations of MNIST (e.g., as in Simard et al. (2003)). In addition to the observational context (standard MNIST images), the dataset contains the five contexts corresponding to single-concept interventions on concepts scaled, shear, swel, thic, and thin. The datasets contain 60 000 images per context.
3DIdent
The 3DIdent dataset (Zimmermann et al., 2021) consists of a 3D object, rendered using the Blender rendering engine, under different lighting conditions and orientations. We generated 50 000 samples per context corresponding to single-node interventions on background colour, spotlight colour, and object colour, in addition to the observational context.
CelebA
We use the original CelebA data (Liu et al., 2015), using the attributes included in the original dataset to define different contexts. Since we do not have access to interventions on this dataset, we relied on conditioning to create different contexts for different concepts. To minimize leakage between contexts, we chose 6 attributes by minimizing the between-attribute correlations and also choosing attributes that are easily detected visually (e.g. black or blond hair). This resulted in six total contexts of varying sizes, one observational along with single-node interventions on bags_under_eyes (41 446 samples), black_hair (48 472 samples), blond_hair (29 983 samples), male (84 434 samples), and mouth_slightly_open (97 942 samples).
C.3 Architectures
We tested our context module using two black-box models:
  1. 1.
    lightweight-VAE: A standard 3-layer convolutional network with latent dimension $\dim(\mathbf{z})=128$. This is used for the MNIST and quad experiments in Section 5.1.
  2. 2.
    NVAE: A deep hierarchical VAE; see Vahdat and Kautz (2020) for details. This is used for the 3DIdent and CelebA experiments, as well as some MNIST experiments, in Section 5.2.
The lightweight-VAE was used for comprehensive ablations and experiments on simple datasets (150 total models overall, 5 per column per table in Appendix D.1). Since NVAE requires substantially more resources and time to train, this was used to train 6 total models, 2 per dataset.
To provide an apples-to-apples baseline, we trained MNIST using both lightweight-VAE and NVAE.
C.4 Discussion of experimental results (Section 5 in main paper)
Table 1 and Table 2 are interpreted directly as described in Appendix C.1. However, Table 3 warrants some clarification.
Figures D.2.1-D.2.4 in Appendix D.2, show the results of NVAE+context-module across three different datasets (MNIST, 3DIdent, and CelebA). All images here are generated by the model: that is, we draw a sample from the prior and pass it through the context module and decoder—this is to be contrasted with reconstruction, in which a reference image is passed through the encoder to obtain a sample from the posterior that is then passed through the (context module and) decoder. The top row is the result of asking the model to generate images from the observational context (in which no concepts are intervened or conditioned on), while the second and third rows contain images generated from respective contexts in which one concept has been intervened on (in the case of MNIST and 3DIdent) or conditioned on (in CelebA). These first three rows correspond to contexts the model has seen during training (though this is a generation and not reconstruction task). In contrast, the bottom row contains images resulting from asking the model to compose the concepts it has learned (specifically, the two concepts that were individually intervened/conditioned on in the second and third rows)—importantly, notice that the model has not seen data from this double-concept intervention (in the case of MNIST and 3DIdent).
Table 3 compares two different models (namely, NVAE versus NVAE+context-module) across the three datasets. The different models are indicated by the Type column, with “Observational” denoting the plain NVAE model trained on a dataset in which all contexts are pooled, while “Interventional” and “Conditional” denote NVAE+context-module trained on unpooled datasets containing different contexts. The BPD column reports the reconstruction loss (described in Appendix C.1) in standard units so that it is directly comparable with the results of Vahdat and Kautz (2020) and other papers.
C.5 Additional ablations
We conducted additional ablations to stress test our model and isolate the effect of the context module. This included additional experiments on different models, regularization, and expressivity.
Models
On MNIST and quad, we ran ablation studies on lightweight-VAE+context-module (denoted “Context module” in tables)—this is our module attached to the black-box, making full use of the different contexts. To evaluate the proposed module, we compared its performance to three natural baselines:
  1. 1.
    Ablation 1: lightweight-VAE (without the context module) trained only on the observational context—this is the black-box when denied all (non-observational) context information and data.
  2. 2.
    Ablation 2: lightweight-VAE (without the context module) trained on the full pooled dataset—this is the black-box when given all data but no explicit context information.
  3. 3.
    Ablation 3 is lightweight-VAE+context-module trained on the full pooled dataset—this is our module attached to the black-box given all the data but not making use of context-specific information.
Note that due to differences in the size of the pooled vs. observational-only datasets, loss values from Ablation 1 are not directly comparable to Ablations 2 or 3 (or the context module), but are included anyway for reference. Results are shown in Section 5.1 and Appendix D.1.1.
Regularization
We run lightweight-VAE+context-module on MNIST, independently trying out two different regularizers, varying the regularization weight $\lambda$ for each: group lasso and $\ell_{2}$. These are applied to the weights of the reduced-form SEM $A_{0}$ in Equation 1. We also try varying $\beta$ (as in $\beta$-VAE). Results are shown in Appendix D.1.3.
Expressivity
We run lightweight-VAE+context-module on MNIST, jointly varying three parameters that control expressiveness of our context module: expressive input width $0pt_{\mathrm{exp}}$ (which controls latent dimension $\dim(\mathbf{z})$), depth of the expressive layer $h_{\mathrm{exp}}$, and concept width ($0pt_{c}=\dim(\mathbf{c}_{j})=0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ as in Section 4.2). Results are shown in Appendix D.1.4.
Appendix D Additional results
In this section, we present all additional results, including experimental settings described in Appendix C for our context module with the lightweight-VAE black-box architecture as well as additional results on MNIST, 3DIdent, and CelebA for our context module with NVAE.
D.1 lightweight-VAE
D.1.1 Ablations (MNIST and quad)
A description of the architecture and experiment is given in Appendix C. Table 4 shows ablation results after 200 epochs of training on MNIST, and Table 5 is after 1000 epochs. [^1] All other experiments in Appendix D.1 are accordingly run with no ablation and on 200 epochs. These are comparable, with Table 5 indicating some overfitting. Table 6 shows ablation results on quad after 200 epochs.
Altogether, these ablations studies suggest that the cost of the structured representation learned by the context module is only a slight degradation in some of the numerical metrics, while Figure 7 shows sensible OOD concept composition using our module.
Table 4: Ablation on MNIST (200 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.4243 (0.1312) 0.6468 (0.0135) 1.1859 (0.0576) 1.3886 (0.1374)
Mean Error ↓ 0.0218 (0.0033) 0.0052 (0.0005) 0.0204 (0.0018) 0.0237 (0.0053)
MMD ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
Covariance Error (ivn) ↓ 2.3717 (0.2764) 2.1983 (0.2567) 2.0005 (0.2380) 2.4591 (0.2833)
Mean Error (ivn) ↓ 0.0332 (0.0030) 0.0221 (0.0019) 0.0258 (0.0030) 0.0339 (0.0039)
MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
Generated Distribution Evaluation
MMD ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
Table 5: Ablation on MNIST (1000 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.3150 (0.0085) 0.1889 (0.0020) 0.2393 (0.0087) 0.3145 (0.0116)
Reconstruction Loss (bpd) ↓ 0.2836 (0.0087) 0.1513 (0.0025) 0.2010 (0.0090) 0.2837 (0.0112)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.3057 (0.0635) 0.5542 (0.0061) 1.1227 (0.0940) 1.5646 (0.2203)
Mean Error ↓ 0.0154 (0.0023) 0.0040 (0.0001) 0.0194 (0.0038) 0.0274 (0.0096)
MMD ↓ 0.0089 (0.0010) 0.0000 (0.0000) 0.0022 (0.0014) 0.0172 (0.0040)
Covariance Error (ivn) ↓ 2.2713 (0.2320) 2.2229 (0.2536) 1.9782 (0.2247) 2.6567 (0.2901)
Mean Error (ivn) ↓ 0.0314 (0.0023) 0.0221 (0.0018) 0.0263 (0.0030) 0.0392 (0.0046)
MMD (ivn) ↓ 0.0932 (0.0340) 0.0298 (0.0108) 0.0388 (0.0132) 0.0610 (0.0177)
Generated Distribution Evaluation
MMD ↓ 0.0045 (0.0004) 0.0028 (0.0004) 0.0055 (0.0011) 0.0179 (0.0129)
MMD (ivn) ↓ 0.0211 (0.0058) 0.0156 (0.0037) 0.0159 (0.0036) 0.0231 (0.0034)
Table 6: Ablation on quad
Metric / Evaluation Context module Ablation 2
Validation Metrics
Validation ELBO (bpd) ↓ 5.0031 (0.7180) 3.9357 (0.8504)
Reconstruction Loss (bpd) ↓ 4.9586 (0.7161) 3.8795 (0.8506)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 14.5117 (0.5123) 6.0099 (1.3163)
MAE_Mean ↓ 0.0176 (0.0009) 0.0058 (0.0015)
MMD ↓ 0.0000 (0.0000) 0.0000 (0.0000)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 25.3612 (6.8679) 15.8097 (2.4898)
MAE_Mean ↓ 0.0295 (0.0055) 0.0191 (0.0036)
MMD ↓ 0.0060 (0.0030) 0.0022 (0.0015)
Out-of-Distribution (OOD) Evaluation (Averaged)
MAE_Cov ↓ 43.5994 (2.8442) 36.3404 (3.0105)
MAE_Mean ↓ 0.0570 (0.0036) 0.0440 (0.0037)
MMD ↓ 0.0220 (0.0031) 0.0103 (0.0018)
Generated Distribution Evaluation
MMD ↓ 0.0099 (0.0058) 0.0187 (0.0063)
MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
D.1.2 Examples of concept composition (quad)
Figure 7: Example generated OOD images from a run of Quad.
D.1.3 Regularization (MNIST)
A description of the architecture and experiment is given in Appendix C. Tables 7 and 8 respectively show results of group lasso (Yuan and Lin, 2006) and $\ell_{2}$ regularization, while Table 9 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017). [^2] All other experiments in Appendix D.1 are run with $\lambda=0$ for group lasso and $\ell_{2}$ and with $\beta=1$. Altogether, these results indicate relatively stable performance across regularization weights while suggesting that tuning these regularizers could be worthwhile (e.g., notice that $\lambda=10$ improves performance).
Table 7: Group lasso regularization on MNIST
Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
Validation Metrics
Validation ELBO (bpd) ↓ 0.2983 (0.0053) 0.2907 (0.0081) 0.2869 (0.0064)
Reconstruction Loss (bpd) ↓ 0.2683 (0.0047) 0.2605 (0.0088) 0.2568 (0.0074)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.2976 (0.0608) 1.2175 (0.0611) 1.3087 (0.0469)
Mean Error ↓ 0.0196 (0.0020) 0.0134 (0.0018) 0.0178 (0.0023)
MMD ↓ 0.0125 (0.0018) 0.0069 (0.0023) 0.0109 (0.0015)
Covariance Error (ivn) ↓ 2.2654 (0.2437) 2.2651 (0.2638) 2.2818 (0.2705)
Mean Error (ivn) ↓ 0.0277 (0.0018) 0.0277 (0.0025) 0.0315 (0.0031)
MMD (ivn) ↓ 0.0993 (0.0367) 0.0982 (0.0356) 0.0986 (0.0358)
Generated Distribution Evaluation
MMD ↓ 0.0048 (0.0005) 0.0053 (0.0005) 0.0057 (0.0009)
MMD (ivn) ↓ 0.0283 (0.0101) 0.0274 (0.0076) 0.0211 (0.0061)
Table 8: $\ell_{2}$ regularization on MNIST
Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
Validation Metrics
Validation ELBO (bpd) ↓ 0.3079 (0.0138) 0.2917 (0.0063) 0.2966 (0.0084)
Reconstruction Loss (bpd) ↓ 0.2774 (0.0145) 0.2612 (0.0067) 0.2667 (0.0092)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.3562 (0.1205) 1.2883 (0.0103) 1.4347 (0.0955)
Mean Error ↓ 0.0178 (0.0021) 0.0182 (0.0018) 0.0209 (0.0023)
MMD ↓ 0.0090 (0.0020) 0.0088 (0.0022) 0.0119 (0.0025)
Covariance Error (ivn) ↓ 2.3289 (0.2539) 2.2137 (0.2341) 2.2341 (0.2606)
Mean Error (ivn) ↓ 0.0287 (0.0022) 0.0281 (0.0021) 0.0298 (0.0031)
MMD (ivn) ↓ 0.1023 (0.0377) 0.0957 (0.0348) 0.0881 (0.0317)
Generated Distribution Evaluation
MMD ↓ 0.0051 (0.0005) 0.0056 (0.0006) 0.0067 (0.0012)
MMD (ivn) ↓ 0.0231 (0.0071) 0.0254 (0.0072) 0.0217 (0.0064)
Table 9: Latent space regularization (as in $\beta$-VAE) on MNIST
Metric / Evaluation $\beta$=0.5 $\beta$=2 $\beta$=4
Validation Metrics
Validation ELBO (bpd) ↓ 0.2591 (0.0057) 0.3399 (0.0080) 0.3725 (0.0076)
Reconstruction Loss (bpd) ↓ 0.2367 (0.0059) 0.3016 (0.0094) 0.3317 (0.0083)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.1730 (0.0739) 1.7058 (0.0478) 1.8174 (0.1025)
Mean Error ↓ 0.0180 (0.0026) 0.0232 (0.0018) 0.0149 (0.0016)
MMD ↓ 0.0031 (0.0017) 0.0268 (0.0033) 0.0688 (0.0105)
Covariance Error (ivn) ↓ 1.9956 (0.2056) 2.3965 (0.2786) 2.6715 (0.3147)
Mean Error (ivn) ↓ 0.0258 (0.0017) 0.0359 (0.0038) 0.0408 (0.0053)
MMD (ivn) ↓ 0.0821 (0.0307) 0.1110 (0.0357) 0.1491 (0.0324)
Generated Distribution Evaluation
MMD ↓ 0.0038 (0.0001) 0.0132 (0.0014) 0.0437 (0.0086)
MMD (ivn) ↓ 0.0160 (0.0041) 0.0410 (0.0112) 0.0898 (0.0179)
D.1.4 Expressivity (MNIST)
[…]
Metric / Evaluation (15, 22) (22, 22) (22, 5) (50, 22) (50, 5)
Validation Metrics
Validation ELBO (bpd) ↓ 0.3777 (0.0100) 0.3892 (0.0119) 0.2956 (0.0032) 0.3655 (0.0088) 0.3073 (0.0098)
Reconstruction Loss (bpd) ↓ 0.3532 (0.0095) 0.3641 (0.0123) 0.2653 (0.0034) 0.3413 (0.0083) 0.2781 (0.0098)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 1.6402 (0.1417) 1.7153 (0.1730) 1.4769 (0.0512) 1.6964 (0.1163) 1.5230 (0.1056)
MAE_Mean ↓ 0.0206 (0.0032) 0.0214 (0.0020) 0.0243 (0.0019) 0.0215 (0.0023) 0.0194 (0.0015)
MMD ↓ 0.0183 (0.0016) 0.0199 (0.0015) 0.0119 (0.0008) 0.0241 (0.0050) 0.0138 (0.0010)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 2.8664 (0.3217) 2.9818 (0.3424) 2.3018 (0.2701) 2.7742 (0.3048) 2.3233 (0.2495)
MAE_Mean ↓ 0.0407 (0.0035) 0.0455 (0.0044) 0.0311 (0.0028) 0.0374 (0.0032) 0.0305 (0.0023)
MMD ↓ 0.1036 (0.0313) 0.1064 (0.0337) 0.0907 (0.0321) 0.1149 (0.0372) 0.0882 (0.0302)
Generated Distribution Evaluation
MMD ↓ 0.0094 (0.0006) 0.0091 (0.0015) 0.0061 (0.0008) 0.0110 (0.0021) 0.0070 (0.0017)
MMD (ivn) ↓ 0.0312 (0.0066) 0.0422 (0.0130) 0.0224 (0.0062) 0.0388 (0.0117) 0.0249 (0.0059)
[…]
Metric / Evaluation (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
Validation Metrics
Validation ELBO (bpd) ↓ 0.3423 (0.0089) 0.3147 (0.0105) 0.3779 (0.0178) 0.3027 (0.0063) 0.3702 (0.0208) 0.3160 (0.0124)
Reconstruction Loss (bpd) ↓ 0.3213 (0.0092) 0.2897 (0.0108) 0.3569 (0.0181) 0.2767 (0.0061) 0.3531 (0.0240) 0.2899 (0.0132)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 1.5155 (0.0766) 1.5477 (0.0886) 1.7862 (0.1572) 1.3584 (0.1025) 1.9985 (0.3544) 1.4030 (0.1516)
MAE_Mean ↓ 0.0141 (0.0014) 0.0178 (0.0019) 0.0201 (0.0030) 0.0148 (0.0031) 0.0210 (0.0054) 0.0157 (0.0029)
MMD ↓ 0.0251 (0.0021) 0.0180 (0.0026) 0.0282 (0.0039) 0.0137 (0.0018) 0.2226 (0.1939) 0.0139 (0.0030)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 2.5968 (0.2735) 2.2916 (0.2427) 2.8177 (0.3093) 2.2404 (0.2412) 2.9055 (0.3518) 2.3451 (0.2558)
MAE_Mean ↓ 0.0359 (0.0035) 0.0311 (0.0024) 0.0418 (0.0040) 0.0298 (0.0023) 0.0428 (0.0068) 0.0304 (0.0028)
MMD ↓ 0.0937 (0.0279) 0.1036 (0.0359) 0.1112 (0.0337) 0.0987 (0.0350) 0.2750 (0.0712) 0.0946 (0.0333)
Generated Distribution Evaluation
MMD ↓ 0.0123 (0.0012) 0.0092 (0.0013) 0.0135 (0.0024) 0.0067 (0.0007) 0.2112 (0.1967) 0.0056 (0.0006)
MMD (ivn) ↓ 0.0334 (0.0076) 0.0272 (0.0069) 0.0374 (0.0082) 0.0242 (0.0064) 0.2254 (0.0721) 0.0245 (0.0069)
D.2 Additional NVAE Figures
D.2.1 Additional examples of concept composition
Figure D.2.1 contains additional examples of composing concepts from the NVAE-based context module. Each row corresponds to actual generated samples from the trained model in different contexts:
  1. 1.
    The first row shows generated observational samples;
  2. 2.
    The second and third row contain generated single-node interventions;
  3. 3.
    The final row shows generated samples where two distinct concepts are composed together.
For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon. Recall that CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data.
Additional examples of both single- and multi-concept interventions over more contexts are provided in Appendices D.2.2D.2.4.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 8: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right). In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together. The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
D.2.2 Additional MNIST figures
Figure 9: Examples of single-node concept interventions in MNIST. (top) observational, scaled, shear (bottom) thic, thin, swel.
Figure 10: Examples of OOD multi-node concept interventions in MNIST.
D.2.3 Additional 3DIdent figures
[Uncaptioned image][Uncaptioned image]
Figure 11: Examples of single-node concept interventions in 3DIdent. (top) observational, background (bottom) object, spotlight.
Figure 12: Examples of OOD multi-node concept interventions in 3DIdent.
D.2.4 Additional CelebA figures
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 13: Examples of single-node concept interventions in CelebA. (top) observational, bags under eyes, black hair (bottom) blond hair, male, mouth slightly open.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 14: Examples of OOD multi-node concept interventions in CelebA.

### v2

@@
-22affiliationtext: Booth School of Business
+
 Abstract
 In designing generative models, it is commonly believed that in order to learn useful latent structure, we face a fundamental tension between expressivity and structure.
-In this paper we challenge this view by proposing a new approach to training arbitrarily expressive generative models that simultaneously learn disentangled latent structure.
-This is accomplished by adding a simple decoder-only module to the head of an existing decoder block that can be arbitrarily complex.
-The module learns to process concept information by implicitly inverting linear representations from an encoder.
-Inspired by the notion of intervention in causal graphical models, our module selectively modifies its architecture during training, allowing it to learn a compact joint model over different contexts.
-We show how adding this module leads to disentangled representations that can be composed for out-of-distribution generation.
-To further validate our proposed approach, we prove a new identifiability result that extends existing work on identifying structured representations in nonlinear models.
+In this paper we challenge this view by proposing a new approach to training arbitrarily expressive generative models that simultaneously learn causally disentangled concepts.
+This is accomplished by adding a simple context module to an arbitrarily complex black-box model, which learns to process concept information by implicitly inverting linear representations from the model’s encoder.
+Inspired by the notion of intervention in a causal model, our module selectively modifies its architecture during training, allowing it to learn a compact joint model over different contexts.
+We show how adding this module leads to causally disentangled representations that can be composed for out-of-distribution generation on both real and simulated data.
+The resulting models can be trained end-to-end or fine-tuned from pre-trained models.
+To further validate our proposed approach, we prove a new identifiability result that extends existing work on identifying structured representations.
+keywords:
+causal disentanglement; generative models; compositional generalization; concept learning; interventions.
 1 Introduction
@@
 Recent years have seen increased interest in understanding and exploring the representations produced by these models through evolving lines of work on structured representation learning, identifiability and interpretability, disentanglement, and causal generative models.
-This work is motivated in part by the desire to produce performant generative models that also capture meaningful, semantic latent spaces that enable out-of-distribution generation under perturbations.
-A key conceptual driver of this line of work is the tradeoff between flexibility and structure, or expressivity and interpretability: It is widely believed that to learn structured, interpretable representations, model capacity must be constrained, sacrificing flexibility and expressivity.
-As described in Section 3, this intuition is supported by a growing body of theoretical work on nonlinear ICA (Hyvärinen and Pajunen, 1999), disentanglement (Bengio, 2013), and causal representation learning (Schölkopf et al., 2021).
-On the practical side, methods that have been developed to learn structured latent spaces tend to be bespoke to specific data types and models, and typically impose significant limitations on the flexibility and expressivity of the underlying models.
+This work is motivated in part by the desire to produce performant generative models that also capture meaningful, semantic latent spaces that enable out-of-distribution (OOD) generation.
+A key driver behind this work is the tradeoff between flexibility and structure, or expressivity and interpretability: Conventional wisdom suggests that to learn structured, interpretable representations, model capacity must be constrained, sacrificing flexibility and expressivity.
+This intuition is supported by a growing body of work on nonlinear ICA (Hyvärinen and Pajunen, 1999), disentanglement (Bengio, 2013), and causal representation learning (Schölkopf et al., 2021).
+On the practical side, methods that have been developed to learn structured latent spaces tend to be bespoke to specific data types and models, and typically impose limitations on the flexibility and expressivity of the underlying models.
 Moreover, the most successful methods for learning structured representations typically impose fixed, known structure a priori, as opposed to learning this structure from data.
-At the same time, there is a growing body of work that suggests generative models already learn surprisingly structured latent spaces (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015, see also Section 3).
-This in turn suggests that existing models are already “close” to capturing the desired structure, and perhaps only small modifications are needed.
-Since we already have performant models that achieve state-of-the-art results in generation and discrimination, do we need to re-invent the wheel to achieve these goals?
-Our hypothesis is that we should be able to leverage the expressiveness of these models to build new models that learn latent structure from scratch.
-We emphasize that our goal is not to explain or interpolate the latent space of a pre-trained model, but rather to leverage known architectures to train a new model end-to-end from scratch.
-Figure 1: Overview of our approach.
-Given a black-box encoder-decoder architecture (blue), we propose to append a context module (red) to the head of the decoder.
+Figure 1: Overview (see (4) for notation map).
+Given a black-box encoder-decoder architecture (blue), we propose to prepend a context module (red) to the decoder.
 (left) Instead of passing the output of the encoder directly to the decoder (blue box + blue arrow), the embeddings $\mathbf{e}$ are passed through the context module, consisting of three distinct layers.
 The output of this module is then passed into the decoder.
-(right) The model learns to compose different concepts OOD, in this case object and background colour, to values that never appear in the training data.
-In this paper, we adopt this perspective: We start with a black-box encoder-decoder architecture, upon which we make no assumptions, and then augment this model to learn structured representations.
-The idea is that the black-box is an architecture that is already known to perform well on downstream tasks such as sampling and prediction, and so is flexible enough to capture complex patterns in data.
-We then introduce a modular, end-to-end differentiable architecture for learning structured representations, augmenting the original model with a simple decoder-only module that retains its existing capabilities while enabling concept identification and intervention as well as composing together multiple concept interventions for out-of-distribution (OOD) generation.
-The resulting latent structure is learned fully end-to-end, with no fixed structure imposed a priori.
-The motivation is to provide a framework for taking existing performant architectures and to train a new model that performs just as well, but with the added benefit of learning structured representations without imposing specific prior knowledge or structure a priori.
-These representations may be useful for interpretability, downstream tasks, or other purposes.
-Since the model, unlike previous approaches, imposes no latent structure a priori, we theoretically validate the approach with a novel identifiability result for structured representations recovered by the model.
-To explore the implications of this modification, we run experiments including OOD generation.
-An attractive feature of our approach is that it makes running carefully controlled experiments and ablations particularly simple: We can carefully control hyperparameters across different models, ensuring that any differences in performance are attributable to specifically controlled architectural modifications.
-As a result, we contribute to a growing body of empirical work on the OOD capabilities of generative models by introducing a new set of ablations that can be used to guide the development of new methods.
+(right) The model learns to compose different concepts OOD, e.g., object and background colour, to values that never appear together in the training data.
+At the same time, there is also a growing body of work that suggests generative models already learn surprisingly structured latent spaces (e.g.
+Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015).
+This in turn suggests that existing models are already “close” to capturing useful structure, and perhaps only small modifications are needed.
+Since we already have performant models that achieve state-of-the-art results in generation and prediction, do we need to re-invent the wheel to achieve these goals?
+Our hypothesis is that we should be able to leverage the expressiveness of these models to build new models that learn latent structure from scratch.
+We emphasize that our goal is not to explain or interpolate the latent space of a pre-trained model, but rather to leverage known architectures to train a new model, either end-to-end or by fine-tuning an existing model.
+In this paper, we adopt this perspective: We start with an arbitrarily complex black-box model, upon which we make no assumptions, then augment this model to learn causally disentangled representations.
+The idea is that the black-box model is already known to perform well on downstream tasks (e.g., generation, prediction), and so is flexible enough to capture complex patterns in data.
+We introduce a modular, end-to-end differentiable architecture for learning causally disentangled representations (Bengio, 2013; Schölkopf et al., 2021) that augments the original model with a simple module that retains its existing capabilities while enabling concept identification and intervention as well as composing together multiple concept interventions for OOD generation (Figure 1).
+The resulting latent structure is learned entirely from data, with no fixed structure imposed a priori.
+The motivation is to provide a framework for taking existing performant architectures and to train a new model that performs just as well, but with the added benefit of learning structured representations without imposing specific prior knowledge or structure.
+Since the model, unlike previous approaches, does not impose rigid latent structure, we theoretically validate the approach with a novel identifiability result for structured representations recovered by the model.
+For a detailed discussion of related work, see Appendix A.
 Contributions
-Our main contributions can be summarized as follows:
+More specifically, we make the following contributions:
   1. 1.
-    We introduce a simple module that can be attached to the head of an existing decoder block to learn concept representations by splitting each concept into a tensor slice, where each slice represents an interventional context in a reduced form structural equation model (SEM).
-This module can be attached to the head of any decoder block and trained end-to-end.
+    We introduce a context module that is attached to a decoder to learn concept representations by splitting each concept into a tensor slice, where each slice represents an interventional context in a reduced form structural equation model (SEM).
+This module can be attached to any decoder and trained end-to-end or fine-tuned (Figure 1).
   2. 2.
-    We conduct experiments on disentanglement and OOD generation on several benchmark datasets (3DIdent, CelebA, MNIST) as well as controlled simulations.
-To this end, we introduce a simple simulated visual environment for testing disentanglement and OOD generation.
-This allows us to carefully test different ablations under controlled settings.
+    We evaluate this module through quantitative and qualitative experiments on OOD generation, which is distinct from and significantly more challenging than OOD reconstruction as in prior work (e.g.
+Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022).
+We perform experiments on OOD generation on several real datasets as well as carefully controlled simulations.
   3. 3.
-    To illustrate the adaptivity of our module to different architectures, we include experiments with several different VAE architectures including NVAE (Vahdat and Kautz, 2020).
-As a matter of independent interest, we prove an identifiability result under concept interventions, and discuss how the proposed architecture can be interpreted as an approximation to this model.
-Full technical proofs are deferred to the appendix.
+    We illustrate the adaptivity of our module to different architectures (including NVAE, Vahdat and Kautz, 2020) and against multiple ablations that control model capacity, data contexts, training protocols, and hyperparameters, ensuring that any differences in performance are directly attributable to the context module.
+  4. 4.
+    We fine-tune existing, pre-trained models after attaching the context module, giving substantial computational savings compared to end-to-end training while improving the learned representations and lending OOD generation abilities to pre-trained models.
+As a matter of independent interest, we prove an identifiability result under concept interventions (Theorem 3.3), and discuss how the proposed architecture can be interpreted as an approximation to this model.
+We also introduce quad, a simple simulated visual environment for evaluating causal disentanglement, compositional abilities, and OOD generation.
+Due to space limitations, full technical proofs, along with detailed literature review and experimental details, can be found in the appendix.
 2 Preliminaries
-Let $\mathbf{x}$ denote the observations, e.g., pixels in an image, and $\mathbf{z}$ denote hidden variables, e.g.
-latent variables that are to be inferred from the pixels.
-We are interested in training generative models with an encoder-decoder architecture, such as a variational autoencoder (VAE).
+Let $\mathbf{x}$ denote the observations, e.g., pixels in an image, and $\mathbf{z}$ denote hidden variables, e.g., latent variables that are to be inferred from the pixels.
 A typical generative model consists of a decoder $p_{\theta}(\mathbf{x}\,|\,\mathbf{z})$ and an encoder $q_{\phi}(\mathbf{z}\,|\,\mathbf{x})$.
 After specifying a prior $p_{\theta}(\mathbf{z})$, this defines a likelihood by
-[…]
-Both the decoder and encoder are specified by deep neural networks that will be trained end-to-end via standard techniques (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015).
-These networks are often chosen to capture implicit assumptions about the structure of the latent space, such as hierarchical or graphical structure, equivariance, etc.
-In this work, we focus on VAEs since they are often use to represent structured and semantic latent spaces, which is our focus.
+$p_{\theta}(\mathbf{x})=\int p_{\theta}(\mathbf{x}\,|\,\mathbf{z})p_{\theta}(\mathbf{z})\operatorname{d\!}{\mathbf{z}}.$ (1)
+Both the decoder and encoder are specified by deep neural networks that are trained end-to-end via standard techniques (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015).
+In this work, we focus on variational autoencoders (VAEs) since they are often use to represent structured and semantic latent spaces, which is our focus.
 Unlike traditional structured VAEs, we do not impose a specific structure a priori.
-Our starting point is the following incongruence: On the one hand, it is well-established that the latent spaces of VAEs are typically entangled and semantically misaligned, fail to generalize OOD, and suffer from posterior collapse (which has been tied to latent variable nonidentifiability, Wang et al., 2021).
-On the other hand, VAEs still learn highly structured latent spaces that can be traversed and interpolated, are “nearly” identifiable (Willetts and Paige, 2021; Reizinger et al., 2022), and represent abstract concepts linearly.
-Thus, our hypothesis is that sufficiently flexible models do capture meaningful structure internally, just not in a way that is interpretable or meaningful in practice.
-So, to learn latent structure, we build off the already performant embeddings of a generative model.
-Our set-up is the following: We are given a black-box VAE, consisting of an encoder and a decoder, on which we make no assumptions other than the encoder outputs a latent code corresponding to $\mathbf{x}$.
+Causal disentanglement
+Our goal is to learn distinct latent factors that can be composed together. This is closely related to causal disentanglement, which we recall informally here:
+Definition 2.1 (Bengio, 2013; Thomas et al., 2017).
+A causally disentangled latent space consists of separately controllable (intervenable) latent factors.
+In particular, we want to manipulate and intervene on latent factors in a causally meaningful way; crucially this allows us to compose distinct concepts together without needing to learn the full causal graph in the latent space.
+This notion of disentanglement is also different from the axis-alignment notion (Kim and Mnih, 2018; Chen et al., 2018).
+The “causal” aspect comes from interpreting “separately controllable” to mean “able to intervene on independent causal mechanisms”, as suggested by Locatello et al.
+(2019); see also Schölkopf et al.
+(2021).
+See Appendix B for formal details.
+Rather than attempt to reconstruct the full causal DAG model (as in Yang et al., 2021; Squires et al., 2023; Buchholz et al., 2023), we aim for the more practical goal of learning different interventional contexts through distributional invariances; thus our approach strikes a balance between easy-to-learn unstructured latent spaces and hard-to-learn but desirable fully structured causal DAG latent spaces.
+Measuring causal disentanglement
+A standard approach to quantifying causal disentanglement is to evaluate the recovery of the latent causal graph; since our explicit goal is to avoid learning this graph, we propose a different metric that is more suitable to downstream applications.
+To measure causal disentanglement, we use OOD generation on novel interventions as a metric.
+We train models so that they never see examples of certain latent factors composed together (equivalently, multiple simultaneous interventions on distinct factors), so these examples are genuinely OOD.
+By withholding these examples during training and building a model that is capable of intervening on multiple factors, we can compare the quality of OOD samples generated from our model to the held-out OOD samples.
+In our experiments, we use the sliced Wasserstein (SW) metric (18); see Appendix D.1 for details.
+OOD reconstruction vs. OOD generation
+Previous work has evaluated the OOD capabilities of generative models by using OOD reconstruction as a metric (e.g.
+Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022).
+There is a crucial difference between OOD reconstruction and OOD generation: By OOD generation, we mean the ability to conditionally generate and combine novel interventions on existing concepts.
+For example, suppose during training the model only sees small, red objects along with big, blue objects.
+At test time, we wish to generate OOD samples of big, red objects or small, blue objects that were never seen during training.
+The crucial difference is that while we can always attempt to reconstruct any sample (i.e., OOD or not) and evaluate its reconstruction error, it is not always possible to have the model generate a new OOD sample.
+Unless the model learns specific latent factors corresponding to concepts such as size or colour, we cannot control (i.e., intervene) the size or colour of random samples from the model.
+Thus, OOD generation not only evaluates the ability of a model to compose learned concepts in new ways, but also the ability of a model to identify and capture underlying concepts of interest.
+Accordingly, we argue that OOD generation is more appropriate to evaluate causal disentanglement since it captures both the model’s ability to learn concepts as distinct latent factors as well as compose them together in novel ways.
+Set-up and approach
+Our set-up is the following: We have a black-box encoder and decoder, on which we make no assumptions other than the encoder outputs a latent code corresponding to $\mathbf{x}$.
 In the notation of (1), this corresponds to $\mathbf{z}$, however, to distinguish the black-box embeddings from our model, we will denote the black-box embeddings hereafter by $\mathbf{e}$.
 Our plan is to work solely with the embeddings $\mathbf{e}$ and learn how to extract linear concept representations from $\mathbf{e}$ such that distinct concepts can be intervened upon and composed together to create novel, OOD samples.
-Crucially, we do not modify the black-box architecture in any way.
-3 Related work
-Structured generative models
-To provide structure such as hierarchical, graphical, causal, and disentangled structures as well as other inductive biases in the latent space, there has been a trend towards building structured generative models that directly impose this structure a priori.
-Early work looked at incorporating fixed, known structure into generative models, such as autoregressive, graphical, and hierarchical structure (Germain et al., 2015; Johnson et al., 2016; Sønderby et al., 2016; Webb et al., 2018; Weilbach et al., 2020; Ding et al., 2021; Mouton and Kroon, 2023).
-This was later translated into known causal structure (Kocaoglu et al., 2017; Markham et al., 2023).
-When the latent structure is unknown, several techniques have been developed to learn useful (not necessarily causal) structure from data (Li et al., 2018; He et al., 2019; Wehenkel and Louppe, 2021; Kivva et al., 2022; Moran et al., 2022).
-More recently, based on growing interest in disentangled (Bengio, 2013) and/or causal (Schölkopf et al., 2021) representation learning, methods that learn causal structure have been developed (Markham and Grosse-Wentrup, 2020; Moraffah et al., 2020; Yang et al., 2021; Ashman et al., 2022; Shen et al., 2022; Kaltenpoth and Vreeken, 2023).
-Causal representation learning
-Causal representation learning (Schölkopf et al., 2021), which involves learning causal structure in the latent space, is a rapidly developing area that has produced fundamental results on theoretical aspects of identifiability (Brehmer et al., 2022; Shen et al., 2022; Lachapelle et al., 2022; Moran et al., 2022; Kivva et al., 2022; Buchholz et al., 2023; Gresele et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023; Morioka and Hyvärinen, 2023; Liu et al., 2023; Mameche et al., 2023; Yao et al., 2023; Varici et al., 2023; Sturma et al., 2023; Xu et al., 2024; Zhang et al., 2024; Li et al., 2024; Varici et al., 2024; Bing et al., 2024; Ahuja et al., 2024; Talon et al., 2024; Yao et al., 2024).
-Recent work has also pushed in the direction of identifying concepts (Leemann et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025).
-Our work draws inspiration from this line of work, which articulates precise conditions under which latent structure can be recovered in principle.
-By contrast, our focus is on methodological aspects of exploiting learned causal invariances in practice.
-Linear representations
-Generative models are known to represent concepts linearly in embedding space (e.g.
-Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015); see also Levy and Goldberg (2014); Arora et al.
-(2016); Gittens et al.
-(2017); Allen and Hospedales (2019); Ethayarajh et al.
-(2018); Seonwoo et al.
-(2019).
-This phenomenon has been well-documented over the past decade in both language models (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Conneau et al., 2018; Tenney et al., 2019; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Li et al., 2023; Park et al., 2023; Gurnee et al., 2023; Jiang et al., 2024) and computer vision (Radford et al., 2015; Raghu et al., 2017; Bau et al., 2017; Engel et al., 2017; Kim et al., 2018; Chen et al., 2020; Wang et al., 2023; Trager et al., 2023).
-Our approach actively exploits this tendency by searching for concept representations as linear projections of the embeddings learned by a black-box model.
-OOD generalization
-A growing line of work studies the OOD generalization capabilities of generative models, with the general observation being that existing methods struggle to generalize OOD (Xu et al., 2022; Mo et al., 2024; Montero et al., 2022, 2021; Schott et al., 2022).
-It is worth noting that most if not all of this work evaluates OOD generalization using reconstruction on held-out OOD samples, as opposed to generation.
-For example, a traditional VAE may be able to reconstruct held-out samples, but it is not possible to actively sample OOD.
-See Section 5.1 for more discussion.
+Our approach is motivated by the following incongruence: On the one hand, it is well-established that the latent spaces of generative models are typically entangled and semantically misaligned, fail to generalize OOD, and suffer from posterior collapse (which has been tied to latent variable nonidentifiability, Wang et al., 2021).
+On the other hand, they still learn structured latent spaces that can be interpolated, are “nearly” identifiable (Willetts and Paige, 2021; Reizinger et al., 2022), and represent abstract concepts linearly (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015).
+So, to learn latent structure, we simply need to extract the right concept representations from the already performant embeddings of existing architectures and learn how to compose them together.
 3 Architecture
-We start with a black-box encoder-decoder architecture and assume there are $d_{c}$ concepts of interest, denoted by $\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$.
-Our objectives are two-fold:
+We start with a black-box encoder-decoder architecture along with $d_{c}$ concepts of interest, denoted by $\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$. Our objectives are three-fold:
   1. 1.
-    To learn structure between concepts from black-box embeddings $\mathbf{e}$ as linear projections $C\mathbf{e}$;
+    To extract concepts from black-box embeddings $\mathbf{e}$ as linear projections $C\mathbf{e}$;
   2. 2.
-    To compose concepts together in a single, transparent model that captures how different concepts are related.
-We seek to learn these concepts in an end-to-end, differentiable manner.
+    To compose concepts together in a single, transparent model that captures how different concepts are related;
+  3. 3.
+    To accomplish both of these objectives without restricting the black-box encoder or decoder architectures, and without sacrificing sampling quality.
 A key intuition is that composition can be interpreted as a type of intervention in the latent space.
 This is sensible since interventions in a causal model are a type of nontrivial distribution shift.
-Thus, the problem takes on a causal flavour which we exploit to build our architecture.
-The difficulty with this from the causal modeling perspective is that encoding structural assignments and/or causal mechanisms directly into a feed-forward neural network is tricky, because edges between nodes within the same layer aren’t allowed in a feed-forward network but are required for the usual DAG representation of a causal model.
-To bypass this, we use the reduced form of a causal model which has a clear representation as a bipartite directed graph, with arrows only from exogenous to endogenous variables, making it conducive to being embedded into a feedforward NN.
-The tradeoff is that we do not learn a causal structure (i.e., a causal DAG), however, this enables the model to perform interventions directly in the latent space and to exploit invariances between interventional contexts.
+Thus, the problem takes on a causal flavour which we leverage to build our architecture.
+The difficulty with this from the causal modeling perspective is that encoding structural assignments and/or causal mechanisms directly into a feed-forward neural network is tricky, because edges between nodes within the same layer are not allowed in a feed-forward network but are required for the usual DAG representation of a causal model.
+To bypass this, we use the reduced form of a causal model which has a clear representation as a bipartite directed graph, with arrows only from exogenous to endogenous variables, making it conducive to being embedded into a neural network.
+The tradeoff is that this does not learn a causal DAG, however it still enables the model to perform interventions directly in the latent space and to leverage invariances between interventional contexts.
 3.1 Overview
@@
 Thus,
-[…]
-Consistent with existing empirical work (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015), we represent concepts as linear projections of the embeddings $\mathbf{e}$: Each concept $\mathbf{c}_{j}$ can be approximated as $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$.
+$\mathbf{x}\overset{\textup{encoder}}{\xrightarrow{\hskip 16.38895pt}}\mathbf{e}\overset{\textup{decoder}}{\xrightarrow{\hskip 16.38895pt}}\widehat{\mathbf{x}}$.
+Based on an extensive body of empirical work that shows concepts are linearly represented (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015, see Appendix A for more discussion), we represent concepts as linear projections of the embeddings $\mathbf{e}$: Each concept $\mathbf{c}_{j}$ can be approximated as $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$.
 This is modeled via a linear layer $\mathbf{c}\to\mathbf{e}$ that implicitly inverts this relationship in the decoder.
-As a result, it makes sense to model the relationships between concepts with a linear SEM:
-$\mathbf{c}_{j}=\sum_{k=1}^{d_{c}}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{k},\quad\alpha_{kj}\in\mathbb{R}.$
-By reducing this SEM, we deduce that
-$\mathbf{c}=A_{0}\boldsymbol{\varepsilon},\text{ where }\left\{\begin{aligned} \mathbf{c}&=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})\\ \boldsymbol{\varepsilon}&=(\boldsymbol{\varepsilon}_{1},\ldots,\boldsymbol{\varepsilon}_{d_{c}})\end{aligned}\right..$ (1)
+We further model the relationships between concepts with a linear SEM:
+$\mathbf{c}_{j}=\sum_{k=1}^{d_{c}}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{k},\quad\alpha_{kj}\in\mathbb{R}.$ (2)
+By reducing this SEM and solving for $\mathbf{c}$, we deduce that
+$\mathbf{c}=A_{0}\boldsymbol{\varepsilon},\text{ where }\left\{\begin{aligned} \mathbf{c}&=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})\\ \boldsymbol{\varepsilon}&=(\boldsymbol{\varepsilon}_{1},\ldots,\boldsymbol{\varepsilon}_{d_{c}})\end{aligned}\right.$.
+(3)
+This is known as the reduced form of the SEM (2).
 We model this with a linear layer $\boldsymbol{\varepsilon}\to\mathbf{c}$, where the weights in this layer correspond to the matrix $A_{0}$.
@@
 This allows for $\dim(\mathbf{z})$ to be larger and more expressive than $d_{c}$ (in practice, we set $\dim(\mathbf{z})$ to be a multiple of $\dim(\boldsymbol{\varepsilon})$), and for $\boldsymbol{\varepsilon}$ to be potentially non-Gaussian.
-The final decoder architecture can be decomposed, at a high-level, as follows:
+The final decoder architecture can be decomposed at a high-level as follows (Figure 1):
 $\mathbf{z}\longrightarrow\boldsymbol{\varepsilon}\longrightarrow\mathbf{c}\longrightarrow\mathbf{e}\longrightarrow\mathbf{x}.$ (4)
@@
 3.2 Details
-Our goal is to augment the decoder of a generative model (without modifying the internals of the existing decoder) in such a way to identify $d_{c}$ latent concepts from a black-box decoder.
-Moreover, we seek to accomplish this without sacrificing expressivity: The input layer to the decoder as well as the output layer of the encoder should both be arbitrary.
-This not only allows for arbitrary flexibility in both the latent dimension as well as the structure of the black-box encoder-decoder and its latent space, but also makes implementation significantly simpler: The learning of latent concepts is completely abstracted away from the encoding step.
-To accomplish this, we assume given a black-box encoder-decoder pair, denoted by $\operatorname{enc}(\mathbf{x})$ and $\operatorname{dec}(\mathbf{z})$, respectively.
-We modify the decoder by appending a context module to the head of the decoder.
-This context module is divided into three layers: A representation layer, an intervention layer, and an expressive layer.
+We assume given a black-box encoder-decoder pair, denoted by $\operatorname{enc}(\mathbf{x})$ and $\operatorname{dec}(\mathbf{z})$, respectively.
+We modify the decoder by appending a context module to the decoder.
+The context module has three layers: A representation layer, an intervention layer, and an expressive layer.
   1. 1.
     The first representation layer, as the name suggests, learns to represent concepts by implicitly inverting their linear representations $\mathbf{c}_{j}=C_{j}\mathbf{e}$ from the embeddings of the encoder.
-This is a linear layer between $\mathbf{c}\to\mathbf{e}$.
+This is a linear layer between $\mathbf{c}\to\mathbf{e}$ in (4).
   2. 2.
-    The intervention layer embeds these concept representations into an overall reduced form SEM as suggested by (1).
-Each concept corresponds to its own context where it has been intervened upon.
+    The intervention layer embeds these concept representations into the reduced form SEM (3).
+Each concept has its own context where it has been intervened upon.
 A key part of this layer is how it can be used to learn and enforce interventional semantics through this shared SEM.
@@
 In particular, we do not assume a known causal DAG or even a known causal order.
-This stands in contrast to previous work on causal generative models such as CausalGAN (Kocaoglu et al., 2017) and CausalVAE (Yang et al., 2021).
+This stands in contrast to previous work on causal generative models (e.g.
+Kocaoglu et al., 2017; Yang et al., 2021).
 Concept interventions
-For simplicity, assume here that the concepts are each one-dimensional with $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$; generalization to multi-dimensional concepts is more or less straightforward and explained below.
+Assume for now that the concepts are each one-dimensional with $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$; extensions to multi-dimensional concepts are straightforward and explained below.
 The structural coefficients $\alpha_{kj}\in\mathbb{R}$ capture direct causal effects between concepts, with $\alpha_{kj}\neq 0$ indicating the presence of an edge $\mathbf{c}_{k}\to\mathbf{c}_{j}$.
-An intervention on the $j$th concept requires deleting all these incoming edges; i.e., setting $\alpha_{\cdot j}=0$, updating the outgoing edges from $\boldsymbol{\varepsilon}_{j}$, as well as replacing $\boldsymbol{\varepsilon}_{j}$ with a new $\boldsymbol{\varepsilon}_{j}^{\prime}$, which results from updating the incoming edges to $\boldsymbol{\varepsilon}_{j}$ in the preceding layer.
+An intervention on the $j$th concept requires deleting each incoming edge; i.e., setting $\alpha_{\cdot j}=0$, updating the outgoing edges from $\boldsymbol{\varepsilon}_{j}$, as well as replacing $\boldsymbol{\varepsilon}_{j}$ with a new $\boldsymbol{\varepsilon}_{j}^{\prime}$, which results from updating the incoming edges to $\boldsymbol{\varepsilon}_{j}$ in the preceding layer.
 Thus, during training, when we intervene on $\mathbf{c}_{j}$, we zero out the row $\alpha_{\cdot j}$ while replacing the column $\alpha_{j\cdot}$ with a new column $\beta_{j\cdot}$ that is trained specifically for this interventional setting.
 The result is $d_{c}+1$ intervention-specific layers: $A_{0}$ for the observational setting, and $A_{1},\ldots,A_{d_{c}}$ for each interventional setting.
-This is captured as a three-way tensor where each slice of this tensor captures a different interventional context that corresponds to intervening on different concepts.
-At inference time, to generate interventional samples, we simply swap out the $A_{0}$ for $A_{j}$.
-Moreover, multi-target interventions can be sampled by zero-ing out multiple rows and substituting multiple $\beta_{j\cdot}$’s and $\boldsymbol{\varepsilon}_{j}$’s.
+This yields a three-way tensor where each slice of this tensor captures a different context that corresponds to intervening on different concepts.
+At inference time, to generate interventional samples we simply swap out the $A_{0}$ for $A_{j}$.
+Moreover, multi-target interventions can be sampled by zeroing out multiple rows and substituting multiple $\beta_{j\cdot}$’s and $\boldsymbol{\varepsilon}_{j}$’s.
 Multi-dimensional concepts
-In the causal model described above, we assumed $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$ for simplicity.
-Everything carries through if we allow additional flexibility with $\dim(\mathbf{c}_{j})>1$ and $\dim(\boldsymbol{\varepsilon}_{j})>1$, potentially even with $\dim(\mathbf{c}_{j})\neq\dim(\boldsymbol{\varepsilon}_{j})$.
+Everything above goes through if we allow additional flexibility with $\dim(\mathbf{c}_{j})>1$ and $\dim(\boldsymbol{\varepsilon}_{j})>1$, potentially even with $\dim(\mathbf{c}_{j})\neq\dim(\boldsymbol{\varepsilon}_{j})$.
 In practice, we implement this via two width parameters $0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ and $0pt_{c}=\dim(\mathbf{c}_{j})$.
@@
 Identifiability
-An appealing aspect of this architecture is that under certain assumptions, it can be viewed as an approximation to an identifiable model over disentangled concepts.
+An appealing aspect of this architecture is that it can be viewed as an approximation to an identifiable model over causally disentangled concepts.
 Formally, we assume given observations $\mathbf{x}$, concepts $\mathbf{c}$, and a collection of embeddings $\mathbf{e}$ such that $\mathbf{x}=f(\mathbf{e})$.
@@
 Then we have the following identifiability result:
-Theorem 4.1 (Identifiability of linear representations).
-Assume that the rows of each $C_{j}$ are chosen from a linearly independent set and $f$ is injective and differentiable.
-Then, given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$.
-We do not claim to identify a causal graph, only the latent distribution $p(\mathbf{c})$ over the concepts and its functional structure $C_{j}$ within the data-generating process.
-As opposed to solving a causal representation learning problem, the causal semantics used in this paper are simply being ported into the generative modeling framework to give solutions to the problem of learning latent distribution structure in a completely unsupervised fashion.
-Theorem 4.1 shows that the set problems where there exists a structured latent distribution to be identified is nonempty.
-Moreover, the structure of these identifiable distributions exactly fits our model architecture, meaning that set of problems in which the proposed method will approximate a fully disentangled latent space is nonempty.
-In other words, there exist data-modeling problems that are solvable by the novel approach introduced in this paper.
-The technical difficulty in the proof of Theorem 4.1 is analyzing the behaviour of interventions in this model; since our main focus is on methodological aspects and the technical details are fairly involved, these details are deferred to the appendix.
+Theorem 3.3 (Identifiability).
+Assume that the rows of each $C_{j}$ come from a linearly independent set and $f$ is injective and differentiable.
+Then, given single-node interventions on each concept $\mathbf{c}_{j}$, the representations $C_{j}$ and latent concept distribution $p(\mathbf{c})$ are identifiable.
+For a formal statement, see Theorem B.1 in Appendix B.
+As opposed to solving a causal representation learning problem, the causal semantics used in this paper have been ported into a generative model that give solutions to the problem of learning latent distribution structure, as in Theorem 3.3.
+We note also that the proof of Theorem 3.3 itself is nontrivial, and involves tracing the effects of interventions through the linear representations; since our focus is on implementation and practical aspects, we defer further discussion to Appendix B.
 3.3 Summary of architecture
@@
   1. 1.
+    Black-box.
 There is no coupling between the embedding dimension and the number of concepts, or the complexity of the SEM.
-In particular, this architecture allows for arbitrary black-box encoder-decoder architectures to be used to learn embeddings, from which a causal model is then trained on top of.
-This can be trained end-to-end, fine-tuned after pre-training, or even frozen, allowing for substantial computational savings.
-In our experiments, we illustrate this with both standard convolutional networks as well as NVAE (Vahdat and Kautz, 2020), which is a deep hierarchical VAE with millions of parameters.
+This allows for arbitrary black-box encoder-decoder architectures to be used to learn embeddings, from which a causal model is then trained on top of.
+This can be trained end-to-end, fine-tuned after pre-training, or even partially frozen, allowing for substantial computational savings.
   2. 2.
-    The intervention layer is a genuine causal model that provides rigourous causal semantics that allow for sampling from arbitrary concept interventions, including interventions that have not been seen during training.
-But note that we do not claim to learn the structural coefficients, only the reduced model which is sufficient for concept interventions and sampling.
+    Causality.
+The intervention layer is a genuine causal model that provides rigorous causal semantics that allow sampling from arbitrary concept interventions, including interventions that have not been seen during training.
   3. 3.
+    Identifiability.
 The architecture is based on an approximation to an identifiable concept model (Theorem 3.3), which provides formal justification for the intervention layer as well as reproducibility assurances.
   4. 4.
-    The context module is itself arbitrarily flexible, meaning that there is no risk of information loss in representing concepts with the embeddings $\mathbf{e}$.
-Of course, information loss is possible if we compress this layer too much (e.g., by choosing $d_{c}$, $k$, $0pt_{\varepsilon}$, or $0pt_{c}$ too small), but this is a design choice and not a constraint of the architecture itself.
-As a consequence, the only tradeoff between representational capacity and causal semantics is design-based: The architecture itself imposes no such constraints.
-The causal model can be arbitrarily flexible and chosen independent of the black-box encoder-decoder pair, which is also allowed to be arbitrarily flexible.
-5 Empirical results
-We divide the presentation here into two main parts: small models on (semi-)synthetic datasets (Section 5.1) versus larger models on more challenging datasets (Section 5.2).
-We emphasize key findings here but include complete results and experimental details in the appendix.
-5.1 Small models on (semi-)synthetic data
-To systematically investigate the interplay between architectural design, learned representations, and OOD performance, we begin with small-scale experiments with controlled environments.
-This setting enables rapid iteration, comprehensive ablation studies, and access to ground-truth metrics unavailable in complex real-world data.
-We evaluate models on two benchmarks:
-  • MNIST: We use Morpho-MNIST (Castro et al., 2019) to allow for different interventional contexts, such as making digits thicker or thinner.
-  • quad: A novel semi-synthetic visual environment with disentangled latents (colour, shape, size, orientation) and explicit OOD compositionality challenges (e.g., generating unseen combinations of background quadrant colours and central object attributes).
-MNIST
-On MNIST, we see (Table 1) results of our module across different ablations.
-The first column (Context module) is our module incorporated into a lightweight convolutional VAE.
-Ablation 1 is an analogously lightweight standard implementation of $\beta$-VAE trained only on the observational context.
-Ablation 2 is $\beta$-VAE trained on the full pooled dataset.
-Ablation 3 is similar to the context module, except that it ignores the contexts and pools everything into a single context (like Ablation 2).
-These results indicate that the cost of the structured representation learned by the context module is only a slight degradation in the numerical metrics.
-Table 1: Ablation on MNIST (200 epochs)
-Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
-Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
-Reconstructed Distribution Evaluation
-MMD (obs) ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
-MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
-Generated Distribution Evaluation
-MMD (obs) ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
-MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
-Figure 2: Example images obtained from the quad dataset.
-(top) Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention).
-(bottom) Each subfigure corresponds to a different double-concept intervention context, indicated by the title.
-Double-concept interventions were not included in the training data, allowing for genuinely OOD evaluation.
-Controlled simulations on quad
-While evaluations on unsimulated data are the gold standard for evaluating models, simulated environments are valuable for conducting controlled experiments under given conditions.
-To provide a controllable, synthetic test bed for evaluating composition in a visual environment, we developed a simple semi-synthetic benchmark, visualized in Figure 2.
-quad is a visual environment defined by 8 concepts:
+    Flexibility.
+The context module is itself arbitrarily flexible, so there is no risk of information loss in representing concepts with the embeddings $\mathbf{e}$.
+Of course, information loss is possible if we compress this layer too much (e.g., by choosing $d_{c}$, $k$, $0pt_{\varepsilon}$, or $0pt_{c}$ too small), but this is a design choice and not a constraint of the module itself.
+Thus the only tradeoff between representational capacity and causal semantics is design-based: The architecture itself imposes no constraints.
+The causal model can be arbitrarily flexible and chosen independent of the encoder-decoder pair, which are also allowed to be arbitrary.
+4 Empirical evaluation
+[…]
   1. 1.
-    quad1: The colour of the first quadrant;
+    Concept learning (Section 4.1): How well does the context module learn representations that correspond to each concept?
+To measure this, we generate interventional samples by intervening on a single concept, and directly compare these interventional samples to held-out samples using the sliced Wasserstein (SW) metric (18).
   2. 2.
-    quad2: The colour of the second quadrant;
+    Composition (Section 4.2): How well does the context module learn to compose concepts together in novel ways that have not been seen during training?
+To measure this, we generate samples by intervening on multiple concepts, and directly compare these new samples to held-out samples using the SW metric.
+By design, interventions on multiple concepts are never seen during training, so this is genuinely OOD.
   3. 3.
-    quad3: The colour of the third quadrant;
+    Reconstruction (Section 4.3): How well does the context module reconstruct samples from each context?
+This can be evaluated using standard measures of generation quality; we use bits per dimension (BPD) for comparison with prior work.
+Since this is computed across all contexts, this captures the model’s ability to reconstruct samples from diverse contexts.
   4. 4.
-    quad4: The colour of the fourth quadrant;
-  5. 5.
-    size: The size of the center object;
-  6. 6.
-    orientation: The orientation (angle) of the center object;
-  7. 7.
-    object: The colour of the center object;
-  8. 8.
-    shape: The shape of the center object (circle, square, pill, triangle).
-With the exception of shape, which is discrete, the remaining concepts take values in $[0,1]$.
-Table 2: Ablation on Quad
-Metric / Evaluation Context module Ablation 2
-Generated Sample Evaluation
-MMD (obs) ↓ 0.0099 (0.0058) 0.0187 (0.0063)
-MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
-MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
-Figure 3 shows OOD samples from a lightweight VAE trained on this dataset.
-Table 2 shows numerical results comparing our module (incorporated into a lightweight convolutional VAE) to Ablation 2 from above (analogously lightweight $\beta$-VAE trained on the full pooled dataset) for sample generation across the observational and interventional data (same distribution seen during training) along with OOD data.
-The quad data allows evaluating the sampled OOD distribution (which our module composes from seen contexts) with a ground truth OOD distribution (unseen during training).
-Figure 3: Example generated OOD images from a simple, lightweight VAE on the quad benchmark.
-This shows a proof-of-concept that the concept module successfully composes concepts using even simple architectures.
-For performant architectures, see the NVAE results in Figure 5.2.
-OOD generation vs. reconstruction
-There is a crucial difference between OOD reconstruction and OOD generation.
-By OOD generation, we mean the ability to conditionally generate and combine novel interventions.
-For example, suppose during training the model only sees small, red objects along with big, blue objects.
-At test time, we wish to generate OOD samples of big, red objects or small, blue objects.
-The crucial difference is that a model may be capable of OOD reconstruction, but not OOD generation.
-For example, we can always attempt to reconstruct an OOD sample and evaluate its reconstruction error.
-This is a common metric that has been used in previous work to evaluate the OOD capabilities of generative models (e.g.
-Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022).
-But unless the model learns specific latent concepts corresponding to size or colour, then we have no control over the size or colour of random samples from the model.
-We can always condition on the representations learned by the model, but these will not be disentangled or interpretable.
-Thus, OOD generation not only evaluates the ability of a model to compose learned concepts in new ways, but implicitly also the ability of a model to identify and capture underlying concepts of interest.
-Accordingly, we argue that OOD generation is a more appropriate evaluation metric.
-While the numerical results on synthetic data above indicate better performance on OOD generation using our module, a more intuitive visual evaluation on benchmark datasets is provided in Section 5.2, Figure 5.2.
-5.2 Large-scale models and benchmarks
-[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Figure 4: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right).
-In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together.
-The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
-We now evaluate performance when our module is incorporated into a complex black-box model (NVAE of Vahdat and Kautz (2020)) on more challenging datasets, again including MNIST but now also 3DIdent (Zimmermann et al., 2021) and CelebA (Liu et al., 2015).
-Since CelebA does not have interventional data, we use this as an important ablation to test how well our module performs when used with conditional data (i.e.
-by conditioning on attributes in CelebA).
-Table 3: Comparison of models (BPD = bits per dimension)
-Dataset Type Black-box BPD
-MNIST Observational NVAE 0.144
-MNIST Interventional NVAE 0.149
-3DIdent Observational NVAE 0.673
-3DIdent Interventional NVAE 0.754
-CelebA Observational NVAE 2.08
-CelebA Conditional NVAE 2.13
-Table 3 summarizes the results of these training runs, where “observational” means that the standard NVAE (i.e.
-without a context module) is trained on the entire datasets, whereas “interventional” means that the context module was used.
-We observe competitive reconstruction, with only slight degradation when the context module is attached to the NVAE decoder.
-More importantly, Figure 5.2 demonstrates generation capabilities enabled by our module.
-Each row corresponds to generated samples from the trained model in different contexts:
-  1. 1.
-    The first row shows generated observational samples;
-  2. 2.
-    The second and third row contain generated single-node interventions;
-  3. 3.
-    The final row shows generated samples where two distinct concepts are composed together.
-For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon.
-CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data.
-Although there is a slight drop in OOD perceptual accuracy on CelebA, this is to be expected since the model was trained on conditional data as opposed to interventions.
-Finally, observe the last row of Figure 5.2, which demonstrates the OOD generation capability enabled by our module: these contexts did not appear in the training data but our module is able to sensibly generate them by composing and generalizing from the concepts it learned according to different data contexts.
-Acknowledgements
+    Fine-tuning (Section 4.4): Can pre-trained models leverage the context module via fine-tuning to improve their learned representations and enable compositional generation?
+Instead of training end-to-end, we start with a pre-trained model and fine-tune it with the context module attached.
+This is evaluated on the concept learning, composition, and reconstruction tasks.
+For concept learning and composition, including OOD generation (Section 2), we report the sliced Wasserstein (SW) distance (Bonneel et al., 2015; Flamary et al., 2021) between a held-out sample from the ground-truth distribution and a sample generated (not reconstructed) by the trained model.
+For reconstruction, we report the standard ELBO loss (Kingma and Welling, 2014) in bits per dimension.
+Methods, ablations, baselines
+[…]
+Datasets
+We compare these models to our context module on four datasets: 3DIdent (Zimmermann et al., 2021), Morpho MNIST (Castro et al., 2019), and two versions of a new semi-synthetic environment called quad—one with independent concepts and another with (causally) dependent concepts.
+While 3DIdent and Morpho MNIST are established baselines, quad is a novel visual environment with intervenable latents (colour, shape, size, orientation) that explicitly enables sampling from ground-truth OOD composed contexts.
+See Appendix C for more details on this dataset.
+We introduce quad in order to carefully evaluate model performance in a controlled environment over different random seeds (and hence assess statistical variability).
+Thus, our evaluations use quad to assess reproducibility and statistical significance, whereas 3DIdent and MNIST are more appropriate for testing our approach on downstream tasks such as image generation.
+Due to space constraints, we report only a representative slice of the experiments here; full details and additional experiments can be found in Appendices CE.
+This includes detailed performance breakdowns across contexts; the effects of regularization and architectural capacity; and additional figures.
+4.1 Concept learning
+Table 1, in the Concept columns, shows a quantitative evaluation of concept learning on the quad, MNIST, and 3DIdent datasets.
+We draw three main conclusions from this table, by comparing our context module variants to each of the other methods (rows) in the ablations and baselines.
+First, comparing the three CM rows to the ablations in Table 1 (specifically, base and base pooled), we see that the context module excels at learning individual concepts across all datasets and models, despite the base models achieving slightly better average reconstruction performance (see Section 4.3).
+Moreover, Figure 2 shows that there is essentially no perceptual difference between samples from the base models and the context module.
+Additionally, across multiple training runs over different random seeds, our context module consistently outperforms the base models across quad concepts—in this case the simpler model/data allows more replicates to be run to evaluate statistical significance and standard errors.
+Second, comparing the CM rows to the base row of Table 1, we see that the context module outperforms the base model trained only on the observational (rather than pooled) data.
+This confirms that the context module is indeed successfully exploiting learned invariances across the different contexts, providing it with performance and computational resource advantages compared to models trained independently on the different contexts.
+Table 1: Context module (CM) vs.
+fine-tuning, ablations, and baselines on three tasks across four datasets.
+Best performance per task/dataset in bold.
+Base model indicated by $^{*}=$ lightweight VAE, $^{**}=$ NVAE.
+quad* (independent) quad* (dependent) MNIST** 3DIdent**
+Concept Compo Recon Concept Compo Recon Concept Compo Recon Concept Compo Recon
+ours CM (end-to-end) 0.063 0.081 0.519 0.116 0.127 0.828 0.035 - 0.135 0.014 0.051 0.528
+CM (fine-tune) 0.059 0.086 0.518 0.112 0.121 0.840 0.037 - 0.136 0.031 0.063 0.549
+CM (frozen) 0.086 0.132 0.559 0.121 0.134 0.889 0.052 - 0.161 0.048 0.063 0.609
+ablations base 0.175 0.288 0.444 0.170 0.222 0.767 0.102 - 0.271 0.056 0.092 4.966
+base pooled 0.166 0.248 0.446 0.167 0.212 0.746 0.105 - 0.139 0.049 0.056 0.559
+CM pooled 0.202 0.306 0.840 0.166 0.203 0.752 0.099 - 0.143 0.041 0.068 0.516
+baselines Ada-GVAE 0.335 0.355 2.017 0.375 0.377 2.443 0.500 - 1.035 0.249 0.251 3.267
+BetaTCVAE 0.330 0.346 1.973 0.366 0.374 2.408 0.484 - 1.036 0.257 0.254 3.261
+TVAE 0.334 0.341 2.048 0.370 0.368 2.455 0.486 - 1.037 0.254 0.256 3.282
+Third, comparing the CM rows to the ablation in CM pooled—which pools all the data into a single context in order to circumvent the context module’s intervention mechanism while leaving the rest of the context module architecture intact—we see that CM consistently outperforms this ablation.
+This demonstrates that the increased performance of the context module is due our module’s intervention mechanism and its resulting ability to better leverage context-separated data, as opposed to spuriously benefiting from a different architecture or increased complexity.
+Finally, for a qualitative evaluation, we look at Figure 2, in the Concept 1 and Concept 2 rows, to see concept learning in CM+NVAE across three different datasets.
+For example, looking at the 3DIdent results in Figure 2 (middle), the Observation row shows the learned observational distribution of images to contain background colours and object colours both in the orange–green range ( ).
+The Concept 1 row, produced by intervening on the learned concept of background colour, shows that the background colour is shifted to the blue–red range while other features, like object colour, remain invariant with the observational samples.
+Analogously, the Concept 2 row shows a shift of object colour to the blue–red range ( ) while the other features remain invariant.
+These illustrate that CM+NVAE successfully learned the concepts of background colour and object colour, in line with the notion of causal disentanglement in Definition 2.1.
+4.2 Concept composition
+Table 1, in the Compo columns, provides a quantitative evaluation of concept composition on the quad and 3DIdent datasets, while the Compo row of Figure 2 provides a qualitative evaluation across the MNIST and 3DIdent datasets.
+Since Morpho MNIST does not include ground truth compositional samples, we are not able to evaluate OOD generation on this dataset.
+In each evaluation, the model uses the intervention layer (Section 3.2) to intervene on pairs of learned concepts, i.e., to compose two concepts together.
+By holding out from training examples where concepts are composed, we are assured that the model has not seen any of these compositions.
+The only way to compose these concepts together is through the implicit causal model (via the reduced form SEM (3)) that is learned during training.
+We evaluate model performance with the SW metric (18), comparing concept compositions from the model against ground-truth hold out compositions.
+The same patterns in performance hold here for concept composition as we saw for concept learning in Section 4.1—namely, our context module consistently achieves better performance than baselines and ablations, demonstrating successful concept composition due to the module’s intervention mechanism.
+For example, looking again at the 3DIdent results in Figure 2 (middle), we see that the colour range of both the background and the object are shifted from the orange–green range ( ) to the blue–red range ( ).
+These experiments demonstrate the context module’s ability to learn causally disentangled representations through both concept learning and composition.
+(a)      (b)
+Figure 2: (a) Examples of concept learning and composition in MNIST (left) and 3DIdent (right) using CM (end-to-end), our context module augmenting the base NVAE—these images are generated, not reconstructed.
+The first row shows generated observational samples; The second and third rows show samples of learned concepts, generated by intervening on individual concepts, e.g., ‘Concept 1’ is ‘scaled’ for MNIST; The final row shows OOD composition, generated by simultaneously intervening on pairs of learned concepts.
+(b) Observational samples from the context module vs.
+the base model show no perceptual loss.
+4.3 Reconstruction
+The Recon columns of Table 1 compare the validation reconstruction performance of our approach against the alternatives.
+We draw two conclusions from these results.
+First, adding the context module incurs only a minor quantitative cost in reconstruction performance, while granting additional concept learning and composition abilities explored in Sections 4.1-4.2.
+Second, Figure 2 demonstrates that despite this small quantitative cost, the perceptual quality of the samples is still very high and indistinguishable from the base model.
+This validates the central motivation behind the context module, namely that it can imbue existing black-box models with additional concept learning and causal composition abilities without sacrificing perceptual metrics or downstream performance.
+4.4 Fine-tuning
+Finally, an important aspect of our method is that it need not be trained end-to-end but can instead be successfully fine-tuned from a frozen base mode.
+The second (CM (fine-tune)) and third (CM (frozen)) rows of Table 1 evaluate the context module on concept leaning and composition using two fine-tuning variants.
+Both variants start with a pre-trained model and then insert the context module and resume training, either only updating the context module weights in the frozen variant, or also fine-tuning the non-context module weights in the fine-tune variant.
+[…]
+5 Conclusion and Limitations
+Designing generative models with structured latent spaces that enable intervention, composition, and OOD generation is an outstanding challenge.
+Existing approaches either sacrifice structure for flexibility, or flexibility for structure.
+We argue that such a tradeoff is not necessary and that arbitrarily flexible models can indeed learn useful structure.
+To this end, we propose a simple context module that can be appended to a black-box decoder that learns causally disentangled latent spaces.
+The context module enables genuinely OOD, compositional generation without hardcoding prior knowledge or structure into the model, and it avoids the difficult problem of latent causal discovery.
+Our experiments provide both quantitative and qualitative evidence for this.
+While existing evaluations have focused on reconstruction as a metric to quantify OOD performance, we argue that OOD generation is in fact more appropriate since it also measures the ability of a model to learn useful concepts that can be manipulated at inference time.
+Since our module imposes no hard constraints, it may be useful in combination with other modules.
+Understanding how the context module interacts with other architectures is a natural next step and a promising direction for future work.
+We also encourage the community to explore alternative metrics for evaluating genuine OOD generalization.
+Finally, understanding the cost and tradeoffs of variational approximations compared to, say, exact maximum likelihood estimation is important, especially in causal applications.
+Indeed, this can make estimating the exact causal graph much more challenging.
+This motivates our fundamental design principle: to use a “causal enough” architecture to obtain composability and OOD generation while otherwise leveraging already empirically proven non-causal architectures.
+Acknowledgments
 The work by AM was supported by Novo Nordisk Foundation Grant NNF20OC0062897.
 This research was supported in part through the computational resources and staff contributions provided for the Mercury high performance computing cluster at The University of Chicago Booth School of Business which is supported by the Office of the Dean.
-Appendix A Identifiability of the proposed model
+Supplementary material:
+Intervening to learn and compose
+causally disentangled representations
+Appendix A Related work
+Our work is closely related to several parallel lines of work on structured generative models, causal representation learning, linear representations, and OOD generalization.
+Below we compare and contrast our contributions against this literature.
+Structured generative models
+To provide structure such as hierarchical, graphical, causal, and disentangled structures as well as other inductive biases in the latent space, there has been a trend towards building structured generative models that directly impose this structure a priori.
+Early work looked at incorporating fixed, known structure into generative models, such as autoregressive, graphical, and hierarchical structure (Germain et al., 2015; Johnson et al., 2016; Sønderby et al., 2016; Webb et al., 2018; Weilbach et al., 2020; Ding et al., 2021; Mouton and Kroon, 2023).
+This was later translated into known causal structure (Kocaoglu et al., 2017; Markham et al., 2023).
+When the latent structure is unknown, several techniques have been developed to learn useful (not necessarily causal) structure from data (Li et al., 2018; He et al., 2019; Wehenkel and Louppe, 2021; Kivva et al., 2022; Moran et al., 2022).
+More recently, based on growing interest in disentangled (Bengio, 2013) and/or causal (Schölkopf et al., 2021) representation learning, methods that learn causal structure have been developed (Markham and Grosse-Wentrup, 2020; Moraffah et al., 2020; Yang et al., 2021; Ashman et al., 2022; Shen et al., 2022; Kaltenpoth and Vreeken, 2023).
+In contrast to this line of work, which emphasizes hard graphical constraints, we neither learn nor impose any fixed graphical structure.
+Instead, we emphasize performance on concept learning and composition as downstream tasks.
+Causal disentanglement and representation learning
+Causal representation learning (Schölkopf et al., 2021) is a rapidly developing area that involves, among other goals, two key objectives: 1) Learning disentangled latent factors along with 2) Learning latent causal structure.
+The former is what we refer to as causal disentanglement (see Definition 2.1 and its causal interpretation).
+The latter problem has seen tremendous theoretical progress in recent years (Brehmer et al., 2022; Shen et al., 2022; Lachapelle et al., 2022; Moran et al., 2022; Kivva et al., 2022; Buchholz et al., 2023; Gresele et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023; Morioka and Hyvärinen, 2023; Liu et al., 2023; Mameche et al., 2023; Yao et al., 2023; Varici et al., 2023; Sturma et al., 2023; Xu et al., 2024; Zhang et al., 2024; Li et al., 2024; Varici et al., 2024; Bing et al., 2024; Ahuja et al., 2024; Talon et al., 2024; Yao et al., 2024).
+Recent work has also pushed in the direction of identifying concepts (Leemann et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025).
+Our work draws inspiration from these lines of work, which articulate precise conditions under which latent causal discovery is possible in principle.
+By contrast, our focus is on methodological aspects of causal disentanglement in practice.
+While the former is a notoriously difficult task, the latter only needs to exploit learned causal invariances: In particular, causal disentanglement is possible without learning latent causal structure.
+Linear representations
+Generative models are known to represent concepts linearly in embedding space (e.g.
+Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015); see also Levy and Goldberg (2014); Arora et al.
+(2016); Gittens et al.
+(2017); Allen and Hospedales (2019); Ethayarajh et al.
+(2018); Seonwoo et al.
+(2019).
+This phenomenon has been well-documented over the past decade in both language models (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Conneau et al., 2018; Tenney et al., 2019; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Li et al., 2023; Park et al., 2023; Gurnee et al., 2023; Cunningham et al., 2024; Jiang et al., 2024) and computer vision (Radford et al., 2015; Raghu et al., 2017; Bau et al., 2017; Engel et al., 2017; Kim et al., 2018; Chen et al., 2020; Wang et al., 2023; Trager et al., 2023).
+Our approach actively exploits this tendency by searching for concept representations as linear projections of the embeddings learned by a black-box model.
+In contrast to this prior work, our emphasis is on applying these empirical observations for causal disentanglement, compositional generalization, and OOD generation.
+OOD generalization
+A growing line of work studies the OOD generalization capabilities of generative models, with the general observation being that existing methods struggle to generalize OOD (Xu et al., 2022; Mo et al., 2024; Montero et al., 2022, 2021; Schott et al., 2022).
+It is worth noting that most if not all of this work evaluates OOD generalization using reconstruction on held-out OOD samples, as opposed to generation.
+For example, a traditional VAE may be able to reconstruct held-out samples, but it is not possible to actively sample OOD.
+See Section 2 for more discussion.
+Appendix B Identifiability of the proposed model
 In this appendix we set up and prove Theorem 4.1.
-A.1 Set-up
+B.1 Set-up
 Our setup is the following: Let $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n}),\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$ and $\mathbf{e}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{d_{e}})$ be random vectors taking values in $\mathcal{X},\mathcal{C}$ and $\mathcal{E}$, respectively.
@@
 that may not be perfectly represented by any particular embedding in $\mathbf{e}$.
-The linear representation hypothesis (Section 3) is an empirical observation that abstract concepts can be approximately represented as linear projections of a sufficiently flexible latent space.
+The linear representation hypothesis (Section A) is an empirical observation that abstract concepts can be approximately represented as linear projections of a sufficiently flexible latent space.
 This is operationalized by assuming that $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$ (see (A1) below).
 The matrix $C_{j}$ is referred to as the representation of the concept $\mathbf{c}_{j}$.
-If $\mathbf{c}_{j}\in\mathbb{R}$ (i.e.
-$C_{j}\in\mathbb{R}^{1\times d_{e}}$), then $\mathbf{c}_{j}$ is referred to as an atom and $C_{j}$ its atomic representation.
+If $\mathbf{c}_{j}\in\mathbb{R}$ (i.e., $C_{j}\in\mathbb{R}^{1\times d_{e}}$), then $\mathbf{c}_{j}$ is referred to as an atom and $C_{j}$ its atomic representation.
 We will often denote atoms by $\mathbf{a}_{j}$ for notational clarity.
@@
 Note that these assumptions are not necessary in practice—as evidenced by our experiments in Section 4 which clearly violate these assumptions—and are merely used here to provide a proof-of-concept identifiability result based on our architectural choices.
+The representation (5) implies, in particular, the usual factorization
+$p(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})=\prod_{j=1}^{d_{c}}p(\mathbf{c}_{j}\,|\,\operatorname{pa}_{\mathsf{G}}(j)),$
+which formalizes the notion of causal disentanglement as in Schölkopf et al. (2021).
 Our goal is to identify the concept marginal $p(\mathbf{c})$ and concept representations $(C_{1},\ldots,C_{d_{c}})$ from concept interventions.
 Concept interventions are well-defined through the structural causal model defined by the DAG $\mathsf{G}$ and its structural equations.
-We consider single-node concept interventions where $\mathbf{c}_{j}$ is stochastically set to be centered at $\boldsymbol{\mu}_{j}$ with variance $\Omega_{j}\succ 0$ (making this precise requires some set-up; see Appendix A.3.1 for details).
+We consider single-node concept interventions where $\mathbf{c}_{j}$ is stochastically set to be centered at $\boldsymbol{\mu}_{j}$ with variance $\Omega_{j}\succ 0$ (making this precise requires some set-up; see Appendix B.3.1 for details).
 By identifiability, we mean the usual notion of identifiability up to permutation, shift, and scaling from the literature.
@@
 As a result, the concept marginal $p(\mathbf{c})$ is also identifiable (again up to permutation, shift, and scale as well).
-The shift ambiguity can of course be resolved by assuming the concepts are centered (zero-mean), and the scale ambiguity can be resolved by assuming concepts are normalized (e.g.
-unit norm).
+The shift ambiguity can of course be resolved by assuming the concepts are centered (zero-mean), and the scale ambiguity can be resolved by assuming concepts are normalized (e.g., unit norm).
 We make the following assumptions, which are adapted from Rajendran et al.
@@
 (2024) do not consider a structural causal model over $\mathbf{c}$, and thus the notion of a concept intervention there is not well-defined.
-Assumption 1 (Atomic representations).
+Assumption 1 (Atomic representations)
 There exists a set of atomic representations $\mathcal{A}=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{n}\}$ of linearly independent vectors $\mathbf{a}_{j}\in\mathbb{R}^{d_{e}}$ such that the rows of each representation $C_{j}$ are in $\mathcal{A}$.
-We denote the indices of $\mathcal{A}$ that appear as rows of $C_{j}$ by $S^{j}$, i.e.
-$S^{j}=\{i:\mathbf{a}_{i}\text{ is a row in $C_{j}$}\}$.
-We assume that $\cup_{j}S^{j}=\mathcal{A}$, i.e.
-every atom in $\mathcal{A}$ appears in some concept representation $C_{j}$.
+We denote the indices of $\mathcal{A}$ that appear as rows of $C_{j}$ by $S^{j}$, i.e., $S^{j}=\{i\mathrel{\mathop{\ordinarycolon}}\mathbf{a}_{i}\text{ is a row in $C_{j}$}\}$.
+We assume that $\cup_{j}S^{j}=\mathcal{A}$, i.e., every atom in $\mathcal{A}$ appears in some concept representation $C_{j}$.
 Define a matrix $M\in\mathbb{R}^{n\times d_{c}}$ to track which concepts use which atoms, i.e.
@@
 where $\omega_{j}^{2}$ are the diagonal entries of $\Omega_{j}$. Similarly, we define a matrix $B\in\mathbb{R}^{n\times d_{c}}$ given by
-$B_{ij}=\begin{cases}\frac{(\boldsymbol{\mu}_{j})_{k}}{\omega_{j}^{2}}&\text{if $i\in S^{j}$ and the $k$th row of $C_{j}$ is $\mathbf{a}_{i}$,}\\ 0&\text{otherwise.}\end{cases}$ (29)
+[…]
 The second assumption ensures that the concepts and associated interventional environments are sufficiently diverse.
-Assumption 2 (Concept diversity).
+Assumption 2 (Concept diversity)
 For every pair of atoms $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$ with $i\neq j$ there is a concept $\mathbf{c}_{k}$ such that $i\in S^{k}$ and $j\notin S^{k}$.
 Furthermore, $\operatorname{rank}(M)=n$ and there exists $v\in\mathbb{R}^{d_{c}}$ such that $v^{\top}M=0$ and $(v^{\top}B)_{i}\neq 0$ for each coordinate $i$.
-A.2 Formal statement of Theorem 4.1
+B.2 Formal statement of Theorem 4.1
 With this setup, we can now state and prove a more formal version of Theorem 4.1.
-Theorem A.1.
-Under Assumptions 1-2 and given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$, in the sense of (26-27).
+Theorem B.1.
+Under Assumptions 1-2 and given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$, in the sense of (6-7).
 In the context of learning disentangled representations, identifying the concept marginal $p(\mathbf{c})$ is particularly relevant, since this implies we also learn certain (in)dependence relationships between the concepts.
-In other words, if the “true” concept representations are disentangled (i.e.
-independent), then we identify disentangled concepts.
+In other words, if the “true” concept representations are disentangled (i.e., independent), then we identify disentangled concepts.
 In relation to existing work, we make the following remarks:
   • While our proof ultimately relies on techniques from Rajendran et al.
-(2024), Theorem A.1 does not trivially follow from their results, which do not cover interventions over the concept marginal $p(\mathbf{c})$.
+(2024), Theorem B.1 does not trivially follow from their results, which do not cover interventions over the concept marginal $p(\mathbf{c})$.
 Extending their results to the case of concept interventions, including making sure these interventions are well-defined causal objects, is the main technical difficulty in our proof.
@@
   • Various results in CRL (e.g.
-Buchholz et al., 2023; von Kügelgen et al., 2023; Varici et al., 2024, see Section 3 for more references) prove identifiability results under interventions, however, these results require interventions directly on the embeddings $\mathbf{e}$, as opposed to the concepts $\mathbf{c}$.
-Theorem A.1, by contrast, requires only $d_{c}\ll d_{e}$ total interventions and in doing so directly address the technical challenges associated with intervening directly on concepts as opposed to embeddings.
+Buchholz et al., 2023; von Kügelgen et al., 2023; Varici et al., 2024, see Section A for more references) prove identifiability results under interventions, however, these results require interventions directly on the embeddings $\mathbf{e}$, as opposed to the concepts $\mathbf{c}$.
+Theorem B.1, by contrast, requires only $d_{c}\ll d_{e}$ total interventions and in doing so directly addresses the technical challenges associated with intervening directly on concepts as opposed to embeddings.
   • Leemann et al.
@@
 On the other hand, we have left finite-sample aspects to future work.
-A.3 Proof of Theorem 4.1
+B.3 Proof of Theorem 4.1
 To prove Theorem 4.1, we begin with some notation.
@@
 Finally, let $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ denote the collection of all marginal distributions over $\mathbf{c}$ induced by the distributions in $\mathcal{M}(\mathsf{G})$.
-A.3.1 Concept interventions
+B.3.1 Concept interventions
 In practice, the edge structure of the graph $\mathsf{G}$ is unknown, and we do not assume this is known in advance.
@@
 Although immediate from these definitions, we record this as a lemma for future use:
-Lemma A.2.
+Lemma B.2.
 If $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ then there is a distribution $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ such that $\widetilde{p}(\mathbf{c})=p(\mathbf{c})$ for all $\mathbf{c}\in\mathcal{C}$.
@@
 The following is obvious from our assumptions:
-Lemma A.3.
-For any distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, the interventional distribution $P^{j}$ defined by (31) is also an element of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, i.e.
-$P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
+Lemma B.3.
+For any distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, the interventional distribution $P^{j}$ defined by (11) is also an element of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, i.e., $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
 Thus, we have the usual, well-defined notion of intervention over the concepts $\mathbf{c}$. The next step is to translate the effect of these interventions onto $\mathbf{e}$ and $\mathbf{x}$.
 The following proposition shows that an interventional distribution $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ always extends to a well-defined post-interventional joint distribution over $(\mathbf{x},\mathbf{c},\mathbf{e})$ that can be interpreted as an interventional distribution for the appropriate choice of joint observational distribution over $(\mathbf{x},\mathbf{c},\mathbf{e})$.
-Proposition A.4.
+Proposition B.4.
 Let $P^{j}$ be an interventional distribution for $j$ and $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Then there exists a distribution $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
@@
     If $\widetilde{P}$ is induced by the functional relation $\mathbf{x}=f(\mathbf{e})$, and similarly $\widetilde{P}^{j}$ with $\mathbf{x}=f^{j}(\mathbf{e})$ then $f^{j}=f$.
-Proof.
-Lemma A.3 implies that $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Thus Lemma A.2 implies the existence of joint distributions $\widetilde{P},\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
-$\widetilde{p}(\mathbf{c})=p(\mathbf{c})\quad\text{ and }\quad\widetilde{p}^{j}(\mathbf{c})=p^{j}(\mathbf{c})\text{ for all }\mathbf{c}\in\mathcal{C}.$ (32)
-(1) follows from the definition of $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ for $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ from Lemma A.2 above; i.e. (32) above.
+Proof B.5.
+Lemma B.3 implies that $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Thus Lemma B.2 implies the existence of joint distributions $\widetilde{P},\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
+[…]
+(1) follows from the definition of $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ for $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ from Lemma B.2 above; i.e., (12) above.
 (2) follows from the definition of an interventional distribution and the factorization
@@
 satisfied by any distribution $P\in\mathcal{M}(\mathsf{G})$. Namely, the interventional distribution $P^{j}$ satisfies
-$\begin{split}\widetilde{p}^{j}(\mathbf{x},\mathbf{c},\mathbf{e})&=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ &=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e}){p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ \end{split}$
+[…]
 according to the definition of $\widetilde{P}$.
 Since the intervention only perturbs the conditional factors $\widetilde{p}(\mathbf{c}_{j}\mid\mathbf{c}_{\textrm{pa}_{\mathsf{G}}(j)},\mathbf{e})$ in $\widetilde{p}(\mathbf{c}\mid\mathbf{e})$ for $j$, we can always choose $\widetilde{P}$ and $\widetilde{P}^{j}$ such that $\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})$.
-(3) follows similarly to (2). Namely, since the intervention only perturbs the conditional factors specified above, we can always choose $f^{j}=f$. ∎
+(3) follows similarly to (2). Namely, since the intervention only perturbs the conditional factors specified above, we can always choose $f^{j}=f$.
 We additionally have the following:
-Proposition A.5.
-Let $I_{j}:=\{j\}$ be the collection of single-node intervention targets, and consider the interventional distributions $P^{1},\ldots,P^{d_{c}}$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
+Proposition B.6.
+Let $I_{j}\mathrel{\mathop{\ordinarycolon}}=\{j\}$ be the collection of single-node intervention targets, and consider the interventional distributions $P^{1},\ldots,P^{d_{c}}$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
 Then there exists $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}\in\mathcal{M}(\mathsf{G})$ such that
@@
 and $f^{j}=f$ for all $j$.
-Proof.
+Proof B.7.
 Fix a choice of $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ for $P$, which specifies a function $f$.
-For the chosen $\widetilde{P}$, we construct $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}$ according to Proposition A.4.
-The result follows by applying Proposition A.4 for every $j$.
-∎
-Proposition A.5 will allow us (below) to define more formally the environments that are used to identify concepts.
+For the chosen $\widetilde{P}$, we construct $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}$ according to Proposition B.4.
+The result follows by applying Proposition B.4 for every $j$.
+Proposition B.6 will allow us (below) to define more formally the environments that are used to identify concepts.
 Intuitively, each environment corresponds to single-node interventions on a single concept $\mathbf{c}_{j}$, however, we can only observe the effect this intervention has on the observed $\mathbf{x}$.
-The purpose of Propositions A.4-A.5 are to trace the dependence between $\mathbf{x}$ and $\mathbf{c}$—via the embeddings $\mathbf{e}$—according to the model (A1)-(A3).
-Formally, let $I_{j}=\{j\}$ with $\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2})$ and $\mathbf{c}_{j}=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime}$ according to (31), i.e.
-each environment corresponds to a single-node intervention on the $j$th concept $\mathbf{c}_{j}$.
-By Proposition A.5, there exist post-intervention joint distributions $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ such that
+The purpose of Propositions B.4-B.6 are to trace the dependence between $\mathbf{x}$ and $\mathbf{c}$—via the embeddings $\mathbf{e}$—according to the model (A1)-(A3).
+Formally, let $I_{j}=\{j\}$ with $\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2})$ and $\mathbf{c}_{j}=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime}$ according to (11), i.e., each environment corresponds to a single-node intervention on the $j$th concept $\mathbf{c}_{j}$.
+By Proposition B.6, there exist post-intervention joint distributions $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ such that
 $\widetilde{p}^{j}(\mathbf{e})=\widetilde{p}(\mathbf{e})\qquad\textrm{ for all $j\in[d_{c}]$}$
 and $f^{j}=f$ for all $j$.
-Then the $j$th environment corresponds to sampling $\mathbf{x}=f(\mathbf{e})$ where $\mathbf{e}\sim p_{j}(\mathbf{e}):=\widetilde{p}^{j}(\mathbf{e}\,|\,\mathbf{c}_{j}=\boldsymbol{\mu}_{j})$.
-A.3.2 Reduction to concept identification
+Then the $j$th environment corresponds to sampling $\mathbf{x}=f(\mathbf{e})$ where $\mathbf{e}\sim p_{j}(\mathbf{e})\mathrel{\mathop{\ordinarycolon}}=\widetilde{p}^{j}(\mathbf{e}\,|\,\mathbf{c}_{j}=\boldsymbol{\mu}_{j})$.
+B.3.2 Reduction to concept identification
 We begin by computing the log-likelihood ratio between the observational and interventional environments.
 The idea is that the concept representations are revealed through fluctuations in this likelihood ratio.
-Using Proposition A.5, we have
+Using Proposition B.6, we have
 $\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\log p_{0}(\mathbf{e})-\log\widetilde{p}^{j}(\mathbf{c}_{j}=\boldsymbol{\mu}_{j}\,|\,\mathbf{e})-\log\widetilde{p}^{j}(\mathbf{e})$ (13)
@@
 After this change of coordinates, the system can be rewritten as
-$h_{j}(\mathbf{e}):=\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}\mathbf{e}_{i}^{2}-B_{ij}\mathbf{e}_{i}+O(1).$ (36)
+$h_{j}(\mathbf{e})\mathrel{\mathop{\ordinarycolon}}=\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}\mathbf{e}_{i}^{2}-B_{ij}\mathbf{e}_{i}+O(1).$ (16)
 These functions are convex, which allows us to identify the $|S_{T}|$ where $S_{T}=\cup_{j\in T}S_{j}$ by minimizing the convex function $\sum_{j\in T}h_{j}(\mathbf{e})$.
@@
 Let $\mathbf{h}=(h_{1},\ldots,h_{n})$ and write $\mathbf{e}^{c}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{n})$ and $\mathbf{e}=(\mathbf{e}^{c},\mathbf{e}^{\perp})\in\mathbb{R}^{n\times(d_{c}-n)}$.
-Let $\iota^{\perp}(\mathbf{e}^{c})=(\mathbf{e}^{c},0)$, $\pi^{c}(\mathbf{e})=\mathbf{e}^{c}$, and $\varphi^{\perp}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\varphi^{\perp}(\mathbf{e}^{c})_{i}=\varphi(\mathbf{e},0)_{i}$.
+Let $\iota^{\perp}(\mathbf{e}^{c})=(\mathbf{e}^{c},0)$, $\pi^{c}(\mathbf{e})=\mathbf{e}^{c}$, and $\varphi^{\perp}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\varphi^{\perp}(\mathbf{e}^{c})_{i}=\varphi(\mathbf{e},0)_{i}$.
 Note that via the change of variables formula, we can write $\mathbf{h}(\mathbf{e})=\mathbf{H}(\mathbf{x})$ for some function $\mathbf{H}$.
@@
 From here, identifiability follows from a straightforward but lengthy linear algebraic argument based on the first-order conditions for minimizing $\mathbf{h}$ and the identifiability of $M,B$ proved above.
-The proof of Theorem A.1 follows.
-Appendix B quad: A semi-synthetic benchmark for OOD composition
+The proof of Theorem B.1 follows.
+Appendix C quad: A semi-synthetic benchmark for compositional generation
 To provide a controllable, synthetic test bed for evaluating composition in a visual environment, we developed a simple semi-synthetic benchmark, visualized in Figures 34.
@@
   • This can be simulated on any $n\times n$ grid (as long as $n$ is large enough to distinguish different shapes; $n\geq 16$ is large enough in practice). In our experiments, we used $n=64$.
-These properties make quad especially useful for both quantitative and qualitative evaluation; indeed, it was created precisely for this reason. Example images are shown in Figure 5.
-To describe our experimental setting, let $\mathbf{c}_{j}$ indicate the $j$th concept as listed above.
-We used seven different contexts, including an observational context and one context each for single-concept interventions on $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$, i.e.
-quad1, quad2, quad3, quad4, size, and orientation.
+These properties make quad especially useful for both quantitative and qualitative evaluation; indeed, it was created precisely for this reason. Example images are shown in Figures 3-4.
+To describe the experimental setting used in the quad (independent) experiments, let $\mathbf{c}_{j}$ indicate the $j$th concept as listed above.
+We used seven different contexts, including an observational context and one context each for single-concept interventions on $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$, i.e., quad1, quad2, quad3, quad4, size, and orientation.
 We use a sample size of 50 000 images per context.
@@
 Similarly, we can compose multiple interventions together, as in Figure 4.
-This facilitates evaluation of the concept composition capabilities of models trained on single interventions, providing a ground truth distribution to compare against generated OOD samples like in Figure 7.
+This facilitates evaluation of the concept composition capabilities of models trained on single interventions, providing a ground truth distribution to compare against generated OOD samples like in Figure LABEL:fig:quad-ood-full.
 Figure 3: Example images obtained from the quad dataset.
 Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention).
+For example, in quad1, the first (top left) quadrant has been manipulated from orange-green hues to blue-red hues.
 We also made use of the easy-to-generate and perceptually distinct concept compositions in this dateset by evaluating models on held-out contexts (each containing 10 000 images) with double-concept interventions, including quad2_quad3, quad2_quad4, quad2_size, quad1_quad2, quad1_quad3, quad1_quad4, and quad1_size.
 Examples can be seen in Figure 4.
-For comparison, see Figure D.1.2 for examples of generated multi-concept compositions (OOD) from a simple 3-layer convolutional model. The training data did not contain any multi-concept compositions.
-Figure 6: Example images obtained from the quad dataset. Each subfigure corresponds to a different double-concept intervention context, indicated by the title.
-Appendix C Experimental details
-In this appendix we provide additional details on our experimental protocol and set-up.
-C.1 Evaluation metrics
+For comparison, see Figure E.1.2 for examples of generated multi-concept compositions (OOD) from a simple 3-layer convolutional model. The training data did not contain any multi-concept compositions.
+Figure 4: Example images obtained from the quad dataset.
+Each subfigure corresponds to a different double-concept intervention context, indicated by the title.
+For example, in quad2_quad3, the second (top-right) and third (bottom-left) quadrants have been manipulated from orange-green hues to blue-red hues.
+For quad dependent, a random SEM is generated using the sempler package (Gamella et al., 2022), which also allows for data generation under shift interventions.
+Appendix D Experimental details
+In this appendix we provide additional details on our experimental protocol and set-up. We have used Snakemake (Mölder et al., 2025) for reproducibility.
+D.1 Evaluation metrics
 We evaluated each model on the following metrics:
-  • Validation ELBO loss: This is the standard ELBO loss used for both training and validating VAEs.
+  • Validation ELBO loss: This is the standard ELBO loss used for validating VAEs.
 We report this in bits per dimension (bpd), which is the loss scaled by a factor of $\frac{1}{\mathrm{num\_pixels}\cdot\ln 2}$.
-  • Reconstruction loss: This approximates the negative log-likelihood. We use binary cross-entropy and report bpd.
-  • Covariance error: Unlike the losses above, which evaluate individual reconstructed images with respect to their corresponding ground truth images, this metric (and the ones below) evaluates a reconstructed distribution with respect to a ground truth distribution.
-Specifically, denoting the empirical covariance matrix across the ground truth datasets as $S^{T}$ and that over the reconstructed dataset as $S^{R}$, we report $\frac{1}{d^{2}}\sum_{i,j=1}^{d}|S^{T}_{ij}-S^{R}_{ij}|$.
-  • Mean error: This is analogous to the previous metric but using sample means instead of covariances:$\frac{1}{d}\sum_{i=1}^{d}|\mu^{T}_{i}-\mu^{R}_{i}|$.
-  • Maximum mean discrepancy (mmd): Another metric comparing distributions, MMD (Gretton et al., 2012) is a kernel-based metric for comparing two general distributions that may not be distinguishable by merely mean and covariance.
-When reporting values for these metrics in tables, we show the mean, averaged over different runs/seeds, followed in parentheses by the standard error.
-We generally use a 70/30 training/validation split.
-When appropriate, evaluations are split up so that metrics are computed with respect to the observational context (obs), or averaged over single-concept interventional contexts (ivn), or averaged over double-concept interventional contexts (ood)—with the exception of Table 3, which is detailed in Section 5.2 and Appendix C.4.
-C.2 Datasets
-In addition to the quad benchmark introduced in Appendix B, we used the following benchmark datasets.
+  • $p$-sliced Wasserstein (SW) distance: Unlike the loss above, which evaluates individual reconstructed images with respect to their corresponding ground truth images, this metric (Bonneel et al., 2015) evaluates a generated distribution $\mu_{\textrm{gen}}$ with respect to a ground truth distribution $\mu$.
+Specifically, we use the implementation of Flamary et al.
+(2021) to compute a Monte-Carlo approximation of $\mathcal{SWD}_{p}(\mu_{\textrm{gen}},\mu)$, where
+    $\mathcal{SWD}_{p}(\mu,\nu)=\underset{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}{\mathbb{E}}\left(\mathcal{W}_{p}^{p}(\theta_{\#}\mu,\theta_{\#}\nu)\right)^{\frac{1}{p}},$ (18)
+    and $\theta_{\#}\mu$ stands for the pushforwards of the projection $X\in\mathbb{R}^{d}\mapsto\langle\theta,X\rangle$ and $p=2$.
+When reporting values for these metrics in tables, we show the mean, averaged over different runs/seeds, followed by the standard error. We generally use a 70/30 training/validation split.
+D.2 Datasets
+In addition to the quad benchmark introduced in Appendix C, we used the following benchmark datasets.
+3DIdent
+The 3DIdent dataset (Zimmermann et al., 2021) consists of a 3D object, rendered using the Blender rendering engine, under different lighting conditions and orientations.
+We generated 50 000 samples per context corresponding to single-node interventions on background colour, spotlight colour, and object colour, in addition to the observational context.
 MNIST
@@
 The datasets contain 60 000 images per context.
-3DIdent
-The 3DIdent dataset (Zimmermann et al., 2021) consists of a 3D object, rendered using the Blender rendering engine, under different lighting conditions and orientations.
-We generated 50 000 samples per context corresponding to single-node interventions on background colour, spotlight colour, and object colour, in addition to the observational context.
-CelebA
-We use the original CelebA data (Liu et al., 2015), using the attributes included in the original dataset to define different contexts.
-Since we do not have access to interventions on this dataset, we relied on conditioning to create different contexts for different concepts.
-To minimize leakage between contexts, we chose 6 attributes by minimizing the between-attribute correlations and also choosing attributes that are easily detected visually (e.g.
-black or blond hair).
-This resulted in six total contexts of varying sizes, one observational along with single-node interventions on bags_under_eyes (41 446 samples), black_hair (48 472 samples), blond_hair (29 983 samples), male (84 434 samples), and mouth_slightly_open (97 942 samples).
-C.3 Architectures
-We tested our context module using two black-box models:
+D.3 Architectures
+We tested our context module end-to-end using two black-box models:
   1. 1.
-    lightweight-VAE: A standard 3-layer convolutional network with latent dimension $\dim(\mathbf{z})=128$. This is used for the MNIST and quad experiments in Section 5.1.
+    (lightweight) $\beta$-VAE: A standard 3-layer convolutional network with latent dimension $\dim(\mathbf{z})=128$.
+This is used for the quad experiments in Section 4 and the quad and MNIST experiments in Section E.1.
   2. 2.
-    NVAE: A deep hierarchical VAE; see Vahdat and Kautz (2020) for details. This is used for the 3DIdent and CelebA experiments, as well as some MNIST experiments, in Section 5.2.
-The lightweight-VAE was used for comprehensive ablations and experiments on simple datasets (150 total models overall, 5 per column per table in Appendix D.1).
-Since NVAE requires substantially more resources and time to train, this was used to train 6 total models, 2 per dataset.
-To provide an apples-to-apples baseline, we trained MNIST using both lightweight-VAE and NVAE.
-C.4 Discussion of experimental results (Section 5 in main paper)
-Table 1 and Table 2 are interpreted directly as described in Appendix C.1. However, Table 3 warrants some clarification.
-Figures D.2.1-D.2.4 in Appendix D.2, show the results of NVAE+context-module across three different datasets (MNIST, 3DIdent, and CelebA).
-All images here are generated by the model: that is, we draw a sample from the prior and pass it through the context module and decoder—this is to be contrasted with reconstruction, in which a reference image is passed through the encoder to obtain a sample from the posterior that is then passed through the (context module and) decoder.
-The top row is the result of asking the model to generate images from the observational context (in which no concepts are intervened or conditioned on), while the second and third rows contain images generated from respective contexts in which one concept has been intervened on (in the case of MNIST and 3DIdent) or conditioned on (in CelebA).
-These first three rows correspond to contexts the model has seen during training (though this is a generation and not reconstruction task).
-In contrast, the bottom row contains images resulting from asking the model to compose the concepts it has learned (specifically, the two concepts that were individually intervened/conditioned on in the second and third rows)—importantly, notice that the model has not seen data from this double-concept intervention (in the case of MNIST and 3DIdent).
-Table 3 compares two different models (namely, NVAE versus NVAE+context-module) across the three datasets.
-The different models are indicated by the Type column, with “Observational” denoting the plain NVAE model trained on a dataset in which all contexts are pooled, while “Interventional” and “Conditional” denote NVAE+context-module trained on unpooled datasets containing different contexts.
-The BPD column reports the reconstruction loss (described in Appendix C.1) in standard units so that it is directly comparable with the results of Vahdat and Kautz (2020) and other papers.
-C.5 Additional ablations
+    NVAE: A deep hierarchical VAE; see Vahdat and Kautz (2020) for details. This is used for the MNIST and 3DIdent experiments in Sections 4 and E.2.
+The lightweight-VAE was used for comprehensive ablations and experiments on simple datasets, totaling more than 350 models overall, with 5 per column per table in Appendix E.1.
+Since NVAE requires substantially more resources and time to train, this was used to train 12 total models, shown in Table 1, the three CM variations plus three ablations on both MNIST and 3DIdent.
+D.4 Additional ablations
 We conducted additional ablations to stress test our model and isolate the effect of the context module. This included additional experiments on different models, regularization, and expressivity.
 Models
-On MNIST and quad, we ran ablation studies on lightweight-VAE+context-module (denoted “Context module” in tables)—this is our module attached to the black-box, making full use of the different contexts.
+We ran ablations over different base VAEs augmented with the context module, denoted CM (end-to-end)—this is our module attached to the black-box, making full use of the different contexts, trained end-to-end.
 To evaluate the proposed module, we compared its performance to three natural baselines:
   1. 1.
-    Ablation 1: lightweight-VAE (without the context module) trained only on the observational context—this is the black-box when denied all (non-observational) context information and data.
+    CM pooled: CM (end-to-end) but trained with a single context on the full pooled dataset—this is our module attached to the black-box given all the data but not making use of context-specific information, providing an ablation that circumvents our intervention layer (Section 3.2) while maintaining the same architectural complexity.
   2. 2.
-    Ablation 2: lightweight-VAE (without the context module) trained on the full pooled dataset—this is the black-box when given all data but no explicit context information.
+    base: Base VAE (without the context module) trained only on a single observational context—this is the black-box when denied all (non-observational) context information and data.
   3. 3.
-    Ablation 3 is lightweight-VAE+context-module trained on the full pooled dataset—this is our module attached to the black-box given all the data but not making use of context-specific information.
-Note that due to differences in the size of the pooled vs.
-observational-only datasets, loss values from Ablation 1 are not directly comparable to Ablations 2 or 3 (or the context module), but are included anyway for reference.
-Results are shown in Section 5.1 and Appendix D.1.1.
+    base pooled: Base VAE (without the context module) trained with a single context on the full pooled dataset—this is the black-box when given all data but no explicit context information.
+[…]
 Regularization
-We run lightweight-VAE+context-module on MNIST, independently trying out two different regularizers, varying the regularization weight $\lambda$ for each: group lasso and $\ell_{2}$.
-These are applied to the weights of the reduced-form SEM $A_{0}$ in Equation 1.
+We run CM (end-to-end) with $\beta$-VAE base on MNIST, independently trying out two different regularizers, varying the regularization weight $\lambda$ for each: group lasso and $\ell_{2}$.
+These are applied to the weights of the reduced-form SEM $A_{0}$ in 2.
 We also try varying $\beta$ (as in $\beta$-VAE).
-Results are shown in Appendix D.1.3.
+Results are shown in Appendix E.1.3.
 Expressivity
-We run lightweight-VAE+context-module on MNIST, jointly varying three parameters that control expressiveness of our context module: expressive input width $0pt_{\mathrm{exp}}$ (which controls latent dimension $\dim(\mathbf{z})$), depth of the expressive layer $h_{\mathrm{exp}}$, and concept width ($0pt_{c}=\dim(\mathbf{c}_{j})=0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ as in Section 4.2).
-Results are shown in Appendix D.1.4.
-Appendix D Additional results
-In this section, we present all additional results, including experimental settings described in Appendix C for our context module with the lightweight-VAE black-box architecture as well as additional results on MNIST, 3DIdent, and CelebA for our context module with NVAE.
-D.1 lightweight-VAE
-D.1.1 Ablations (MNIST and quad)
-A description of the architecture and experiment is given in Appendix C.
-Table 4 shows ablation results after 200 epochs of training on MNIST, and Table 5 is after 1000 epochs.
-[^1] All other experiments in Appendix D.1 are accordingly run with no ablation and on 200 epochs.
-These are comparable, with Table 5 indicating some overfitting.
-Table 6 shows ablation results on quad after 200 epochs.
-Altogether, these ablations studies suggest that the cost of the structured representation learned by the context module is only a slight degradation in some of the numerical metrics, while Figure 7 shows sensible OOD concept composition using our module.
-Table 4: Ablation on MNIST (200 epochs)
-Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
-Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
-Reconstructed Distribution Evaluation
-Covariance Error ↓ 1.4243 (0.1312) 0.6468 (0.0135) 1.1859 (0.0576) 1.3886 (0.1374)
-Mean Error ↓ 0.0218 (0.0033) 0.0052 (0.0005) 0.0204 (0.0018) 0.0237 (0.0053)
-MMD ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
-Covariance Error (ivn) ↓ 2.3717 (0.2764) 2.1983 (0.2567) 2.0005 (0.2380) 2.4591 (0.2833)
-Mean Error (ivn) ↓ 0.0332 (0.0030) 0.0221 (0.0019) 0.0258 (0.0030) 0.0339 (0.0039)
-MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
-Generated Distribution Evaluation
-MMD ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
-MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
-Table 5: Ablation on MNIST (1000 epochs)
-Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.3150 (0.0085) 0.1889 (0.0020) 0.2393 (0.0087) 0.3145 (0.0116)
-Reconstruction Loss (bpd) ↓ 0.2836 (0.0087) 0.1513 (0.0025) 0.2010 (0.0090) 0.2837 (0.0112)
-Reconstructed Distribution Evaluation
-Covariance Error ↓ 1.3057 (0.0635) 0.5542 (0.0061) 1.1227 (0.0940) 1.5646 (0.2203)
-Mean Error ↓ 0.0154 (0.0023) 0.0040 (0.0001) 0.0194 (0.0038) 0.0274 (0.0096)
-MMD ↓ 0.0089 (0.0010) 0.0000 (0.0000) 0.0022 (0.0014) 0.0172 (0.0040)
-Covariance Error (ivn) ↓ 2.2713 (0.2320) 2.2229 (0.2536) 1.9782 (0.2247) 2.6567 (0.2901)
-Mean Error (ivn) ↓ 0.0314 (0.0023) 0.0221 (0.0018) 0.0263 (0.0030) 0.0392 (0.0046)
-MMD (ivn) ↓ 0.0932 (0.0340) 0.0298 (0.0108) 0.0388 (0.0132) 0.0610 (0.0177)
-Generated Distribution Evaluation
-MMD ↓ 0.0045 (0.0004) 0.0028 (0.0004) 0.0055 (0.0011) 0.0179 (0.0129)
-MMD (ivn) ↓ 0.0211 (0.0058) 0.0156 (0.0037) 0.0159 (0.0036) 0.0231 (0.0034)
-Table 6: Ablation on quad
-Metric / Evaluation Context module Ablation 2
-Validation Metrics
-Validation ELBO (bpd) ↓ 5.0031 (0.7180) 3.9357 (0.8504)
-Reconstruction Loss (bpd) ↓ 4.9586 (0.7161) 3.8795 (0.8506)
-Reconstructed Distribution Evaluation (Observations)
-MAE_Cov ↓ 14.5117 (0.5123) 6.0099 (1.3163)
-MAE_Mean ↓ 0.0176 (0.0009) 0.0058 (0.0015)
-MMD ↓ 0.0000 (0.0000) 0.0000 (0.0000)
-Reconstructed Distribution Evaluation (Others)
-MAE_Cov ↓ 25.3612 (6.8679) 15.8097 (2.4898)
-MAE_Mean ↓ 0.0295 (0.0055) 0.0191 (0.0036)
-MMD ↓ 0.0060 (0.0030) 0.0022 (0.0015)
-Out-of-Distribution (OOD) Evaluation (Averaged)
-MAE_Cov ↓ 43.5994 (2.8442) 36.3404 (3.0105)
-MAE_Mean ↓ 0.0570 (0.0036) 0.0440 (0.0037)
-MMD ↓ 0.0220 (0.0031) 0.0103 (0.0018)
-Generated Distribution Evaluation
-MMD ↓ 0.0099 (0.0058) 0.0187 (0.0063)
-MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
-MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
-D.1.2 Examples of concept composition (quad)
+We run CM (end-to-end) on MNIST, jointly varying three parameters that control expressiveness of our context module: expressive input width $0pt_{\mathrm{exp}}$ (which controls latent dimension $\dim(\mathbf{z})$), depth of the expressive layer $h_{\mathrm{exp}}$, and concept width ($0pt_{c}=\dim(\mathbf{c}_{j})=0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ as in Section 3.2).
+Results are shown in Appendix E.1.5.
+D.5 Additional baselines
+In addition to the above base models and ablations, we include comparisons against three disentanglement methods, including the unsupervised $\beta$-TCVAE (Chen et al., 2018), the weakly supervised Ada-GVAE (Locatello et al., 2020), and the supervised TVAE (Ishfaq et al., 2018).
+We gratefully acknowledge the open source implementation of these within the disent Python package (Michlo, 2021).
+In addition to comparisons against the quad ablations, these baselines are run on MNIST and 3DIdent in Appendix E.1.1.
+Appendix E Additional results
+In this section, we present complete results, including experimental settings described in Appendix D for our context module with the lightweight-VAE black-box architecture as well as additional results on MNIST and 3DIdent for our context module with NVAE.
+E.1 lightweight-VAE
+E.1.1 Ablations (MNIST and quad)
+A description of the architecture and experiment is given in Appendix D.
+Tables 25 are already summarized in the main text in Table 1, but here we use the extra space here to additionally present standard errors and show a breakdown of the results over the different contexts.
+Table 2: Evaluation of the context module, ablations, and baselines in terms of reconstruction and concept learning on MNIST.
+CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
+reconstruction (ELBO BPD)
+in-distribution 0.259 ± 0.005 0.235 ± 0.004 0.231 ± 0.003 0.181 ± 0.000 0.214 ± 0.003 0.228 ± 0.002 1.035 ± 0.002 1.036 ± 0.012 1.037 ± 0.001
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.099 ± 0.006 0.085 ± 0.002 0.082 ± 0.001 0.067 ± 0.002 0.085 ± 0.003 0.082 ± 0.003 0.521 ± 0.009 0.502 ± 0.014 0.463 ± 0.020
+scaled 0.062 ± 0.002 0.100 ± 0.007 0.109 ± 0.011 0.179 ± 0.006 0.142 ± 0.010 0.165 ± 0.008 0.526 ± 0.017 0.492 ± 0.012 0.490 ± 0.016
+shear 0.119 ± 0.006 0.110 ± 0.002 0.110 ± 0.003 0.105 ± 0.002 0.109 ± 0.002 0.109 ± 0.003 0.510 ± 0.013 0.476 ± 0.012 0.488 ± 0.024
+shift 0.113 ± 0.005 0.112 ± 0.003 0.105 ± 0.003 0.110 ± 0.004 0.115 ± 0.002 0.116 ± 0.004 0.494 ± 0.014 0.447 ± 0.012 0.501 ± 0.009
+swel 0.126 ± 0.007 0.113 ± 0.005 0.115 ± 0.002 0.102 ± 0.003 0.129 ± 0.009 0.114 ± 0.003 0.473 ± 0.011 0.492 ± 0.007 0.481 ± 0.011
+thic 0.178 ± 0.007 0.184 ± 0.002 0.193 ± 0.010 0.196 ± 0.008 0.231 ± 0.014 0.207 ± 0.014 0.474 ± 0.008 0.475 ± 0.012 0.472 ± 0.013
+thin 0.064 ± 0.004 0.080 ± 0.006 0.092 ± 0.010 0.133 ± 0.011 0.107 ± 0.006 0.126 ± 0.007 0.501 ± 0.014 0.506 ± 0.013 0.504 ± 0.011
+concept learning mean 0.109 ± 0.006 0.112 ± 0.004 0.115 ± 0.006 0.127 ± 0.005 0.131 ± 0.007 0.131 ± 0.006 0.500 ± 0.012 0.484 ± 0.012 0.486 ± 0.015
+Table 3: Evaluation of the context module, ablations, and baselines in terms of reconstruction, concept learning, and composition on quad (independent).
+CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
+reconstruction (ELBO BPD)
+in-distribution 0.519 ± 0.022 0.518 ± 0.006 0.559 ± 0.011 0.444 ± 0.000 0.446 ± 0.002 0.840 ± 0.072 2.017 ± 0.006 1.973 ± 0.001 2.048 ± 0.008
+out-of-distribution 0.720 ± 0.069 0.675 ± 0.039 0.774 ± 0.039 3.095 ± 0.295 0.487 ± 0.009 0.830 ± 0.042 2.092 ± 0.007 1.980 ± 0.003 2.115 ± 0.019
+concept learning (sliced Wasserstein)
+obs 0.056 ± 0.003 0.051 ± 0.004 0.078 ± 0.005 0.061 ± 0.004 0.100 ± 0.007 0.110 ± 0.013 0.332 ± 0.007 0.317 ± 0.009 0.349 ± 0.007
+quad1 0.062 ± 0.003 0.056 ± 0.005 0.075 ± 0.003 0.233 ± 0.004 0.201 ± 0.014 0.266 ± 0.020 0.342 ± 0.009 0.335 ± 0.012 0.324 ± 0.013
+quad2 0.061 ± 0.005 0.058 ± 0.007 0.086 ± 0.009 0.234 ± 0.014 0.196 ± 0.012 0.254 ± 0.017 0.346 ± 0.007 0.323 ± 0.006 0.334 ± 0.010
+quad3 0.063 ± 0.002 0.064 ± 0.005 0.085 ± 0.006 0.251 ± 0.013 0.210 ± 0.008 0.244 ± 0.020 0.333 ± 0.013 0.343 ± 0.010 0.334 ± 0.011
+quad4 0.065 ± 0.007 0.057 ± 0.004 0.100 ± 0.011 0.227 ± 0.010 0.227 ± 0.005 0.246 ± 0.011 0.347 ± 0.016 0.334 ± 0.011 0.324 ± 0.016
+size 0.075 ± 0.007 0.073 ± 0.006 0.101 ± 0.008 0.157 ± 0.009 0.122 ± 0.001 0.177 ± 0.020 0.327 ± 0.004 0.307 ± 0.010 0.336 ± 0.009
+orientation 0.057 ± 0.003 0.053 ± 0.003 0.077 ± 0.006 0.060 ± 0.003 0.104 ± 0.009 0.114 ± 0.016 0.316 ± 0.011 0.349 ± 0.013 0.334 ± 0.010
+concept learning mean 0.063 ± 0.004 0.059 ± 0.005 0.086 ± 0.007 0.175 ± 0.008 0.166 ± 0.008 0.202 ± 0.017 0.335 ± 0.010 0.330 ± 0.010 0.334 ± 0.011
+composition (sliced Wasserstein)
+(quad1, quad2) 0.100 ± 0.028 0.108 ± 0.016 0.137 ± 0.014 0.311 ± 0.009 0.276 ± 0.014 0.343 ± 0.018 0.346 ± 0.012 0.346 ± 0.004 0.334 ± 0.007
+(quad1, quad3) 0.078 ± 0.016 0.107 ± 0.011 0.146 ± 0.013 0.323 ± 0.011 0.273 ± 0.007 0.349 ± 0.008 0.367 ± 0.016 0.372 ± 0.009 0.337 ± 0.008
+(quad1, quad4) 0.094 ± 0.032 0.091 ± 0.010 0.179 ± 0.014 0.346 ± 0.017 0.277 ± 0.017 0.325 ± 0.023 0.377 ± 0.018 0.343 ± 0.012 0.346 ± 0.010
+(quad1, size) 0.101 ± 0.010 0.087 ± 0.005 0.111 ± 0.010 0.266 ± 0.012 0.226 ± 0.013 0.284 ± 0.011 0.343 ± 0.005 0.331 ± 0.009 0.336 ± 0.009
+(quad1, orientation) 0.060 ± 0.003 0.059 ± 0.004 0.080 ± 0.002 0.236 ± 0.012 0.194 ± 0.017 0.240 ± 0.004 0.367 ± 0.016 0.360 ± 0.007 0.356 ± 0.007
+(quad2, quad3) 0.067 ± 0.007 0.098 ± 0.017 0.148 ± 0.014 0.338 ± 0.017 0.283 ± 0.014 0.325 ± 0.016 0.350 ± 0.010 0.349 ± 0.012 0.350 ± 0.008
+(quad2, quad4) 0.068 ± 0.004 0.078 ± 0.010 0.194 ± 0.018 0.302 ± 0.013 0.291 ± 0.015 0.332 ± 0.012 0.382 ± 0.007 0.342 ± 0.005 0.340 ± 0.013
+(quad2, size) 0.099 ± 0.008 0.083 ± 0.008 0.102 ± 0.007 0.248 ± 0.008 0.197 ± 0.004 0.281 ± 0.021 0.334 ± 0.007 0.328 ± 0.015 0.336 ± 0.006
+(quad2, orientation) 0.063 ± 0.004 0.058 ± 0.003 0.090 ± 0.012 0.225 ± 0.011 0.214 ± 0.010 0.274 ± 0.013 0.329 ± 0.011 0.348 ± 0.010 0.333 ± 0.008
+composition mean 0.081 ± 0.013 0.086 ± 0.009 0.132 ± 0.012 0.288 ± 0.012 0.248 ± 0.012 0.306 ± 0.014 0.355 ± 0.011 0.346 ± 0.009 0.341 ± 0.008
+Table 4: Evaluation of the context module, ablations, and baselines in terms of reconstruction, concept learning, and composition on quad (dependent).
+CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
+reconstruction (ELBO BPD)
+in-distribution 0.828 ± 0.004 0.840 ± 0.005 0.889 ± 0.017 0.767 ± 0.001 0.746 ± 0.000 0.752 ± 0.000 2.443 ± 0.002 2.408 ± 0.000 2.455 ± 0.002
+out-of-distribution 0.863 ± 0.006 0.908 ± 0.005 0.959 ± 0.011 0.815 ± 0.005 0.746 ± 0.001 0.753 ± 0.001 2.450 ± 0.002 2.406 ± 0.000 2.463 ± 0.002
+concept learning (sliced Wasserstein)
+obs 0.107 ± 0.005 0.103 ± 0.008 0.116 ± 0.003 0.110 ± 0.005 0.104 ± 0.006 0.113 ± 0.004 0.365 ± 0.007 0.375 ± 0.008 0.357 ± 0.005
+quad1 0.100 ± 0.004 0.102 ± 0.004 0.107 ± 0.008 0.158 ± 0.003 0.163 ± 0.004 0.137 ± 0.007 0.373 ± 0.004 0.352 ± 0.004 0.374 ± 0.008
+quad2 0.103 ± 0.004 0.100 ± 0.006 0.110 ± 0.009 0.179 ± 0.007 0.165 ± 0.008 0.173 ± 0.009 0.373 ± 0.004 0.359 ± 0.003 0.366 ± 0.005
+quad3 0.127 ± 0.017 0.127 ± 0.008 0.139 ± 0.008 0.200 ± 0.010 0.214 ± 0.009 0.213 ± 0.015 0.379 ± 0.005 0.374 ± 0.006 0.372 ± 0.009
+quad4 0.145 ± 0.010 0.140 ± 0.007 0.131 ± 0.011 0.305 ± 0.012 0.277 ± 0.010 0.278 ± 0.015 0.401 ± 0.015 0.386 ± 0.008 0.388 ± 0.013
+size 0.109 ± 0.006 0.105 ± 0.006 0.117 ± 0.004 0.116 ± 0.004 0.121 ± 0.007 0.119 ± 0.004 0.361 ± 0.006 0.362 ± 0.005 0.367 ± 0.005
+orientation 0.123 ± 0.003 0.109 ± 0.007 0.124 ± 0.004 0.120 ± 0.009 0.123 ± 0.008 0.130 ± 0.007 0.369 ± 0.006 0.353 ± 0.002 0.365 ± 0.007
+concept learning mean 0.116 ± 0.007 0.112 ± 0.007 0.121 ± 0.007 0.170 ± 0.007 0.167 ± 0.007 0.166 ± 0.009 0.375 ± 0.007 0.366 ± 0.005 0.370 ± 0.007
+composition (sliced Wasserstein)
+(quad1, quad2) 0.111 ± 0.008 0.097 ± 0.007 0.123 ± 0.011 0.229 ± 0.003 0.214 ± 0.005 0.196 ± 0.013 0.366 ± 0.006 0.381 ± 0.010 0.367 ± 0.007
+(quad1, quad3) 0.134 ± 0.008 0.117 ± 0.003 0.140 ± 0.008 0.227 ± 0.009 0.228 ± 0.008 0.198 ± 0.011 0.390 ± 0.009 0.367 ± 0.011 0.374 ± 0.004
+(quad1, quad4) 0.133 ± 0.007 0.163 ± 0.010 0.163 ± 0.008 0.284 ± 0.015 0.292 ± 0.006 0.276 ± 0.013 0.403 ± 0.012 0.376 ± 0.013 0.361 ± 0.006
+(quad1, size) 0.108 ± 0.003 0.106 ± 0.004 0.123 ± 0.007 0.172 ± 0.009 0.170 ± 0.009 0.157 ± 0.011 0.360 ± 0.008 0.357 ± 0.008 0.367 ± 0.003
+(quad1, orientation) 0.110 ± 0.007 0.108 ± 0.007 0.125 ± 0.006 0.167 ± 0.006 0.157 ± 0.007 0.143 ± 0.007 0.365 ± 0.010 0.371 ± 0.007 0.369 ± 0.004
+(quad2, quad3) 0.158 ± 0.007 0.150 ± 0.013 0.164 ± 0.005 0.222 ± 0.008 0.200 ± 0.011 0.195 ± 0.010 0.367 ± 0.007 0.378 ± 0.006 0.379 ± 0.008
+(quad2, quad4) 0.130 ± 0.011 0.122 ± 0.009 0.119 ± 0.010 0.299 ± 0.008 0.282 ± 0.025 0.279 ± 0.008 0.403 ± 0.010 0.404 ± 0.013 0.382 ± 0.014
+(quad2, size) 0.126 ± 0.009 0.109 ± 0.006 0.120 ± 0.005 0.191 ± 0.011 0.166 ± 0.010 0.181 ± 0.010 0.368 ± 0.004 0.365 ± 0.005 0.363 ± 0.004
+(quad2, orientation) 0.135 ± 0.004 0.120 ± 0.010 0.133 ± 0.006 0.208 ± 0.011 0.195 ± 0.020 0.198 ± 0.010 0.368 ± 0.005 0.366 ± 0.005 0.351 ± 0.006
+composition mean 0.127 ± 0.007 0.121 ± 0.008 0.134 ± 0.007 0.222 ± 0.009 0.212 ± 0.011 0.203 ± 0.010 0.377 ± 0.008 0.374 ± 0.009 0.368 ± 0.006
+Table 5: Evaluation of the baselines in terms of reconstruction, concept learning, and composition on 3DIdent.
+Ada-GVAE BetaTCVAE TVAE
+reconstruction (ELBO BPD)
+in-distribution 3.267 ± 0.001 3.261 ± 0.000 3.282 ± 0.005
+out-of-distribution 3.274 ± 0.001 3.264 ± 0.001 3.291 ± 0.006
+concept learning (sliced Wasserstein)
+obs 0.267 ± 0.011 0.257 ± 0.012 0.245 ± 0.003
+bg 0.232 ± 0.012 0.254 ± 0.009 0.262 ± 0.006
+obj 0.243 ± 0.007 0.263 ± 0.010 0.259 ± 0.007
+sl 0.253 ± 0.013 0.254 ± 0.007 0.250 ± 0.009
+concept learning mean 0.249 ± 0.011 0.257 ± 0.010 0.254 ± 0.006
+composition (sliced Wasserstein)
+(bg, obj) 0.252 ± 0.014 0.261 ± 0.011 0.256 ± 0.010
+(bg, sl) 0.227 ± 0.009 0.255 ± 0.012 0.241 ± 0.013
+(obj, sl) 0.274 ± 0.010 0.247 ± 0.005 0.269 ± 0.009
+composition mean 0.251 ± 0.011 0.254 ± 0.009 0.256 ± 0.011
+E.1.2 Examples of concept composition (quad)
+Figure LABEL:fig:quad-ood-full depicts samples generated from the context module appended to the lightweight-VAE described in Appendix D.3.
+[…]
+fig:quad-ood-full
 Figure 5: Example generated OOD images from a run of quad.
-D.1.3 Regularization (MNIST)
-A description of the architecture and experiment is given in Appendix C.
-Tables 7 and 8 respectively show results of group lasso (Yuan and Lin, 2006) and $\ell_{2}$ regularization, while Table 9 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017).
-[^2] All other experiments in Appendix D.1 are run with $\lambda=0$ for group lasso and $\ell_{2}$ and with $\beta=1$.
-Altogether, these results indicate relatively stable performance across regularization weights while suggesting that tuning these regularizers could be worthwhile (e.g., notice that $\lambda=10$ improves performance).
-Table 7: Group lasso regularization on MNIST
-Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.2983 (0.0053) 0.2907 (0.0081) 0.2869 (0.0064)
-Reconstruction Loss (bpd) ↓ 0.2683 (0.0047) 0.2605 (0.0088) 0.2568 (0.0074)
-Reconstructed Distribution Evaluation
-Covariance Error ↓ 1.2976 (0.0608) 1.2175 (0.0611) 1.3087 (0.0469)
-Mean Error ↓ 0.0196 (0.0020) 0.0134 (0.0018) 0.0178 (0.0023)
-MMD ↓ 0.0125 (0.0018) 0.0069 (0.0023) 0.0109 (0.0015)
-Covariance Error (ivn) ↓ 2.2654 (0.2437) 2.2651 (0.2638) 2.2818 (0.2705)
-Mean Error (ivn) ↓ 0.0277 (0.0018) 0.0277 (0.0025) 0.0315 (0.0031)
-MMD (ivn) ↓ 0.0993 (0.0367) 0.0982 (0.0356) 0.0986 (0.0358)
-Generated Distribution Evaluation
-MMD ↓ 0.0048 (0.0005) 0.0053 (0.0005) 0.0057 (0.0009)
-MMD (ivn) ↓ 0.0283 (0.0101) 0.0274 (0.0076) 0.0211 (0.0061)
-Table 8: $\ell_{2}$ regularization on MNIST
-Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.3079 (0.0138) 0.2917 (0.0063) 0.2966 (0.0084)
-Reconstruction Loss (bpd) ↓ 0.2774 (0.0145) 0.2612 (0.0067) 0.2667 (0.0092)
-Reconstructed Distribution Evaluation
-Covariance Error ↓ 1.3562 (0.1205) 1.2883 (0.0103) 1.4347 (0.0955)
-Mean Error ↓ 0.0178 (0.0021) 0.0182 (0.0018) 0.0209 (0.0023)
-MMD ↓ 0.0090 (0.0020) 0.0088 (0.0022) 0.0119 (0.0025)
-Covariance Error (ivn) ↓ 2.3289 (0.2539) 2.2137 (0.2341) 2.2341 (0.2606)
-Mean Error (ivn) ↓ 0.0287 (0.0022) 0.0281 (0.0021) 0.0298 (0.0031)
-MMD (ivn) ↓ 0.1023 (0.0377) 0.0957 (0.0348) 0.0881 (0.0317)
-Generated Distribution Evaluation
-MMD ↓ 0.0051 (0.0005) 0.0056 (0.0006) 0.0067 (0.0012)
-MMD (ivn) ↓ 0.0231 (0.0071) 0.0254 (0.0072) 0.0217 (0.0064)
-Table 9: Latent space regularization (as in $\beta$-VAE) on MNIST
-Metric / Evaluation $\beta$=0.5 $\beta$=2 $\beta$=4
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.2591 (0.0057) 0.3399 (0.0080) 0.3725 (0.0076)
-Reconstruction Loss (bpd) ↓ 0.2367 (0.0059) 0.3016 (0.0094) 0.3317 (0.0083)
-Reconstructed Distribution Evaluation
-Covariance Error ↓ 1.1730 (0.0739) 1.7058 (0.0478) 1.8174 (0.1025)
-Mean Error ↓ 0.0180 (0.0026) 0.0232 (0.0018) 0.0149 (0.0016)
-MMD ↓ 0.0031 (0.0017) 0.0268 (0.0033) 0.0688 (0.0105)
-Covariance Error (ivn) ↓ 1.9956 (0.2056) 2.3965 (0.2786) 2.6715 (0.3147)
-Mean Error (ivn) ↓ 0.0258 (0.0017) 0.0359 (0.0038) 0.0408 (0.0053)
-MMD (ivn) ↓ 0.0821 (0.0307) 0.1110 (0.0357) 0.1491 (0.0324)
-Generated Distribution Evaluation
-MMD ↓ 0.0038 (0.0001) 0.0132 (0.0014) 0.0437 (0.0086)
-MMD (ivn) ↓ 0.0160 (0.0041) 0.0410 (0.0112) 0.0898 (0.0179)
-D.1.4 Expressivity (MNIST)
-[…]
-[…]
-Metric / Evaluation (15, 22) (22, 22) (22, 5) (50, 22) (50, 5)
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.3777 (0.0100) 0.3892 (0.0119) 0.2956 (0.0032) 0.3655 (0.0088) 0.3073 (0.0098)
-Reconstruction Loss (bpd) ↓ 0.3532 (0.0095) 0.3641 (0.0123) 0.2653 (0.0034) 0.3413 (0.0083) 0.2781 (0.0098)
-Reconstructed Distribution Evaluation (Observations)
-MAE_Cov ↓ 1.6402 (0.1417) 1.7153 (0.1730) 1.4769 (0.0512) 1.6964 (0.1163) 1.5230 (0.1056)
-MAE_Mean ↓ 0.0206 (0.0032) 0.0214 (0.0020) 0.0243 (0.0019) 0.0215 (0.0023) 0.0194 (0.0015)
-MMD ↓ 0.0183 (0.0016) 0.0199 (0.0015) 0.0119 (0.0008) 0.0241 (0.0050) 0.0138 (0.0010)
-Reconstructed Distribution Evaluation (Others)
-MAE_Cov ↓ 2.8664 (0.3217) 2.9818 (0.3424) 2.3018 (0.2701) 2.7742 (0.3048) 2.3233 (0.2495)
-MAE_Mean ↓ 0.0407 (0.0035) 0.0455 (0.0044) 0.0311 (0.0028) 0.0374 (0.0032) 0.0305 (0.0023)
-MMD ↓ 0.1036 (0.0313) 0.1064 (0.0337) 0.0907 (0.0321) 0.1149 (0.0372) 0.0882 (0.0302)
-Generated Distribution Evaluation
-MMD ↓ 0.0094 (0.0006) 0.0091 (0.0015) 0.0061 (0.0008) 0.0110 (0.0021) 0.0070 (0.0017)
-MMD (ivn) ↓ 0.0312 (0.0066) 0.0422 (0.0130) 0.0224 (0.0062) 0.0388 (0.0117) 0.0249 (0.0059)
-[…]
-Metric / Evaluation (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
-Validation Metrics
-Validation ELBO (bpd) ↓ 0.3423 (0.0089) 0.3147 (0.0105) 0.3779 (0.0178) 0.3027 (0.0063) 0.3702 (0.0208) 0.3160 (0.0124)
-Reconstruction Loss (bpd) ↓ 0.3213 (0.0092) 0.2897 (0.0108) 0.3569 (0.0181) 0.2767 (0.0061) 0.3531 (0.0240) 0.2899 (0.0132)
-Reconstructed Distribution Evaluation (Observations)
-MAE_Cov ↓ 1.5155 (0.0766) 1.5477 (0.0886) 1.7862 (0.1572) 1.3584 (0.1025) 1.9985 (0.3544) 1.4030 (0.1516)
-MAE_Mean ↓ 0.0141 (0.0014) 0.0178 (0.0019) 0.0201 (0.0030) 0.0148 (0.0031) 0.0210 (0.0054) 0.0157 (0.0029)
-MMD ↓ 0.0251 (0.0021) 0.0180 (0.0026) 0.0282 (0.0039) 0.0137 (0.0018) 0.2226 (0.1939) 0.0139 (0.0030)
-Reconstructed Distribution Evaluation (Others)
-MAE_Cov ↓ 2.5968 (0.2735) 2.2916 (0.2427) 2.8177 (0.3093) 2.2404 (0.2412) 2.9055 (0.3518) 2.3451 (0.2558)
-MAE_Mean ↓ 0.0359 (0.0035) 0.0311 (0.0024) 0.0418 (0.0040) 0.0298 (0.0023) 0.0428 (0.0068) 0.0304 (0.0028)
-MMD ↓ 0.0937 (0.0279) 0.1036 (0.0359) 0.1112 (0.0337) 0.0987 (0.0350) 0.2750 (0.0712) 0.0946 (0.0333)
-Generated Distribution Evaluation
-MMD ↓ 0.0123 (0.0012) 0.0092 (0.0013) 0.0135 (0.0024) 0.0067 (0.0007) 0.2112 (0.1967) 0.0056 (0.0006)
-MMD (ivn) ↓ 0.0334 (0.0076) 0.0272 (0.0069) 0.0374 (0.0082) 0.0242 (0.0064) 0.2254 (0.0721) 0.0245 (0.0069)
-D.2 Additional NVAE Figures
-D.2.1 Additional examples of concept composition
-Figure D.2.1 contains additional examples of composing concepts from the NVAE-based context module. Each row corresponds to actual generated samples from the trained model in different contexts:
+E.1.3 Regularization (MNIST)
+A description of the architecture and experiment is given in Appendix D.
+Tables 6 and 7 respectively show results of group lasso (Yuan and Lin, 2006) and $L_{2}$ regularization, while Table 8 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017).
+[^1] All other MNIST experiments in Appendix E.1 are run with $\lambda=0$ for group lasso and $L_{2}$ and with $\beta=1$.
+Altogether, the results are relatively stable performance across regularization strengths, with slight increase in reconstruction performance for higher group lasso or $L_{2}$ regularization strength but no other statistically significant differences.
+Table 6: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different group lasso regularization strengths.
+$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
+reconstruction (ELBO BPD)
+in-distribution 0.259 ± 0.005 0.259 ± 0.003 0.253 ± 0.003 0.248 ± 0.002
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.099 ± 0.006 0.092 ± 0.005 0.091 ± 0.006 0.092 ± 0.006
+scaled 0.062 ± 0.002 0.069 ± 0.002 0.075 ± 0.003 0.118 ± 0.006
+shear 0.119 ± 0.006 0.118 ± 0.005 0.115 ± 0.008 0.113 ± 0.004
+shift 0.113 ± 0.005 0.112 ± 0.007 0.116 ± 0.005 0.113 ± 0.003
+swel 0.126 ± 0.007 0.122 ± 0.004 0.131 ± 0.007 0.127 ± 0.006
+thic 0.178 ± 0.007 0.183 ± 0.008 0.216 ± 0.015 0.220 ± 0.005
+thin 0.064 ± 0.004 0.072 ± 0.005 0.082 ± 0.009 0.111 ± 0.008
+concept learning mean 0.109 ± 0.006 0.110 ± 0.005 0.118 ± 0.008 0.128 ± 0.005
+Table 7: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different $L_{2}$-norm regularization strengths.
+$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
+reconstruction (ELBO BPD)
+in-distribution 0.259 ± 0.005 0.254 ± 0.003 0.255 ± 0.004 0.253 ± 0.001
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.099 ± 0.006 0.091 ± 0.006 0.102 ± 0.005 0.086 ± 0.004
+scaled 0.062 ± 0.002 0.068 ± 0.005 0.061 ± 0.003 0.073 ± 0.004
+shear 0.119 ± 0.006 0.119 ± 0.007 0.127 ± 0.005 0.111 ± 0.006
+shift 0.113 ± 0.005 0.111 ± 0.006 0.121 ± 0.005 0.108 ± 0.003
+swel 0.126 ± 0.007 0.121 ± 0.009 0.133 ± 0.005 0.115 ± 0.004
+thic 0.178 ± 0.007 0.186 ± 0.015 0.202 ± 0.010 0.183 ± 0.005
+thin 0.064 ± 0.004 0.072 ± 0.008 0.062 ± 0.003 0.074 ± 0.006
+concept learning mean 0.109 ± 0.006 0.110 ± 0.008 0.115 ± 0.005 0.107 ± 0.005
+Table 8: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different latent space regularization strengths, as $\beta$ varies.
+0.5 1.0 2.0 4.0
+reconstruction (ELBO BPD)
+in-distribution 0.222 ± 0.002 0.259 ± 0.005 0.298 ± 0.003 0.341 ± 0.007
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.085 ± 0.003 0.099 ± 0.006 0.102 ± 0.003 0.108 ± 0.005
+scaled 0.077 ± 0.006 0.062 ± 0.002 0.077 ± 0.004 0.098 ± 0.009
+shear 0.108 ± 0.003 0.119 ± 0.006 0.125 ± 0.005 0.117 ± 0.004
+shift 0.105 ± 0.003 0.113 ± 0.005 0.125 ± 0.004 0.130 ± 0.007
+swel 0.110 ± 0.007 0.126 ± 0.007 0.129 ± 0.006 0.134 ± 0.004
+thic 0.182 ± 0.011 0.178 ± 0.007 0.190 ± 0.016 0.192 ± 0.008
+thin 0.069 ± 0.006 0.064 ± 0.004 0.074 ± 0.005 0.117 ± 0.005
+concept learning mean 0.105 ± 0.006 0.109 ± 0.006 0.118 ± 0.006 0.128 ± 0.006
+E.1.4 Regularization (quad)
+This is like Appendix E.1.3 but applied to the quad dataset rather than MNIST.
+A description of the architecture and experiment is given in Appendix D.
+Tables 9 and 10 respectively show results of group lasso (Yuan and Lin, 2006) and $L_{2}$ regularization, while Table 11 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017).
+[^2] All other quad experiments in Appendix E.1 are run with $\lambda=0$ for group lasso and $L_{2}$ and with $\beta=1$.
+Table 9: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different group norm regularization strengths.
+$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
+reconstruction (ELBO BPD)
+in-distribution 0.519 ± 0.022 0.511 ± 0.023 0.528 ± 0.030 0.492 ± 0.009
+out-of-distribution 0.720 ± 0.069 0.703 ± 0.052 0.868 ± 0.147 0.993 ± 0.181
+concept learning (sliced Wasserstein)
+obs 0.056 ± 0.003 0.049 ± 0.001 0.054 ± 0.004 0.055 ± 0.002
+quad1 0.062 ± 0.003 0.058 ± 0.004 0.056 ± 0.005 0.060 ± 0.001
+quad2 0.061 ± 0.005 0.057 ± 0.006 0.057 ± 0.003 0.055 ± 0.004
+quad3 0.063 ± 0.002 0.058 ± 0.004 0.058 ± 0.003 0.061 ± 0.006
+quad4 0.065 ± 0.007 0.054 ± 0.003 0.052 ± 0.003 0.056 ± 0.003
+size 0.075 ± 0.007 0.076 ± 0.007 0.073 ± 0.007 0.069 ± 0.005
+orientation 0.057 ± 0.003 0.057 ± 0.003 0.055 ± 0.006 0.065 ± 0.006
+concept learning mean 0.063 ± 0.004 0.058 ± 0.004 0.058 ± 0.004 0.060 ± 0.004
+composition (sliced Wasserstein)
+(quad1, quad2) 0.100 ± 0.028 0.061 ± 0.008 0.092 ± 0.023 0.101 ± 0.038
+(quad1, quad3) 0.078 ± 0.016 0.080 ± 0.019 0.101 ± 0.017 0.134 ± 0.019
+(quad1, quad4) 0.094 ± 0.032 0.081 ± 0.017 0.127 ± 0.040 0.132 ± 0.043
+(quad1, size) 0.101 ± 0.010 0.098 ± 0.006 0.115 ± 0.011 0.127 ± 0.025
+(quad1, orientation) 0.060 ± 0.003 0.055 ± 0.004 0.057 ± 0.004 0.063 ± 0.005
+(quad2, quad3) 0.067 ± 0.007 0.072 ± 0.008 0.075 ± 0.008 0.123 ± 0.042
+(quad2, quad4) 0.068 ± 0.004 0.087 ± 0.029 0.075 ± 0.006 0.145 ± 0.024
+(quad2, size) 0.099 ± 0.008 0.097 ± 0.004 0.111 ± 0.019 0.108 ± 0.022
+(quad2, orientation) 0.063 ± 0.004 0.059 ± 0.004 0.058 ± 0.004 0.065 ± 0.008
+composition mean 0.081 ± 0.013 0.077 ± 0.011 0.090 ± 0.015 0.111 ± 0.025
+Table 10: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different $L_{2}$-norm regularization strengths.
+$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
+reconstruction (ELBO BPD)
+in-distribution 0.519 ± 0.022 0.529 ± 0.029 0.550 ± 0.029 0.507 ± 0.024
+out-of-distribution 0.720 ± 0.069 0.662 ± 0.058 0.676 ± 0.023 0.685 ± 0.040
+concept learning (sliced Wasserstein)
+obs 0.056 ± 0.003 0.054 ± 0.001 0.055 ± 0.002 0.054 ± 0.001
+quad1 0.062 ± 0.003 0.058 ± 0.003 0.061 ± 0.003 0.053 ± 0.004
+quad2 0.061 ± 0.005 0.054 ± 0.003 0.062 ± 0.002 0.056 ± 0.002
+quad3 0.063 ± 0.002 0.057 ± 0.002 0.059 ± 0.005 0.055 ± 0.005
+quad4 0.065 ± 0.007 0.053 ± 0.003 0.058 ± 0.002 0.052 ± 0.004
+size 0.075 ± 0.007 0.073 ± 0.009 0.076 ± 0.007 0.070 ± 0.004
+orientation 0.057 ± 0.003 0.058 ± 0.004 0.063 ± 0.004 0.056 ± 0.005
+concept learning mean 0.063 ± 0.004 0.058 ± 0.003 0.062 ± 0.003 0.057 ± 0.004
+composition (sliced Wasserstein)
+(quad1, quad2) 0.100 ± 0.028 0.070 ± 0.007 0.069 ± 0.003 0.058 ± 0.003
+(quad1, quad3) 0.078 ± 0.016 0.073 ± 0.014 0.096 ± 0.021 0.063 ± 0.005
+(quad1, quad4) 0.094 ± 0.032 0.101 ± 0.029 0.081 ± 0.012 0.079 ± 0.019
+(quad1, size) 0.101 ± 0.010 0.095 ± 0.011 0.098 ± 0.004 0.093 ± 0.007
+(quad1, orientation) 0.060 ± 0.003 0.056 ± 0.004 0.061 ± 0.002 0.058 ± 0.003
+(quad2, quad3) 0.067 ± 0.007 0.071 ± 0.009 0.080 ± 0.006 0.071 ± 0.009
+(quad2, quad4) 0.068 ± 0.004 0.072 ± 0.009 0.067 ± 0.003 0.063 ± 0.006
+(quad2, size) 0.099 ± 0.008 0.084 ± 0.006 0.097 ± 0.008 0.089 ± 0.005
+(quad2, orientation) 0.063 ± 0.004 0.058 ± 0.003 0.061 ± 0.003 0.061 ± 0.003
+composition mean 0.081 ± 0.013 0.076 ± 0.010 0.079 ± 0.007 0.071 ± 0.007
+Table 11: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different latent space regularization strengths, as $\beta$ varies.
+0.5 1.0 2.0 4.0
+reconstruction (ELBO BPD)
+in-distribution 0.474 ± 0.003 0.519 ± 0.022 0.599 ± 0.020 0.640 ± 0.029
+out-of-distribution 0.632 ± 0.028 0.720 ± 0.069 0.821 ± 0.047 0.865 ± 0.061
+concept learning (sliced Wasserstein)
+obs 0.050 ± 0.001 0.056 ± 0.003 0.048 ± 0.002 0.054 ± 0.004
+quad1 0.051 ± 0.003 0.062 ± 0.003 0.055 ± 0.002 0.064 ± 0.005
+quad2 0.052 ± 0.002 0.061 ± 0.005 0.052 ± 0.003 0.064 ± 0.011
+quad3 0.059 ± 0.002 0.063 ± 0.002 0.047 ± 0.002 0.064 ± 0.009
+quad4 0.054 ± 0.002 0.065 ± 0.007 0.050 ± 0.002 0.058 ± 0.006
+size 0.061 ± 0.003 0.075 ± 0.007 0.066 ± 0.005 0.082 ± 0.005
+orientation 0.050 ± 0.002 0.057 ± 0.003 0.051 ± 0.002 0.068 ± 0.008
+concept learning mean 0.054 ± 0.002 0.063 ± 0.004 0.053 ± 0.003 0.065 ± 0.007
+composition (sliced Wasserstein)
+(quad1, quad2) 0.059 ± 0.005 0.100 ± 0.028 0.080 ± 0.005 0.082 ± 0.015
+(quad1, quad3) 0.070 ± 0.008 0.078 ± 0.016 0.071 ± 0.007 0.082 ± 0.014
+(quad1, quad4) 0.076 ± 0.010 0.094 ± 0.032 0.073 ± 0.012 0.123 ± 0.027
+(quad1, size) 0.099 ± 0.007 0.101 ± 0.010 0.118 ± 0.008 0.115 ± 0.010
+(quad1, orientation) 0.052 ± 0.002 0.060 ± 0.003 0.057 ± 0.003 0.073 ± 0.005
+(quad2, quad3) 0.056 ± 0.000 0.067 ± 0.007 0.074 ± 0.008 0.094 ± 0.021
+(quad2, quad4) 0.087 ± 0.018 0.068 ± 0.004 0.101 ± 0.027 0.067 ± 0.009
+(quad2, size) 0.086 ± 0.015 0.099 ± 0.008 0.121 ± 0.018 0.131 ± 0.011
+(quad2, orientation) 0.055 ± 0.003 0.063 ± 0.004 0.056 ± 0.006 0.068 ± 0.009
+composition mean 0.071 ± 0.008 0.081 ± 0.013 0.083 ± 0.010 0.093 ± 0.013
+E.1.5 Expressivity (MNIST)
+A description of the architecture and experiment is given in Appendix D.
+Tables 12 and 13 show varying widths for an expressive layer depth respectively of 2 and 5: the tuples in the column headers indicate $(0pt_{\mathrm{exp}},\ 0pt_{c})$.
+[^3] All other experiments on MNIST in Appendix E.1 are run with $h_{\mathrm{exp}}=2$, $0pt_{\mathrm{exp}}=15$, and $0pt_{c}=5$.
+These results suggest relatively stable results across different configurations for concept learning while the best reconstruction performance is obtained from a moderate bottleneck, with expressive layer input width $0pt_{\mathrm{exp}}=22$ and a concept width of $0pt_{c}=5$.
+Table 12: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different capacities, with depth $h_{exp}=2$ as $(w_{exp},w_{c})$ varies.
+$(w_{exp},w_{c})$ (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
+reconstruction (ELBO BPD)
+in-distribution 0.260 ± 0.005 0.259 ± 0.005 0.255 ± 0.003 0.252 ± 0.003 0.265 ± 0.001 0.265 ± 0.006
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.087 ± 0.003 0.099 ± 0.006 0.082 ± 0.005 0.095 ± 0.006 0.086 ± 0.004 0.092 ± 0.008
+scaled 0.073 ± 0.011 0.062 ± 0.002 0.065 ± 0.004 0.067 ± 0.005 0.068 ± 0.003 0.072 ± 0.004
+shear 0.115 ± 0.006 0.119 ± 0.006 0.108 ± 0.006 0.122 ± 0.006 0.116 ± 0.004 0.118 ± 0.004
+shift 0.106 ± 0.009 0.113 ± 0.005 0.108 ± 0.008 0.111 ± 0.002 0.096 ± 0.003 0.115 ± 0.005
+swel 0.119 ± 0.006 0.126 ± 0.007 0.107 ± 0.005 0.118 ± 0.006 0.111 ± 0.007 0.118 ± 0.007
+thic 0.202 ± 0.016 0.178 ± 0.007 0.180 ± 0.007 0.171 ± 0.014 0.176 ± 0.010 0.178 ± 0.015
+thin 0.065 ± 0.006 0.064 ± 0.004 0.069 ± 0.005 0.072 ± 0.010 0.073 ± 0.007 0.067 ± 0.006
+concept learning mean 0.110 ± 0.008 0.109 ± 0.006 0.103 ± 0.006 0.108 ± 0.007 0.104 ± 0.005 0.109 ± 0.007
+Table 13: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different capacities, with depth $h_{exp}=4$ as $(w_{exp},w_{c})$ varies.
+$(w_{exp},w_{c})$ (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
+reconstruction (ELBO BPD)
+in-distribution 0.269 ± 0.003 0.272 ± 0.006 0.276 ± 0.004 0.255 ± 0.004 0.281 ± 0.005 0.279 ± 0.008
+out-of-distribution
+concept learning (sliced Wasserstein)
+obs 0.096 ± 0.004 0.088 ± 0.004 0.089 ± 0.001 0.092 ± 0.004 0.090 ± 0.005 0.093 ± 0.003
+scaled 0.063 ± 0.003 0.071 ± 0.004 0.071 ± 0.003 0.085 ± 0.018 0.067 ± 0.001 0.079 ± 0.003
+shear 0.119 ± 0.006 0.110 ± 0.004 0.116 ± 0.004 0.113 ± 0.004 0.107 ± 0.004 0.103 ± 0.004
+shift 0.114 ± 0.004 0.109 ± 0.006 0.108 ± 0.005 0.107 ± 0.005 0.113 ± 0.006 0.104 ± 0.002
+swel 0.120 ± 0.006 0.113 ± 0.007 0.110 ± 0.001 0.122 ± 0.006 0.116 ± 0.005 0.115 ± 0.003
+thic 0.179 ± 0.012 0.178 ± 0.014 0.175 ± 0.011 0.185 ± 0.011 0.188 ± 0.007 0.184 ± 0.011
+thin 0.066 ± 0.005 0.076 ± 0.004 0.072 ± 0.003 0.070 ± 0.003 0.076 ± 0.007 0.083 ± 0.006
+concept learning mean 0.108 ± 0.005 0.106 ± 0.006 0.106 ± 0.004 0.111 ± 0.007 0.108 ± 0.005 0.109 ± 0.005
+E.1.6 Expressivity (quad)
+This is like Appendix E.1.5 but applied to the quad dataset rather than MNIST.
+A description of the architecture and experiment is given in Appendix D.
+Tables 14 and 15 show varying widths for an expressive layer depth respectively of 2 and 5: the tuples in the column headers indicate $(0pt_{\mathrm{exp}},\ 0pt_{c})$.
+[^4] All other experiments on quad in Appendix E.1 are run with $h_{\mathrm{exp}}=2$, $0pt_{\mathrm{exp}}=32$, and $0pt_{c}=8$.
+These results suggest relatively stable results across different configurations.
+Table 14: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different capacities, with depth $h_{exp}=2$ as $(w_{exp},w_{c})$ varies.
+$(w_{exp)},w_{c)})$ (16, 16) (16, 32) (16, 8) (32, 16) (32, 32) (32, 8)
+reconstruction (ELBO BPD)
+in-distribution 0.494 ± 0.014 0.497 ± 0.012 0.494 ± 0.010 0.512 ± 0.023 0.557 ± 0.016 0.519 ± 0.022
+out-of-distribution 0.676 ± 0.024 0.665 ± 0.025 0.598 ± 0.025 0.626 ± 0.030 0.752 ± 0.067 0.720 ± 0.069
+concept learning (sliced Wasserstein)
+obs 0.055 ± 0.005 0.056 ± 0.002 0.056 ± 0.004 0.053 ± 0.003 0.063 ± 0.005 0.056 ± 0.003
+quad1 0.057 ± 0.004 0.061 ± 0.005 0.059 ± 0.004 0.058 ± 0.004 0.067 ± 0.007 0.062 ± 0.003
+quad2 0.061 ± 0.006 0.053 ± 0.003 0.060 ± 0.003 0.056 ± 0.005 0.063 ± 0.005 0.061 ± 0.005
+quad3 0.058 ± 0.004 0.058 ± 0.005 0.065 ± 0.007 0.055 ± 0.003 0.066 ± 0.006 0.063 ± 0.002
+quad4 0.060 ± 0.007 0.063 ± 0.004 0.063 ± 0.006 0.055 ± 0.003 0.068 ± 0.006 0.065 ± 0.007
+size 0.072 ± 0.008 0.072 ± 0.007 0.076 ± 0.005 0.071 ± 0.005 0.082 ± 0.011 0.075 ± 0.007
+orientation 0.061 ± 0.006 0.055 ± 0.002 0.058 ± 0.003 0.055 ± 0.003 0.062 ± 0.006 0.057 ± 0.003
+concept learning mean 0.061 ± 0.006 0.060 ± 0.004 0.062 ± 0.005 0.057 ± 0.004 0.067 ± 0.007 0.063 ± 0.004
+composition (sliced Wasserstein)
+(quad1, quad2) 0.091 ± 0.010 0.086 ± 0.010 0.066 ± 0.006 0.076 ± 0.014 0.131 ± 0.017 0.100 ± 0.028
+(quad1, quad3) 0.064 ± 0.008 0.083 ± 0.006 0.075 ± 0.012 0.059 ± 0.005 0.102 ± 0.010 0.078 ± 0.016
+(quad1, quad4) 0.082 ± 0.019 0.069 ± 0.006 0.064 ± 0.007 0.076 ± 0.012 0.092 ± 0.013 0.094 ± 0.032
+(quad1, size) 0.084 ± 0.007 0.098 ± 0.009 0.094 ± 0.005 0.107 ± 0.013 0.117 ± 0.011 0.101 ± 0.010
+(quad1, orientation) 0.059 ± 0.004 0.065 ± 0.006 0.060 ± 0.004 0.055 ± 0.003 0.068 ± 0.011 0.060 ± 0.003
+(quad2, quad3) 0.083 ± 0.017 0.080 ± 0.014 0.076 ± 0.015 0.066 ± 0.008 0.083 ± 0.012 0.067 ± 0.007
+(quad2, quad4) 0.079 ± 0.010 0.074 ± 0.007 0.071 ± 0.010 0.065 ± 0.009 0.077 ± 0.008 0.068 ± 0.004
+(quad2, size) 0.100 ± 0.014 0.085 ± 0.011 0.087 ± 0.006 0.093 ± 0.006 0.116 ± 0.011 0.099 ± 0.008
+(quad2, orientation) 0.062 ± 0.006 0.057 ± 0.003 0.060 ± 0.005 0.057 ± 0.003 0.067 ± 0.008 0.063 ± 0.004
+composition mean 0.078 ± 0.011 0.077 ± 0.008 0.072 ± 0.008 0.073 ± 0.008 0.095 ± 0.011 0.081 ± 0.013
+Table 15: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different capacities, with depth $h_{exp}=4$ as $(w_{exp},w_{c})$ varies.
+$(w_{exp},w_{c})$ (16, 16) (16, 32) (16, 8) (32, 16) (32, 32) (32, 8)
+reconstruction (ELBO BPD)
+in-distribution 0.511 ± 0.008 0.516 ± 0.015 0.573 ± 0.045 0.590 ± 0.009 0.613 ± 0.020 0.580 ± 0.017
+out-of-distribution 0.761 ± 0.040 0.750 ± 0.033 0.674 ± 0.049 0.706 ± 0.035 0.759 ± 0.081 0.636 ± 0.017
+concept learning (sliced Wasserstein)
+obs 0.059 ± 0.004 0.052 ± 0.002 0.052 ± 0.003 0.053 ± 0.002 0.059 ± 0.006 0.065 ± 0.004
+quad1 0.061 ± 0.005 0.054 ± 0.004 0.061 ± 0.005 0.063 ± 0.002 0.081 ± 0.013 0.075 ± 0.005
+quad2 0.057 ± 0.004 0.057 ± 0.002 0.058 ± 0.003 0.064 ± 0.004 0.080 ± 0.014 0.071 ± 0.005
+quad3 0.059 ± 0.006 0.058 ± 0.003 0.064 ± 0.003 0.066 ± 0.004 0.078 ± 0.014 0.073 ± 0.002
+quad4 0.056 ± 0.004 0.053 ± 0.002 0.064 ± 0.003 0.058 ± 0.004 0.075 ± 0.012 0.071 ± 0.002
+size 0.069 ± 0.003 0.081 ± 0.005 0.077 ± 0.008 0.097 ± 0.004 0.111 ± 0.016 0.087 ± 0.006
+orientation 0.059 ± 0.004 0.055 ± 0.003 0.061 ± 0.005 0.060 ± 0.003 0.071 ± 0.012 0.071 ± 0.003
+concept learning mean 0.060 ± 0.004 0.059 ± 0.003 0.063 ± 0.004 0.066 ± 0.003 0.079 ± 0.012 0.073 ± 0.004
+composition (sliced Wasserstein)
+(quad1, quad2) 0.071 ± 0.008 0.090 ± 0.010 0.059 ± 0.004 0.089 ± 0.009 0.091 ± 0.016 0.074 ± 0.005
+(quad1, quad3) 0.062 ± 0.006 0.062 ± 0.004 0.080 ± 0.023 0.075 ± 0.005 0.086 ± 0.014 0.079 ± 0.006
+(quad1, quad4) 0.112 ± 0.025 0.072 ± 0.010 0.063 ± 0.009 0.077 ± 0.004 0.086 ± 0.010 0.083 ± 0.005
+(quad1, size) 0.083 ± 0.003 0.119 ± 0.008 0.102 ± 0.007 0.106 ± 0.003 0.133 ± 0.008 0.112 ± 0.005
+(quad1, orientation) 0.059 ± 0.003 0.057 ± 0.003 0.062 ± 0.006 0.064 ± 0.003 0.079 ± 0.011 0.074 ± 0.004
+(quad2, quad3) 0.077 ± 0.011 0.072 ± 0.002 0.061 ± 0.004 0.083 ± 0.008 0.109 ± 0.028 0.076 ± 0.006
+(quad2, quad4) 0.092 ± 0.016 0.083 ± 0.017 0.065 ± 0.004 0.081 ± 0.005 0.102 ± 0.021 0.075 ± 0.007
+(quad2, size) 0.091 ± 0.006 0.108 ± 0.009 0.101 ± 0.016 0.128 ± 0.012 0.135 ± 0.021 0.108 ± 0.007
+(quad2, orientation) 0.058 ± 0.004 0.062 ± 0.003 0.064 ± 0.002 0.067 ± 0.005 0.081 ± 0.015 0.073 ± 0.005
+composition mean 0.078 ± 0.009 0.080 ± 0.007 0.073 ± 0.008 0.086 ± 0.006 0.100 ± 0.016 0.084 ± 0.005
+E.2 Additional NVAE Results
+E.2.1 Additional examples of concept composition
+Figure 15 contains additional examples of composing concepts from the NVAE-based context module. Each row corresponds to actual generated samples from the trained model in different contexts:
   1. 1.
@@
     The final row shows generated samples where two distinct concepts are composed together.
-For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon.
-Recall that CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data.
-Additional examples of both single- and multi-concept interventions over more contexts are provided in Appendices D.2.2D.2.4.
-[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Figure 8: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right).
-In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together.
-The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
-D.2.2 Additional MNIST figures
-Figure 9: Examples of single-node concept interventions in MNIST. (top) observational, scaled, shear (bottom) thic, thin, swel.
-Figure 10: Examples of OOD multi-node concept interventions in MNIST.
-D.2.3 Additional 3DIdent figures
-[Uncaptioned image][Uncaptioned image]
-Figure 11: Examples of single-node concept interventions in 3DIdent. (top) observational, background (bottom) object, spotlight.
-Figure 12: Examples of OOD multi-node concept interventions in 3DIdent.
-D.2.4 Additional CelebA figures
-[Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Figure 13: Examples of single-node concept interventions in CelebA. (top) observational, bags under eyes, black hair (bottom) blond hair, male, mouth slightly open.
-[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Figure 14: Examples of OOD multi-node concept interventions in CelebA.
+The final row of compositional generations is genuinely OOD, in that the training data does not contain any examples where both concepts have been intervened upon.
+Additional examples of both concept learning (sampling after single-target interventions) and composition (sampling after multi-target interventions) are provided in Appendices E.2.2E.2.3.
+Figure 15: Additional examples of concept composition in MNIST (left) and 3DIdent (right).
+E.2.2 Additional MNIST figures
+See Figure 16 for concept learning and Figure 17 for composition.
+Figure 16: Examples of learned concepts (sampling after single-target intervention) in MNIST. (top) observational, scaled, shear (bottom) thic, thin, swel.
+Figure 17: Examples of OOD concept composition (sampling after multi-target intervention) in MNIST.
+E.2.3 Additional 3DIdent figures
+See Figure 18 for concept learning and Figure 19 for composition.
+Figure 18: Examples of learned concepts (sampling after single-target intervention) in 3DIdent. (top) observational, background (bottom) object, spotlight.
+Figure 19: Examples of OOD concept composition (sampling after multi-target intervention) in 3DIdent.